### v1

# 2cm2cm1cm1cm

Abstract
Quantum interactions exchanging different types of particles play a pivotal rôle in quantum many-body theory, but they are not sufficiently investigated from a mathematical perspective. Here, we consider a system made of two fermions and one boson, in order to study the effect of such an off-diagonal interaction term, having in mind the physics of cuprate superconductors. Additionally, our model also includes a generalized Hubbard interaction (i.e., a general local repulsion term for the fermions). Regarding pairing, exponentially localized dressed bound fermion pairs are shown to exist and their effective dispersion relation is studied in detail. Scattering properties of the system are derived for two channels: the unbound and bound pair channels. We give particular attention to the regime of very large on-site (Hubbard)  repulsions, because this situation is relevant for cuprate superconductors.
Keywords: exchange interactions, scattering theory, few-body quantum problem, cuprate superconductivity.
2020 AMS Subject Classification: 47B93, 81Q10
1 Introduction
Off-diagonal interaction terms of the form
$B^{\ast}A+A^{\ast}B\ ,$ (1)
with $A,B$ being two monomials of annihilation operators of two species ($a$) and ($b$) of quantum particles, play a pivotal rôle in the rigorous understanding of quantum many-body systems at low temperatures. Such terms are also named “exchange” terms, because they encode (quantum) processes destroying a set of particles of one specie to create another kind of particles.
For instance, for the Bogoliubov model, an off-diagonal term of the form
$\sum_{k}f_{1}\left(k\right)\left(b_{k}^{\ast}b_{-k}^{\ast}a^{2}+\left(a^{\ast}\right)^{2}b_{k}b_{-k}\right)\ ,\qquad f_{1}\left(k\right)\geq 0\ ,$ (2)
exchanging two bosons ($a=b_{0}$) having zero momentum ($k=0$) with a pair of boson having non-zero momentum of opposite sign ($b_{k\neq 0}$), is shown in [1] to imply a non-conventional Bose condensation. Made of dressed bound pairs of (zero-momentum) bosons, the non-conventional condensate is structurally different from the Bose-Einstein condensate of the ideal Bose gas. In particular, it must be depleted to take advantage of the effective attraction induced by the exchange interaction (2). See [2]. This is reminiscent of liquid helium physics, where 100% superfluid helium occurs at zero temperature with only 9% Bose condensate [3, 4, 5, 6, 7]. Off-diagonal interaction terms (2) are conjectured in [8] to be relevant to explain the macroscopic behavior of weakly interacting Bose gases.
Another example from quantum statistical mechanics is given by the spin-boson model within the so-called “rotating wave approximation”. In this approximation the model has terms of the form
$\sum_{k}f_{2}\left(k\right)\left(b_{k}^{\ast}\sigma_{-}+\sigma_{+}b_{k}\right)\ ,\qquad f_{2}\left(k\right)\geq 0\ ,$
with $\sigma_{\pm}=\sigma_{x}\pm\sigma_{y}$ ($\sigma_{x},\sigma_{y}$ being Pauli matrices) and $b_{k}$ being the annihilation operator of a boson. Note that $\sigma_{-}$ ($\sigma_{+}$) can be related to the annihilation (creation) operator $a$ ($a^{\ast}$) of a fermion, via a so-called Jordan-Wigner transformation. Such off-diagonal terms make impossible the diagonalization of the quantum Hamiltonian with usual methods. In particular, the impact of these interaction terms on the properties of the model is expected to be major. For a general presentation of spin-boson models, see, e.g., [9, Introduction and Section 2.3].
More recently, using the Hubbard model with nearest neighbor interaction near its Hartree-Fock ground state, Bach and Rauch demonstrate [10] that interaction terms of the form
$\sum_{x,y}\sum_{\mathrm{s},\mathrm{t}\in\{\uparrow,\downarrow\}}f_{3}\left(x-y\right)\left(b_{x,\mathrm{s}}^{\ast}b_{y,\mathrm{t}}^{\ast}a_{y,\mathrm{t}}a_{x,\mathrm{s}}+a_{x,\mathrm{s}}^{\ast}a_{y,\mathrm{t}}^{\ast}b_{y,\mathrm{t}}b_{x,\mathrm{s}}\right)\ ,\qquad f_{2}\left(x-y\right)\geq 0\ ,$ (3)
exchanging fermions inside ($a$) and outside ($b$) of the Fermi surface are the only ones that can prevent from getting uniform [^1] We mean a relative bound that is uniform with respect to the length of the discrete $d$-dimensional torus where the Hubbard model is defined. relative bounds of the effective interaction with respect to the effective kinetic energy. See [10, Theorems III.1, III.2 and III.3] for more details. In other words, (3) should again have a drastic impact on the corresponding quantum many-body system.
In the present paper, for a fairly general function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ we study the effect of the off-diagonal interaction term
${\sum\limits_{x,y}}\upsilon\left(x-y\right)\left(c_{y}^{\ast}\,b_{x}+b_{x}^{\ast}c_{y}\right)\ ,$ (4)
where $b_{x}$ is the annihilation operator of a spinless boson on the site $x$ of the two-dimensional (square) lattice $\mathbb{Z}^{2}$, while $c_{y}$ represents the annihilation of a fermion pair of zero total spin, the two components of which are spread around the lattice position $y\in\mathbb{Z}^{2}$. The purely fermionic part of the considered model corresponds to the extended Hubbard Hamiltonian, as used in the context of ultracold atoms, ions, and molecules [11], while the purely bosonic component refers to an ideal gas, i.e., it has only a kinetic part (or “hopping term”), without interbosonic interactions. Because of the fermionic part, which is not exactly diagonalizable, the behavior of the full quantum many-body system, outside perturbative regimes, is almost inaccessible with the mathematical tools at our disposal. We thus consider only a three-body problem, by restricting the model to the sector of one boson and two fermions of opposite spins. In fact, the system restricted to this particular sector is very interesting, both mathematically and physically.
Physically, the model is related to cuprate superconductors, like for instance $\mathrm{La}_{2-x}\mathrm{Sr}_{x}\mathrm{CuO}_{4}$ (LaSr 214) and $\mathrm{La}_{2-x}\mathrm{Ba}_{x}\mathrm{CuO}_{4}$. It is known [12, 13, 14] that in such crystals charge transport occurs within two-dimensional isotropic layers of copper oxides. This is why we consider here quantum particles on $2$-dimensional lattices $\mathbb{Z}^{2}$. A convincing microscopic mechanism behind superconductivity at high critical temperature is still lacking even after almost four decades of intensive theoretical and experimental studies. Many physicists believe that the celebrated Hubbard model could be pivotal, one way or another, in order to get a microscopic theory of high-temperature superconductivity, but many alternative explanations or research directions have also been considered in theoretical physics. For some of the more popular models for cuprate superconductors, see, e.g., [13, Chap. 7].
In many theoretical approaches to this problem, the existence of polaronic quasiparticles in relation with the very strong Jahn-Teller (JT) effect associated with copper ions is neglected, as stressed in [15, Part VII]. The role of polarons is however highlighted in [16], since the JT effect actually led to the discovery of superconductivity in cuprates in 1986. See [17, p. 2] or [18, 19]. Our theoretical approach differs from most popular ones, being based on the existence of JT bipolarons in copper oxides, as is discussed in the literature [20] at least as early as 1990. The physics behind this approach is explained in detail in [21], where a simplified version of the model studied here is considered. In our microscopic model for cuprate superconductors, as presented in [22, 21], the bosonic operator $b_{x}$ ($b_{x}^{\ast}$) in (4) refers to the annihilation (creation) of a JT bipolaron, whereas the fermionic one $c_{y}$ ($c_{y}^{\ast}$) annihilates (creates) a fermion pair, which is reminiscent of Cooper pairs in conventional superconductivity.
As in the present paper, no ad hoc assumptions, in particular concerning anisotropy, are made in [21]. In fact, [21] proves that unconventional pairing may occur, breaking spontaneously discrete symmetries of the model, like the $d$-wave pairing, whose wave function is antisymmetric with respect to $90^{\circ}$-rotations. It turns out that electrostatic (screened Coulomb) repulsion is crucial for such unconventional pairings, which are meanwhile shown to be concomitant with a strong depletion of superconducting pairs. Notice that the results of [21] are coherent with experimental observations on the cuprate LaSr 214: The coherence length at optimal doping and the $d$-wave pair formation in the pseudogap regime, i.e., at temperatures much higher than the superconducting transition temperature, are predicted in good accordance with experimental data.
Mathematically, the present paper improves [21, 22] to get more complete and general rigorous results, including, among other things, extended Hubbard interactions and scattering properties. While [21, 22] focus only on the ground state energy and the unconventional pairings in the limit of large Hubbard interactions, here we provide the full spectral properties of the corresponding Hamiltonian. In particular we study in depth the effective dispersion relation associated with dressed bound fermion pairs. It confirms that off-diagonal interactions of the form (1) produce bounded states by reducing the energy of the system, similar to [2], possibly with a spectral gap.
This was already done in [21, 22], but only for usual (non-extended) Hubbard interactions and one-range creation / annihilation operators of fermion pairs. Even in this specific case, the dispersion relation of dressed bound fermion pairs was analyzed only to a level of detail enough to deduce unconventional pairings near the ground state. By contrast, in the present paper other important properties of the dispersion relation, like its regularity, are studied for the first time and in a more general framework. Using Combes-Thomas estimates we show, among other things, that the dressed bound fermion pairs are localized, in the sense that the fermion-fermion correlation decays very fast in space. Group velocities and tensor masses of dressed bound fermion pairs are also shown to exist under very natural conditions on the (absolutely summable function) $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ appearing in (4). In fact, our analysis allows one to accurately understand which features of the exchange strength function $\upsilon$ can strengthen the stability of the dressed bound fermion pairs. For instance, $\upsilon$ has to be sufficiently strong and localized in Fourier space in order to get a sufficiently strong “gluing effect”. Additionally, the boson should be heavier than two fermions. Notice that this second condition is consistent with the physical interpretation that the boson is a bipolaron, which is known to be (effectively) much heavier than the fermions (electrons or holes), in superconducting cuprates. Observe additionally that the very large mass of bipolarons (and polarons, in general), is one of the main arguments used to discredit theoretical approaches based on bipolarons, because it is known from experiments that the charge carriers in superconducting cuprates have an effective mass comparable to that of electrons and holes. In fact, we prove that the effective mass of bound pairs mainly depends on the properties of the function $\upsilon$, that encodes the fermion-boson exchange processes, but not much on the mass of the boson itself. This issue is discussed in [21], in detail. See also the discussion at the end of Section 3.2. That is why we are interested in results concerning the mass tensor for bound pairs and we think we provide here a convincing solution for the “mass paradox” related to bipolaronic pairing mechanisms in the microscopic theory of cuprate superconductors.
We also study here scattering properties of the three-body model in two channels: the unbound and bound pair channels. The first one corresponds to the case in which two fermions start far apart from each other and only experience a very weak repulsion force due to the extended Hubbard interaction, while the probability that they bind together to form a boson is very small. In this situation, we show that two (almost) freely propagating fermions in the distant past can come together and interact with one another, either via the repulsive electrostatic force or by exchanging a boson, and then propagate away, again freely in the distant future. In this channel, the scattering matrix can be explicitly computed via convergent (Dyson) series, making in particular the study of the scattering effect of the fermion-boson-exchange interaction (1) technically uncomplicated.
The bound pair (scattering) channel corresponds to the case in which dressed bound fermion pairs are formed. In contrast with the first channel, now there is a non-negligible bosonic component related with the exchanged boson that “glues” the two fermions together. We prove that those (spatially localized) dressed bound fermion pairs effectively move like a free (quantum spinless) particle. In this case, strictly speaking in the physical sense, there is no scattering and the pairs evolve freely in space, governed by an effective dispersion relation, the Fourier transform of which is the effective hopping strength for the (spatially localized) dressed bound pairs.
We additionally prove that all these properties hold also true in the limit of large on-site fermionic repulsions, provided that two fermions on two different lattice sites can interact via the fermion-boson exchange interaction. It refers to a hard core limit, preventing two fermions from occupying the same lattice site. For cuprate superconductors, it is an important issue addressed and answered here, because of the undeniable experimental evidence of very strong on-site Coulomb repulsions in cuprates, leading to the universally observed Mott transition at zero doping [23, 24].
To conclude, the mathematical properties of the model studied in the present work are well understood and, as a consequence, the model can serve as a prototypical example of a quantum system including exchange interaction terms of the form (1). From a physics viewpoint, it is also interesting, since dressed bound fermion pairs are good candidates for superconducting charge carriers in cuprate superconductors, as advocated in [21]. Our main results are Theorems 3.1, 3.5, 3.6, 3.9, 3.11 and 3.14. The paper is organized as follows: Section 2 explains in detail the model, while Section 3 gives the main results. Technical outcomes, along with all their proofs, are gathered in Section 4. Section 5 is an appendix collecting some important, albeit well-known, mathematical results used here, making the paper self-contained.
Remark 1.1 (Notation)
For any normed vector space $\mathcal{X}$ over $\mathbb{C}$, we omit the subscript $\mathcal{X}$ to denote its norm $\|\cdot\|\equiv\|\cdot\|_{\mathcal{X}}$, unless there is any risk of confusion. Mutatis mutandis for the scalar product $\langle\cdot,\cdot\rangle\equiv\langle\cdot,\cdot\rangle_{\mathcal{X}}$ in Hilbert spaces. As is usual, $\mathcal{B}(\mathcal{X},\mathcal{Y})$ denotes the set of bounded (linear) operators $\mathcal{X}\rightarrow\mathcal{Y}$ between two normed spaces $\mathcal{X}$ to $\mathcal{Y}$. If $\mathcal{X}=\mathcal{Y}$, $\mathcal{B}(\mathcal{X})\equiv\mathcal{B}(\mathcal{X},\mathcal{X})$ and its (operator) norm and its identity are respectively denoted by $\|\cdot\|_{\mathrm{op}}\equiv\|\cdot\|_{\mathcal{B}(\mathcal{X})}$ and $\mathfrak{1}_{\mathcal{X}}\equiv\mathfrak{1}$. $\mathbb{R}_{0}^{+}$ denotes the set of positive real numbers including zero, whereas $\mathbb{R}^{+}\doteq\mathbb{R}_{0}^{+}\backslash\{0\}$ is the set of strictly positive real numbers.
2 Setup of the Problem
2.1 Background Lattice
As explained in [12, 13, 14], charge transport in cuprates a priori occurs within two-dimensional isotropic $\mathrm{CuO}_{2}$ layers made of $\mathrm{Cu}^{++}$ and $\mathrm{O}^{--}$ ions, even if a weak inter-layer interaction might also be relevant. The (two-dimensional) lattice composed by copper and oxygen ions is supposed to have the symmetries of the square, as is usual. In other words, it is invariant under the group generated by $90^{\circ}$-degree rotations. This is experimentally verified in the important family of tetragonal cuprates like $\mathrm{La}_{2-x}\mathrm{Sr}_{x}\mathrm{CuO}_{4}$ (LaSr 214) and $\mathrm{La}_{2-x}\mathrm{Ba}_{x}\mathrm{CuO}_{4}$. See, e.g., [25, Section 9.1.2], [12, Section 2.3] and [14, Section 6.3.1].
Having in mind these physical observations on cuprates, we consider here quantum particles on lattices $\mathbb{Z}^{2}$. It means in particular that, (disregarding internal degrees of freedom of the quantum particles, like their spin) the (separable) Hilbert space $\ell^{2}(\mathbb{Z}^{2})$ is the “one-particle space” associated with the physical system we are interested in. Its canonical orthonormal basis is $\{\mathfrak{e}_{x}\}_{x\in\mathbb{Z}^{2}}$:
$\mathfrak{e}_{x}\left(y\right)\doteq\delta_{x,y}\ ,\qquad x,y\in\mathbb{Z}^{2}\ ,$ (5)
where $\delta_{\mathfrak{i},\mathfrak{j}}$ is the Kronecker delta.
2.2 Composite of two Fermions and one Boson
We consider a system of two fermions (electrons or holes in cuprates) with opposite spins interacting via the exchange of one boson in a two dimensional square lattice. Physically, the boson that we have in mind in cuprate superconductors is a spinless bipolaron, since the very strong Jahn-Teller (JT) effect associated with copper ions is an important property of such cuprates [16, 15]. See also [17, 18, 19]. However, the exchanged spinless boson could be of any type, like a phonon or a spin wave, depending on the physical system and mechanism one has in mind.
[…]
$\mathfrak{h}_{f}\doteq\bigwedge\nolimits^{2}\ell^{2}(\mathbb{Z}^{2}\times\{\uparrow,\downarrow\})\subseteq\mathfrak{F}_{-}\equiv\mathfrak{F}\left(\ell^{2}(\mathbb{Z}^{2}\times\{\uparrow,\downarrow\})\right)$
[…]
$\mathfrak{F}_{+}\equiv\mathfrak{F}\left(\ell^{2}(\mathbb{Z}^{2})\right)$
associated with $\ell^{2}(\mathbb{Z}^{2})$.
We study here the effect of processes of annihilation of two fermions of opposite spins to create a boson, which can conversely be annihilated to create two new fermions. The Hilbert space associated with this composite system, made of two fermions and one boson, is the direct sum $\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$, and not to the tensor product $\mathfrak{h}_{f}\otimes\ell^{2}(\mathbb{Z}^{2})$. This fact already unveils the strong interdependence of the bosonic and fermionic parts. For this reason, from now on we rather use the term “composite of two fermions and one boson” instead of “three-body system”, in order to avoid any misinterpretation.
The fermionic part of the (infinite volume) Hamiltonian of the composite is defined to be the restriction $H_{f}\in\mathcal{B}(\mathfrak{h}_{f})$ of the formal expression
$-{\frac{\epsilon}{2}}\sum_{s\in\{\uparrow,\downarrow\},\ x,y\in\mathbb{Z}^{2}:|x-y|=1}a_{x,s}^{\ast}a_{y,s}+2\epsilon\sum\limits_{s\in\{\uparrow,\downarrow\},\,x\in\mathbb{Z}^{2}}a_{x,s}^{\ast}a_{x,s}+\mathrm{U}\sum_{x\in\mathbb{Z}^{2}}n_{x,\uparrow}n_{x,\downarrow}+\sum_{x,z\in\mathbb{Z}^{2}}\mathrm{u}\left(z\right)n_{x,\uparrow}n_{x+z,\downarrow}$ (6)
to the Hilbert space $\mathfrak{h}_{f}$. Here, $a_{x,s}$ ($a_{x,s}^{\ast}$) denotes the annihilation (creation) operator acting on the fermionic Fock space $\mathfrak{F}_{-}$ of a fermion at lattice position $x\in\mathbb{Z}^{2}$, the spin of which is $s\in\{\uparrow,\downarrow\}$. As is usual, $n_{x,s}\doteq a_{x,s}^{\ast}a_{x,s}$ stands for the number operator of fermions at lattice position $x\in\mathbb{Z}^{2}$ and spin $s\in\{\uparrow,\downarrow\}$.
The parameter $\epsilon\in\mathbb{R}_{0}^{+}$ quantifies the hopping amplitude of fermions. In high-$T_{c}$ superconductors [23, 24], $\epsilon$ is expected to be much smaller than the fermion-fermion interaction energy, more precisely the on-site repulsion strength $\mathrm{U}\in\mathbb{R}_{0}^{+}$. The function $\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$, which represents the fermion-fermion repulsion at all distances, is absolutely summable and invariant with respect to $90^{\circ}$-rotations, i.e.,
$\sum_{z\in\mathbb{Z}^{2}}\left|\mathrm{u}\left(z\right)\right|<\infty\qquad\text{and}\qquad\mathrm{u}\left(x,y\right)=\mathrm{u}\left(-y,x\right),\qquad x,y\in\mathbb{Z}\ .$ (7)
Clearly, one could set $\mathrm{U}=0$, by redefining the coupling function $\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$. It is however convenient to have a separate parameter $\mathrm{U}\in\mathbb{R}_{0}^{+}$ for the on-site repulsion, because we shall later on consider the “hard-core limit” $\mathrm{U}\rightarrow\infty$ for some fixed coupling function $\mathrm{u}$.
The above fermion-fermion interactions have been extensively studied in condensed matter physics during the last decade, in particular for two dimensional systems. For non-zero functions $\mathrm{u}$, they are named extended Hubbard interactions and they can drastically change the behavior of the system, as compared to the zero-range case (usual Hubbard interaction, $\mathrm{u}=0$). As one example, they are used in the context of ultracold atoms, ions, and molecules [11]. Its bosonic version is also experimentally investigated. See, e.g., [26] published in 2022. In theoretical studies, frequently, only nearest-neighbor interactions added to the on-site (zero-range) Hubbard interactions are considered. Here, we do not need the restriction to one-range (nearest-neighbor) interactions. We only assume that $\mathrm{u}$ is absolutely summable (see (7)), which is physically a very mild restriction, since the effective two-particle repulsive electrostatic potential in crystals is expected to decay exponentially fast in space, because of screening effects.
The rotation invariance in Equation (7) refers to the isotropy of the system under consideration. However, as shown in [22, 21], the system has low energy states that spontaneously break the isotropy. This refers to unconventional parings, typically of $d$-wave type, of electrons one experimentally observes in many high-$T_{c}$ superconductors [27, 24, 13]. In fact, to derive the existence of $d$- and $p$-wave pairings starting from a physically sound microscopic model was the aim of [22, 21]. Here, instead, we keep a broader perspective and do not study this particular question.
Similar to the fermionic part, the bosonic part of the (infinite volume) Hamiltonian of the system is defined to be the restriction $H_{b}\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}))$ to the one-boson Hilbert space $\ell^{2}(\mathbb{Z}^{2})$ of the formal expression
$\epsilon\left(-{\frac{h_{b}}{2}\sum\limits_{x,y\in\mathbb{Z}^{2}\,:\,|x-y|=1}}b_{x}^{\ast}\,b_{y}\,+2h_{b}{\sum\limits_{x\in\mathbb{Z}^{2}}}b_{x}^{\ast}\,b_{x}\right)\text{ }.$ (8)
Here, $b_{x}$ ($b_{x}^{\ast}$) denotes the annihilation (creation) operator acting on the bosonic Fock space $\mathfrak{F}_{+}$ of a boson at lattice position $x\in\mathbb{Z}^{2}$. Observe that the bosonic part only contains a kinetic term. The parameter $h_{b}\in\mathbb{R}_{0}^{+}$ quantifies the ratio of the effective masses of fermions and bosons: Taking $h_{b}$ smaller than one physically means that the bosons are heavier than the fermions. As experimentally found [28, 29, 30, 31] for cuprate superconductors, bipolarons should be much more massive than electrons or holes and, thus, in the physically relevant regime, $h_{b}$ is to be taken very small (or even zero, in an idealized situation). See [21, Section 3.1]. In the sequel, we take $h_{b}\in[0,1/2]$, meaning that the boson mass is at least as big as the mass of two fermions, as discussed in Section 3.
The term of the Hamiltonian that encodes the decay of a boson into two fermions (i.e., one of the two-electron(hole)-bipolaron-exchange interaction of the Hamiltonian) refers to the bounded operator
$W_{\mathrm{b\rightarrow f}}:\ell^{2}\left(\mathbb{Z}^{2}\right)\rightarrow\mathfrak{h}_{f}\text{ },$ (9)
which is defined to be the restriction of the formal expression
${2^{-1/2}\sum\limits_{x,y\in\mathbb{Z}^{2}}}\upsilon\left(x-y\right)c_{y}^{\ast}\,b_{x}$ (10)
to $\ell^{2}(\mathbb{Z}^{2})$, where
$c_{y}^{\ast}\doteq{\sum\limits_{z\in\mathbb{Z}^{2}}}\,\mathfrak{p}\left(z\right)\left(a_{y+z,\uparrow}^{\ast}\,a_{y,\downarrow}^{\ast}+a_{y+z,\uparrow}^{\ast}\,a_{y-z,\downarrow}^{\ast}\right)$ (11)
for some fixed non-zero function $\mathfrak{p}:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ that is invariant under $90^{\circ}$-rotations and exponentially decays in space, i.e.,
${\sum_{z\in\mathbb{Z}^{2}}}\,\mathrm{e}^{\alpha_{0}\left|z\right|}\left|\mathfrak{p}\left(z\right)\right|<\infty\qquad\text{and}\qquad\mathfrak{p}\left(x,y\right)=\mathfrak{p}\left(-y,x\right),\qquad x,y\in\mathbb{Z}\ ,$ (12)
for some $\alpha_{0}>0$. In particular the function $\mathfrak{p}$ is absolutely summable in space. For instance, in [22, 21], $\mathfrak{p}(z)=1$ for $|z|\leq 1$ and $\mathfrak{p}(z)=0$ otherwise. This is the typical example we have in mind, the point here being the fact two fermions on different lattice sites can interact by exchanging a boson. Physically, $c_{y}^{\ast}$ represents the creation of a fermion pair of zero total spin, the two components of which are slightly spread around the lattice position $y\in\mathbb{Z}^{2}$. Such pairs have finite size, because of (12). In fact, as $\mathfrak{p}\neq 0$,
$r_{\mathfrak{p}}\doteq\frac{{\sum_{z\in\mathbb{Z}^{2}}}\,\left|z\right|\left|\mathfrak{p}\left(z\right)\right|}{{\sum_{z\in\mathbb{Z}^{2}}}\,\left|\mathfrak{p}\left(z\right)\right|}\leq\inf_{\alpha_{0}>0}\alpha_{0}^{-1}\sqrt{\frac{{\sum_{z\in\mathbb{Z}^{2}}}\,\mathrm{e}^{\alpha_{0}\left|z\right|}\left|\mathfrak{p}\left(z\right)\right|}{{\sum_{z\in\mathbb{Z}^{2}}}\,\left|\mathfrak{p}\left(z\right)\right|}}<\infty\text{ }$ (13)
is naturally seen as being the actual size of such pairs. Note that the last inequality is a consequence of the Cauchy-Schwarz inequality, along with the bound
$\left|z\right|^{2}\mathrm{e}^{-\alpha_{0}\left|z\right|}\leq\alpha_{0}^{-2},\text{\qquad}\alpha_{0}>0\text{ }.$
Allowing two fermions in different lattice site to interact by exchanging a boson simply means that $r_{\mathfrak{p}}>0$. For the physical significance of this property for cuprates, see [22, 21]. Recall that the exchange strength function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ is only absolutely summable, not necessarily exponentially decaying as $\mathfrak{p}$, and invariant under $90^{\circ}$-rotations, i.e.,
${\sum_{z\in\mathbb{Z}^{2}}}\,\left|\upsilon\left(z\right)\right|<\infty\qquad\text{and}\qquad\upsilon\left(x,y\right)=\upsilon\left(-y,x\right),\qquad x,y\in\mathbb{Z}\ .$ (14)
Note that the Fourier transforms $\hat{\upsilon}$ and $\mathfrak{\hat{p}}$ of $\upsilon$ and $\mathfrak{p}$ are real-valued continuous functions (on the two-dimensional torus $\mathbb{T}^{2}$) that are again invariant under $90^{\circ}$-rotations. Additionally, $\mathfrak{\hat{p}}$ is real analytic, for $\mathfrak{p}$ is exponentially decaying. The reverse process, that is, the annihilation of two unbound fermions to form a boson, is represented by the adjoint operator
$W_{\mathrm{f\rightarrow b}}\doteq W_{\mathrm{b\rightarrow f}}^{\ast}:\mathfrak{h}_{f}\rightarrow\ell^{2}\left(\mathbb{Z}^{2}\right)\ .$ (15)
The infinite sums (6), (8) and (10) defining formally $H_{f}\in\mathcal{B}\left(\mathfrak{h}_{f}\right)$, $H_{b}\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}))$ and $W_{\mathrm{b\rightarrow f}}\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}),\mathfrak{h}_{f})$ (9) are to be understood as follows: If $\psi\in\mathfrak{h}_{f}$ or $\psi\in\ell^{2}(\mathbb{Z}^{2})$ is a finitely supported function, then the sum corresponding to $H_{f}\psi$, $H_{b}\psi$ or $W_{\mathrm{b\rightarrow f}}\psi$ is absolutely convergent. Thus, $H_{f}$, $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$ are well-defined linear operators acting on the dense subspace of such functions. One checks that $H_{f}$, $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$ are all bounded on this subspace and they thus have a unique bounded linear extension to the whole Hilbert space where they are defined, namely $\mathfrak{h}_{f}$ for $H_{f}$, and $\ell^{2}(\mathbb{Z}^{2})$ for $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$. We denote the extensions again by $H_{f}$, $H_{b}$ and $W_{\mathrm{b\rightarrow f}}$. Note that, being a bounded operator, $W_{\mathrm{b\rightarrow f}}$ has an adjoint (15), while $H_{f}$ and $H_{b}$ are clearly symmetric and so, self-adjoint, for they are also bounded.
Finally, the full Hamiltonian for the fermion-boson composite is defined, in matrix notation for the direct sum $\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$, as follows:
$\begin{pmatrix}H_{f}&W_{\mathrm{b\rightarrow f}}\\[5.0pt] W_{\mathrm{f\rightarrow b}}&H_{b}\end{pmatrix}\in\mathcal{B}\left(\mathfrak{h}_{f}\oplus\ell^{2}\left(\mathbb{Z}^{2}\right)\right)\text{ }.$ (16)
Observe that this Hamiltonian is invariant under translations, as well as $90^{\circ}$-rotations.
Using the canonical orthonormal basis [^4] I.e., for any $x,y\in\mathbb{Z}^{2}$ and $s,t\in\{\uparrow,\downarrow\}$, $\mathfrak{e}_{\left(x,s\right)}(y,s)=\delta_{s,t}\delta_{x,y}$, where $\delta_{x,y}$ is the Kronecker delta.
$\left\{\mathfrak{e}_{\left(x,s\right)}:x\in\mathbb{Z}^{2},s\in\{\uparrow,\downarrow\}\right\}\subseteq\ell^{2}\left(\mathbb{Z}^{2}\times\{\uparrow,\downarrow\}\right)$
to define the closed subspace
$\mathfrak{h}_{0}\doteq\text{\ }\overline{\mathrm{span}}\left\{\mathfrak{e}_{(x,\uparrow)}\wedge\mathfrak{e}_{(y,\downarrow)}\,:\,x,y\in\mathbb{Z}^{2}\right\}\subseteq\mathfrak{h}_{f}\ ,$ (17)
we remark that the zero-spin subspace
$\mathfrak{H}\doteq\mathfrak{h}_{0}\oplus\ell^{2}\left(\mathbb{Z}^{2}\right)\subseteq\mathfrak{h}_{f}\oplus\ell^{2}\left(\mathbb{Z}^{2}\right)$ (18)
is invariant under the action of the (full) Hamiltonian (16). We can thus consider its restriction
$H\doteq\left.\begin{pmatrix}H_{f}&W_{\mathrm{b\rightarrow f}}\\[5.0pt] W_{\mathrm{f\rightarrow b}}&H_{b}\end{pmatrix}\right|_{\mathfrak{H}}\in\mathcal{B}\left(\mathfrak{H}\right)$ (19)
to this particular subspace $\mathfrak{H}\subseteq\mathfrak{h}_{f}\oplus\ell^{2}(\mathbb{Z}^{2})$. In fact, as the boson is assumed to be spinless, by the conservation of angular momentum, we have that the total spin of the fermion pair resulting from a bosonic decay must be zero. In other words, the physically relevant (vector) states of the fermion-boson compound system always lie in $\mathfrak{H}$. Note finally that $H$ inherits the symmetries of the Hamiltonian (16), i.e., $H$ is invariant under translations and $90^{\circ}$-rotations. Note that this last symmetry, i.e., the rotation invariance, is mainly relevant for the study of unconventional pairings, which is not done here.
2.3 The Model in Spaces of Quasi-Momenta
We have a composite of two fermions and one boson whose Hamiltonian is translation invariant. In this case, it is a standard procedure (see, e.g. [32, Chapter XIII.16 ]) to use the direct integral decomposition of the Hamiltonian in Fourier space in order to study its spectral properties.
For the two dimensional lattice $\mathbb{Z}^{2}$, the (Fourier) space of quasi-momenta is nothing else than the torus
$\mathbb{T}^{2}\doteq[-\pi,\pi)^{2}\subseteq\mathbb{R}^{2}\ .$
This set is endowed with the metric $d_{\mathbb{T}^{2}}$ defined by
$d_{\mathbb{T}^{2}}\left(k,p\right)\doteq\min\left\{\left|k-p-q\right|:q\in 2\pi\mathbb{Z}^{2}\right\}\ ,$ (20)
where $|k-p-q|$ is the Euclidean distance between $k$ and $p+q$ in $\mathbb{R}^{2}$. This defines a compact metric space $(\mathbb{T}^{2},d_{\mathbb{T}^{2}})$. Observe also that the usual group operation in $\mathbb{T}^{2}$, i.e., the sum in $\mathbb{R}^{2}$ modulo $(2\pi,2\pi)$, is a continuous operation, while any Borel set in $\mathbb{T}^{2}$ is also a Borel set in $\mathbb{R}^{2}$ (endowed with the Euclidean metric). We also need the normalized Haar measure $\nu$ on $\mathbb{T}^{2}$ defined for any Borel set $B\subseteq\mathbb{T}^{2}$ by
$\nu(B)=(2\pi)^{-2}\mathbf{\lambda}(B)\ ,$ (21)
where $\mathbf{\lambda}$ is the Lebesgue measure in $\mathbb{R}^{2}$. This measure appears in relation with direct integrals of constant Hilbert spaces on the two-dimensional torus $\mathbb{T}^{2}$, like the Hilbert space
$L^{2}\left(\mathbb{T}^{2}\right)\equiv L^{2}\left(\mathbb{T}^{2},\mathbb{C}\right)\equiv L^{2}\left(\mathbb{T}^{2},\mathbb{C},\nu\right)\doteq{\int_{\mathbb{T}^{2}}^{\oplus}}\mathbb{C}\,\nu(\mathrm{d}k)$
of square-integrable, complex-valued functions on $\mathbb{T}^{2}$. Since the Haar measure $\nu$ is used in all our direct integrals on $\mathbb{T}^{2}$, for simplicity we often remove the symbol $\nu$ from the notation of $L^{2}$-spaces, unless this information is important to recall.
The Fourier transform can be applied in the fermionic and bosonic sectors. In the fermionic one, there is more than one natural way of implementing the transform, as the corresponding functions have two arguments in $\mathbb{Z}^{2}$. It turns out that to be very useful to extract the total quasi-momentum of fermionic pairs. In fact, we consider the direct integral
$L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)\equiv L^{2}\left(\mathbb{T}^{2},\mathcal{H},\nu\right)\doteq{\int_{\mathbb{T}^{2}}^{\oplus}}L^{2}\left(\mathbb{T}^{2},\mathbb{C},\nu\right)\oplus\mathbb{C}\,\nu(\mathrm{d}k)$ (22)
of the (constant fiber) Hilbert space
$\mathcal{H}\doteq L^{2}\left(\mathbb{T}^{2}\right)\oplus\mathbb{C}\equiv L^{2}\left(\mathbb{T}^{2},\mathbb{C},\nu\right)\oplus\mathbb{C}$ (23)
over the torus $\mathbb{T}^{2}$, and choose a unitary transformation
$U:\mathfrak{H}\longrightarrow L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$ (24)
in such a way that $k\in\mathbb{T}^{2}$, the fiber quasi-momentum, is exactly the total quasi-momentum of the fermion pair. Recall that $\mathfrak{H}$ (18) is Hilbert space on which $H$ is originally defined. More precisely,
$U\doteq U_{f}\oplus\mathcal{F}\ ,$
where
$\mathcal{F}:\ell^{2}\left(\mathbb{Z}^{2}\right)\rightarrow L^{2}\left(\mathbb{T}^{2}\right)$
is the Fourier transform on $\ell^{2}(\mathbb{Z}^{2})$, while the fermionic part
$U_{f}\doteq U_{2}U_{1}:\mathfrak{h}_{0}\rightarrow\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\nu\left(\mathrm{d}k\right)$ (25)
is the composition of two unitary (linear) transformations $U_{1}$ and $U_{2}$, whose exact definitions are given in Section 5.1.
Here, the Fourier transform is defined for any $f\in\ell^{2}(\mathbb{Z}^{2})$ by
$\hat{f}\left(k\right)\equiv\mathcal{F}f\left(k\right)={\sum_{x\in\mathbb{Z}^{2}}}\,\mathrm{e}^{ik\cdot x}f\left(x\right)\ ,\qquad k\in\mathbb{T}^{2}\ ,$ (26)
$k\cdot x$ being the usual scalar product of $k\in\mathbb{T}^{2}$ and $x\in\mathbb{Z}^{2}$, seen as vectors of $\mathbb{R}^{2}$. We often use the symbol $\widehat{\left(\cdot\right)}$ to shorten the notation of the Fourier transform. For instance, for any $x\in\mathbb{Z}^{2}$, we write $\hat{\mathfrak{e}}_{x}$ to denote the function $\mathrm{e}^{i(\cdot)\cdot x}$ on the torus $\mathbb{T}^{2}$. That is, $\{\hat{\mathfrak{e}}_{x}\}_{x\in\mathbb{Z}^{2}}$ is the image under the Fourier transform of the canonical orthonormal basis$\{\mathfrak{e}_{x}\}_{x\in\mathbb{Z}^{2}}$ (5) of $\ell^{2}(\mathbb{Z}^{2})$.
For the reader’s convenience and completeness, in Section 5.3 we gather key results from the theory of direct integrals with constant fiber Hilbert spaces. In the next subsection, we explain how the properties of the Hamiltonian $H\in\mathcal{B}(\mathfrak{H})$ defined by (19) can be studied on the direct integral (22) over total quasi-momenta.
2.4 Fiber Decomposition of the Hamiltonian
By explicit computations, exactly like in [22, 21], we show that the conjugation of the Hamiltonian $H\in\mathcal{B}(\mathfrak{H})$ with the unitary transformation $U$ of Equation (24) is a decomposable operator on the direct integral $L^{2}(\mathbb{T}^{2},\mathcal{H})$. To state this result precisely, we need preliminary definitions allowing to define the so-called “fiber Hamiltonians”, or “fibers” for short, $A(k)\in\mathcal{B}(\mathcal{H})$ of $UHU^{\ast}$ at total quasi-momenta $k\in\mathbb{T}^{2}$. In fact, the mapping $k\mapsto A(k)$ defines an element of the von Neumann algebra
$L^{\infty}\left(\mathbb{T}^{2},\mathcal{B}(\mathcal{H})\right)\equiv L^{\infty}\left(\mathbb{T}^{2},\mathcal{B}(\mathcal{H}),\nu\right)$
of (equivalence classes of) strongly measurable functions $\mathbb{T}^{2}\rightarrow\mathcal{B}(\mathcal{H})$. Recall that the (unique) norm of this von Neumann algebra (as a $C^{\ast}$-algebra) is the essential supremum (with respect to the measure $\nu$ on the torus):
$\left\|B\right\|_{\infty}\doteq\underset{k\in\mathbb{T}^{2}}{\mathrm{ess-sup}}\,\left\|B\left(k\right)\right\|_{\mathrm{op}}<\infty\text{ },$
where $\left\|\cdot\right\|_{\mathrm{op}}$ stands for the operator norm. In fact, $A(\cdot)$ is even a continuous function on the torus $\mathbb{T}^{2}$ as compact metric space. See Section 5.3 for more details.
Given a total quasi-momentum $k\in\mathbb{T}^{2}$ and the parameters $\epsilon,h_{b}\in\mathbb{R}_{0}^{+}$ tuning the strengths of the two (fermionic and bosonic) kinetic parts of the model, we define continuous, real-valued functions $\mathfrak{f}(k),\mathfrak{d}(k),\mathfrak{b}\in C\left(\mathbb{T}^{2}\right)$ on the torus $\mathbb{T}^{2}$ by
$\mathfrak{b}\left(p\right)$ $\doteq$ $h_{b}\epsilon\left(2-\cos\left(p\right)\right)\ ,$ (27)
$\mathfrak{f}\left(k\right)\left(p\right)$ $\doteq$ $\epsilon\left\{4-\cos\left(p+k\right)-\cos\left(p\right)\right\}\ ,$ (28)
$\mathfrak{d}\left(k\right)\left(p\right)$ $\doteq$ $\mathfrak{\hat{p}}\left(k+p\right)\ ,$ (29)
for all $p=(p_{1},p_{2})\in\mathbb{T}^{2}$, where
$\cos\left(q\right)\doteq\cos\left(q_{1}\right)+\cos\left(q_{2}\right)\text{ },\qquad q=(q_{1},q_{2})\in\mathbb{R}^{2}\ .$ (30)
Recall that (the $(2\pi,2\pi)$-periodic function) $\mathfrak{\hat{p}}$ is the Fourier transform of $\mathfrak{p}$, which is the function defining the operator $c_{y}^{\ast}$ in (11), representing the creation of fermion pairs in the model.
Then, at any quasi-momentum $k\in\mathbb{T}^{2}$ and on-site repulsion strength $\mathrm{U}\in\mathbb{R}_{0}^{+}$, we define the bounded operators $B_{1,1}\left(k\right)$ and $A_{1,1}(\mathrm{U},k)$ acting on the Hilbert space $L^{2}\left(\mathbb{T}^{2}\right)$ by
$B_{1,1}\left(k\right)$ $\doteq$ $M_{\mathfrak{f}\left(k\right)}+{\sum\limits_{x\in\mathbb{Z}^{2}}}\,\mathrm{u}\left(x\right)P_{x}\ ,$ (31)
$A_{1,1}\left(\mathrm{U},k\right)$ $\doteq$ $B_{1,1}\left(k\right)+\mathrm{U}P_{0}\ ,$ (32)
where $M_{\mathfrak{f}(k)}$ stands for the multiplication operator by $\mathfrak{f}(k)\in C(\mathbb{T}^{2})$ and $P_{x}$ is the orthogonal projection onto the one-dimensional subspace $\mathbb{C}\mathfrak{\hat{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. Note that the infinite sum defining the bounded operator $B_{1,1}(k)$ is absolutely convergent, for the function $\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ is, by assumption, absolutely summable. See (7). We define next
$\begin{array}[]{cccl}A_{2,1}\left(k\right):&L^{2}\left(\mathbb{T}^{2}\right)&\rightarrow&\mathbb{C}\\ &\varphi&\mapsto&\hat{\upsilon}\left(k\right)\left\langle\mathfrak{d}\left(k\right),\varphi\right\rangle\ ,\end{array}$ (35)
$\begin{array}[]{cccl}A_{1,2}\left(k\right):&\mathbb{C}&\rightarrow&L^{2}\left(\mathbb{T}^{2}\right)\\ &z&\mapsto&\hat{\upsilon}\left(k\right)\mathfrak{d}\left(k\right)z\end{array}$ (38)
as well as
$\begin{array}[]{cccl}A_{2,2}\left(k\right):&\mathbb{C}&\rightarrow&\mathbb{C}\\ &z&\mapsto&\mathfrak{b}\left(k\right)z\end{array}$ (39)
for any fixed $k\in\mathbb{T}^{2}$. By compactness of $\mathbb{T}^{2}$ and continuity (in operator norm) of the mappings $k\mapsto A_{i,j}(k)$ for all $i,j\in\{1,2\}$, we have
$A\left(\cdot\right)\equiv A\left(\mathrm{U},\cdot\right)\doteq\begin{pmatrix}A_{1,1}\left(\mathrm{U},\cdot\right)&A_{1,2}\left(\cdot\right)\\[5.0pt] A_{2,1}\left(\cdot\right)&A_{2,2}\left(\cdot\right)\end{pmatrix}\in L^{\infty}\left(\mathbb{T}^{2},\mathcal{H}\right)\ ,$ (40)
which is meanwhile the fiber decomposition of the operator $UHU^{\ast}$:
Proposition 2.1 (Fiber decomposition of the quantum model)
The conjugation of $H$ by $U$ (24) is decomposable and has $A(\cdot)$ as its fibers, that is,
$UHU^{\ast}={\int_{\mathbb{T}^{2}}^{\oplus}}A(k)\,\nu(\mathrm{d}k)\ .$
Proof. This is proven from explicit computations which are almost the same as those done in [22, 21]. We omit the details.
The fiber decomposition given by Proposition 2.1 is useful because it gives access to spectral properties of $H$. In fact, for an operator that is decomposable on $L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$, i.e., an operator unitarily equivalent to an element of the von Neumann algebra $L^{2}\left(\mathbb{T}^{2},\mathcal{B}(\mathcal{H})\right)$, like the Hamiltonian $H$, the fibers $A(k)$ of which are all self-adjoint, it is known that $\lambda\in\sigma(H)$ if, and only if, for all $\varepsilon>0$,
$\nu\left(\left\{k\in\mathbb{T}^{2}:\sigma\left(A\left(k\right)\right)\cap\left(\lambda-\varepsilon,\lambda+\varepsilon\right)\neq\emptyset\right\}\right)>0\text{ }.$
See Theorem 5.3. As is usual, here, $\sigma(X)$ denotes the spectrum of any operator $X$ acting on some Hilbert space.
3 Main Results
In this section we state our main results, starting with general spectral properties of the Hamiltonian $H$ to finish with results related with scattering.
Recall that the model has parameters $\epsilon,\mathrm{U},h_{b}\in\mathbb{R}_{0}^{+}$ and $\alpha_{0}\in\mathbb{R}^{+}$, and it depends on the choice of functions
$\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}\ ,\quad\mathfrak{p}:\mathbb{Z}^{2}\rightarrow\mathbb{R}\ ,\quad\text{and}\quad\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$
that are absolutely summable and invariant with respect to $90^{\circ}$-rotations. Observe additionally that the function $\mathfrak{p}$ is required to be exponentially decaying, that is, $\mathrm{e}^{\alpha_{0}|\cdot|}\mathfrak{p}$ is absolutely summable for some $\alpha_{0}>0$. See Equations (7), (12) and (14). All details of the Hamiltonian, like the precise choice of its parameters and functions, are not explicitly mentioned in our discussions or statements below, unless it is important for clearness. There is however one important condition to clarify:
While some of our results can be obtained without any other restriction, frequently we fix the parameter $h_{b}$ in the interval $[0,1/2]$. This choice physically means that the boson is heavier than two fermions. As already discussed above, the assumption is perfectly justified when one views the two fermions and the boson of the model as being electrons or holes and a bipolaron, respectively, in a cuprate. In fact, polarons (and thus bipolarons) are charge carriers that are self-trapped inside a strong and local lattice deformation that surrounds them, caused by electrostatic interactions between the carriers and the lattice. A priori, such (strong and local) lattice deformations can barely move, that is, their effective mass is huge. See, e.g., [28, 29, 30]. This is coherent with the assumption of a large mass of JT bipolarons in copper oxides [31], similar to JT polarons [33].
We show that the condition $h_{b}\in[0,1/2]$ is crucial to obtain dressed bound fermion pairs, which are expected to represent the charge carriers below the pseudogap temperature [21].
3.1 Spectral Properties
Having in mind Proposition 2.1 and Theorem 5.3 we start with the spectral properties of fiber Hamiltonians (40) at any quasi-momentum $k\in\mathbb{T}^{2}$. This refers to the following theorem:
Theorem 3.1 (Spectral properties of fiber Hamiltonians)
Fix $\epsilon,\mathrm{U}\in\mathbb{R}_{0}^{+}$, $h_{b}\in[0,1/2]$ and $k\in\mathbb{T}^{2}$.
  1. i.)
    Essential spectrum $\sigma_{\mathrm{ess}}\left(\cdot\right)$ of the fiber Hamiltonian:
    $\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},k\right)\right)=\,\mathfrak{f}\left(k\right)\left(\mathbb{T}^{2}\right)=2\epsilon\cos\left(k/2\right)\left[-1,1\right]+4\epsilon\ .$
  2. ii.)
    Ground state energy: There is a unique, non-degenerate eigenvalue $\mathrm{E}(\mathrm{U},k)\leq\mathfrak{b}(k)$ of $A(\mathrm{U},k)$ below the essential spectrum, with associated eigenvector
    $\Psi(\mathrm{U},k)\doteq(\hat{\psi}_{k}\left(\mathrm{U}\right),-1)\ ,\text{\quad where\quad}\hat{\psi}_{\mathrm{U},k}\doteq\hat{\upsilon}\left(k\right)\left(A_{1,1}\left(\mathrm{U},k\right)-\mathrm{E}\left(\mathrm{U},k\right)\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\in L^{2}\left(\mathbb{T}^{2}\right)\ .$
    In addition, $\mathrm{E}(\mathrm{U},k)=\mathfrak{b}(k)$ iff $\hat{\upsilon}(k)=0$. Recall that $\mathfrak{b}(k)$ is defined by Equation (27).
  3. iii.)
    Spectral gap and Anderson localization: If $\hat{\upsilon}(0)\neq 0$ and $r_{\mathfrak{p}}>0$ then
    $\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\left\{\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},k\right)\right)-\mathrm{E}\left(\mathrm{U},k\right)\right\}>0$
    and there are $C,\alpha\in\mathbb{R}^{+}$ such that, for all $k\in\mathbb{T}^{2}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$,
    $\left|\mathcal{F}^{-1}[\hat{\psi}_{\mathrm{U},k}]\left(x\right)\right|\leq C\mathrm{e}^{-\alpha|x|}\ ,\qquad x\in\mathbb{Z}^{2}\ .$
  4. iv.)
    $\mathrm{E}\left(\mathrm{U},\cdot\right):\mathbb{T}^{2}\rightarrow\mathbb{R}$ is a continuous function and if $\hat{\upsilon}$ is of class [^5] Given $n,d\in\mathbb{N}$ and an open set $\Omega\subseteq\mathbb{R}^{n}$, $C^{d}(\Omega)$ denotes the set of $d$ times continuously differentiable, complex-valued functions on $\Omega$, while $C^{\omega}(\Omega)$ and $C^{a}(\Omega)$ refer to the space of smooth and real analytic functions on $\Omega$, respectively. $C^{d}$ on $(-\pi,\pi)^{2}\backslash\{0\}\subseteq\mathbb{R}^{2}$ with $d\in\mathbb{N}\cup\{\omega,a\}$ then so does $\mathrm{E}\left(\mathrm{U},\cdot\right)$ on $(-\pi,\pi)^{2}\backslash\{0\}$.
Proof. The theorem is a combination of Theorems 4.7, 4.8, 4.17 and 4.19 together with Propositions 4.1, 4.18 and Corollary 4.5 (see (118)).
Remark 3.2
Recall that if, for some natural number $d\geq 1$,
$\sum_{x\in\mathbb{Z}^{2}}\left|x\right|^{d}\left|\upsilon\left(x\right)\right|<\infty$
then the Fourier transform $\hat{\upsilon}$ of the function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$, as defined by (26), is of class $C^{d}$ on the whole torus $\mathbb{T}^{2}$.
Remark 3.3
If $\mathfrak{p}\in\mathbb{C}\mathfrak{e}_{0}$, i.e., $r_{\mathfrak{p}}=0$, then Theorem 3.1 (iii) remains true, but not uniformly in $\mathrm{U}\in\mathbb{R}_{0}^{+}$. That is, in this case, one only has:
$\min_{k\in\mathbb{T}^{2}}\left\{\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},k\right)\right)-\mathrm{E}\left(\mathrm{U},k\right)\right\}>0$
and there are $C_{\mathrm{U}},\alpha_{\mathrm{U}}\in\mathbb{R}^{+}$ such that, for all $k\in\mathbb{T}^{2}$,
$\left|\mathcal{F}^{-1}[\hat{\psi}_{\mathrm{U},k}]\left(x\right)\right|\leq C_{\mathrm{U}}\mathrm{e}^{-\alpha_{\mathrm{U}}|x|}\ ,\qquad x\in\mathbb{Z}^{2}\ .$
Assertion (i) of Theorem 3.1 holds true for all $h_{b}\in\mathbb{R}_{0}^{+}$, but the other assertions need the restriction $h_{b}\in[0,1/2]$ to ensure that the eigenvalues are below the essential spectrum, as stated in Assertion (ii). In fact, $h_{b}\in[0,1/2]$ iff
$\mathfrak{b}\left(k\right)\leq\mathfrak{z}\left(k\right)\doteq 4\epsilon-2\epsilon\cos(k/2)=\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},k\right)\right)$ (41)
for all $k\in\mathbb{T}^{2}$, with equality only at $k=0$. See Equation (27). Therefore, by Assertion (ii), $\mathrm{E}(\mathrm{U},k)$ belongs to the essential spectrum iff $k=0$ and $\hat{\upsilon}(0)=0$. Otherwise, we have a uniform spectral gap, as stated in Assertion (iii).
As is explained in [22, 21], the eigenvalue $\mathrm{E}\left(\mathrm{U},k\right)$ given by Theorem 3.1 is associated with the formation of dressed bound fermion pairs with total quasi-momentum $k\in\mathbb{T}^{2}$. These pairs are generally exponentially localized, thanks to Theorem 3.1 (iii), which basically implies that the two fermions move together confined within some small ball, that is, they are tightly bound in space, provided $\hat{\upsilon}(0)\neq 0$. When $\mathfrak{p}(z)\neq 0$ for some $z\neq 0,$ or, equivalently, $r_{\mathfrak{p}}>0$, the size of the small does not depend upon the Hubbard coupling constant $\mathrm{U}$ and a very large $\mathrm{U}\gg 1$ only prevents two fermions from occupying the same lattice site. The condition $\mathfrak{p}(z)\neq 0$ for some $z\neq 0$ or, equivalently, $r_{\mathfrak{p}}>0$, is therefore pivotal to get (Cooper) fermion pairs, the natural candidates for superconducting charge carriers, in presence of strong on-site Coulomb repulsions, like in cuprates [23, 24].
For the usual (i.e., non-extended) Hubbard interaction ($\mathrm{u}=0$) and one-range [^6] It means here that $c_{y}^{\ast}\doteq\sum_{z\in\mathbb{Z}^{2}:\left|z\right|\leq 1}\,\mathfrak{p}(z)(a_{y+z,\uparrow}^{\ast}\,a_{y,\downarrow}^{\ast}+a_{y+z,\uparrow}^{\ast}\,a_{y-z,\downarrow}^{\ast})$. creation operators $c_{y}^{\ast}$ of fermion pairs (in this case (12) holds true for all $\alpha_{0}\in\mathbb{R}_{0}^{+}$), note that a weak form of pair localization was previously shown in the ground state. See for instance [22, Theorem 3 and Proposition 13]. In this particular case, estimates of $\mathrm{E}\left(\mathrm{U},k\right)$ and $\Psi(\mathrm{U},k)$ are known for large $\mathrm{U}\gg 1$. See, e.g., [22, Theorem 4, Corollary 5, Theorem 16]. Recall that the aim in [22, 21] was to show the existence of $d$- and $p$-wave pairings in the ground state for some physically sound model, and not the systematic study of a general class of models. In [21] we conjecture that such dressed bound fermion pairs represent the charge carriers below the pseudogap temperature in cuprates.
Theorem 3.1 combined with Proposition 2.1 and the theory of direct integrals (cf. Theorem 5.3) has direct consequences for the spectrum of the full Hamiltonian $H\in\mathcal{B}(\mathfrak{H})$, which is defined by Equation (19). Among other things we obtain the following corollary:
Corollary 3.4 (Spectral properties of $H$)
Fix $\epsilon,\mathrm{U}\in\mathbb{R}_{0}^{+}$ and $h_{b}\in[0,1/2]$. Then,
$\left\{\mathrm{E}\left(\mathrm{U},k\right):k\in\mathbb{T}^{2}\right\}\cup\left(0,8\epsilon\right)\subseteq\sigma\left(H\right)\ ,$
where $\sigma\left(H\right)$ denotes, as is usual, the spectrum of $H$, and
$\min\sigma\left(H\right)=E\left(\mathrm{U}\right)\doteq\min_{k\in\mathbb{T}^{2}}\mathrm{E}\left(\mathrm{U},k\right)\leq 0\ .$
If additionally $\hat{\upsilon}(0)\neq 0$ and $r_{\mathfrak{p}}>0$, then
$\sup_{\mathrm{U}\in\mathbb{R}_{0}^{+}}E\left(\mathrm{U}\right)<0\ .$
Proof. To prove the first assertion, it suffices to combine Proposition 2.1 and Theorem 3.1 with Theorem 5.3. The second one can be proven like in [22] by using Kato’s perturbation theory [34]. In Proposition 4.10 we give an alternative and more direct proof of it. Finally, the last assertion is a consequence of the inequalities
$\min_{k\in\mathbb{T}^{2}}\mathrm{E}\left(\mathrm{U},k\right)\leq\mathrm{E}\left(\mathrm{U},0\right)=\mathrm{E}\left(\mathrm{U},0\right)-\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},0\right)\right)$
and Theorem 3.1 (iii).
Physically speaking, the spectral values of $H$ represent the energy levels that are available to the composite of two fermions and one boson, in particular for a fermion pair exchanging a boson. As expected, the minimum energy $E$, also well-known as the ground state energy, is given by minimizing the eigenvalues $\mathrm{E}(\mathrm{U},k)$ over the torus $\mathbb{T}^{2}$.
We now study the model at very large on-site repulsion $\mathrm{U}\gg 1$. In fact, quoting [21], “in all cuprates, there is undeniable experimental evidence of strong on-site Coulomb repulsions, leading to the universally observed Mott transition at zero doping [23, 24]. This phase is characterized by a periodic distribution of fermions (electrons or holes) with exactly one particle per lattice site. Doping copper oxides with holes or electrons can prevent this situation. Instead, at sufficiently small temperatures a superconducting phase is achieved, as first discovered in 1986 for the copper oxide perovskite $\mathrm{La}_{2-x}\mathrm{Ba}_{x}\mathrm{CuO}_{4}$ [16].” However, instead of the usual $s$-wave superconductivity, one experimentally observes $d$-wave superconductivity [27, 24, 13]. The fact that only the $s$-wave pairing is suppressed also advocates for a very local (i.e., on-site) and strong effective repulsion of fermions. For this reason we consider the limit $\mathrm{U}\rightarrow\infty$ in our model. It corresponds to a hard core limit because it prevents two fermions from being on the same lattice site.
In the limit $\mathrm{U}\rightarrow\infty$, it is easy to see that the ground state energy $E(\mathrm{U})$ of Corollary 3.4 defines an increasing function of $\mathrm{U}\in\mathbb{R}_{0}^{+}$, which is bounded from above by 00. Hence,
$E\left(\infty\right)\doteq\lim\limits_{\mathrm{U}\rightarrow\infty}E\left(\mathrm{U}\right)={\sup\limits_{\mathrm{U}\in\mathbb{R}_{0}^{+}}}E\left(\mathrm{U}\right)\leq 0\ .$ (42)
For more details, see Lemma 4.13. The limit $\mathrm{U}\rightarrow\infty$ of the eigenvalue and eigenvector of each fiber, given by Theorem 3.1, is less trivial to obtain and is the object of the next theorem:
Theorem 3.5 (Spectral properties of fiber Hamiltonians – Hard-core limit)
Fix $\epsilon,\mathrm{U}\in\mathbb{R}_{0}^{+}$ and $h_{b}\in[0,1/2]$. The following limits exist:
$\mathrm{E}\left(\infty,k\right)$ $\doteq$ ${\lim\limits_{\mathrm{U}\rightarrow\infty}}\mathrm{E}\left(\mathrm{U},k\right)={\sup\limits_{\mathrm{U}\in\mathbb{R}_{0}^{+}}}\mathrm{E}\left(\mathrm{U},k\right)\leq\mathfrak{b}\left(k\right)\ ,\qquad k\in\mathbb{T}^{2}\ .$
$\Psi\left(\infty,k\right)$ $\doteq$ ${\lim_{\mathrm{U}\rightarrow\infty}}\,\Psi\left(\mathrm{U},k\right)\in\mathcal{H}\backslash\{0\}\ ,\qquad k\in\mathbb{T}^{2}\backslash\{0\}\ .$
Assertion (iv) of Theorem 3.1 also holds true for $\mathrm{U}=\infty$. In addition, when $r_{\mathfrak{p}}>0$, $\mathrm{E}(\infty,k)=\mathfrak{b}(k)$ iff $\hat{\upsilon}(k)=0$. If $r_{\mathfrak{p}}>0$ and $\hat{\upsilon}(0)\neq 0$, then $\Psi(\infty,0)$ exists.
Proof. See Theorems 4.14 and 4.17.
Note that the eigenvalues given by Theorem 3.1 are not explicitly known. The same is of course true in the hard-core limit $\mathrm{U}\rightarrow\infty$. For applications it is important to have a sufficiently good control on these objects to be able to compute them, either analytically or numerically. This is done in [22, 21] for the special case of the usual Hubbard interaction ($\mathrm{u}=0$) and one-range creation operators $c_{y}^{\ast}$ of fermion pairs, by providing estimates for $\mathrm{E}\left(\mathrm{U},k\right)$ and $\Psi\left(\mathrm{U},k\right)$ at large $\mathrm{U}$. See [22, Theorem 4, Corollary 5, Theorem 16].
Recall that $\hat{\upsilon}$ is the Fourier transform of $\upsilon$, which is the function appearing in Equation (10), encoding the (exchange) interaction between fermion pairs and bosons. By Theorem 3.1, if $\hat{\upsilon}(k)=0$ then $\mathrm{E}(\mathrm{U},k)$ is nothing else than the explicit function $\mathfrak{b}(k)$ (27). Hence, we focus on the physically more relevant case $\hat{\upsilon}(k)\neq 0$. Using the Birman-Schwinger principle (Theorem 5.10), we show in this case that the eigenvalue $\mathrm{E}(\mathrm{U},k)$ is the unique solution to a relatively simple equation for real numbers, similar to the characteristic equation used to compute eigenvalues of matrices.
To this end, we define a function $\mathfrak{T}:\mathcal{D}\rightarrow\mathbb{R}$ on the set
$\mathcal{D}\doteq\left\{\left(\mathrm{U},k,x\right)\in\left[0,\infty\right]\times\mathbb{T}^{2}\times\mathbb{R}:x<\mathfrak{z}\left(k\right)\right\}\subseteq\mathbb{R}^{3}$
by
$\mathfrak{T}\left(\mathrm{U},k,x\right)\doteq\left\langle\mathfrak{d}\left(k\right),\left(A_{1,1}\left(\mathrm{U},k\right)-x\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right\rangle$ (43)
for any finite $\mathrm{U}\in\mathbb{R}_{0}^{+}$, $k\in\mathbb{T}^{2}$ and $x\in\left(0,\mathfrak{z}\left(k\right)\right)$, while for the infinite on-site repulsion, $k\in\mathbb{T}^{2}$ and $x\in\left(0,\mathfrak{z}\left(k\right)\right)$,
$\mathfrak{T}\left(\infty,k,x\right)\doteq\lim_{\mathrm{U}\rightarrow\infty}\mathfrak{T}\left(\mathrm{U},k,x\right)\text{ },$
the above limit existing by virtue of Corollary 4.12. Recall that $\mathfrak{z}\left(k\right)$ is defined by Equation (63). In fact, for any $k\in\mathbb{T}^{2}$ and $x\in\left(0,\mathfrak{z}\left(k\right)\right)$,
$\mathfrak{T}\left(\infty,k,x\right)=R_{\mathfrak{s},\mathfrak{s}}^{-1}\left({R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}-}\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}\right)>0\ ,$
where $R_{\mathfrak{s},\mathfrak{s}}$, $R_{\mathfrak{s},\mathfrak{d}}$, $R_{\mathfrak{d},\mathfrak{s}}$ and $R_{\mathfrak{d},\mathfrak{d}}$ are four constants defined by Equations (87)–(90) with $\lambda=x$. When $\mathrm{u}=0$ these constants are given by explicit integrals on the torus $\mathbb{T}^{2}$ [22, 21]. Then, the eigenvalues, the existence of which is stated in Theorem 3.1, as well as their limits (Theorem 3.5), can be studied via the following characteristic equation:
Theorem 3.6 (Characteristic equation for the fiber ground states)
Fix $\epsilon\in\mathbb{R}_{0}^{+}$, $h_{b}\in[0,1/2]$ and $k\in\mathbb{T}^{2}$ such that $\hat{\upsilon}(k)\neq 0$. Then, for any $\mathrm{U}\in\left[0,\infty\right]$, $\mathrm{E}(\mathrm{U},k)$ is the unique solution to the equation
$\hat{\upsilon}\left(k\right)^{2}\mathfrak{T}\left(\mathrm{U},k,x\right)+x-\mathfrak{b}\left(k\right)=0\text{ },\qquad x<\mathfrak{z}\left(k\right)\ .$
Proof. For $\mathrm{U}\in\mathbb{R}_{0}^{+}$, combine Theorem 4.4 with Theorem 4.7, while for $\mathrm{U}=\infty$ use Theorem 4.14.
Notice that, more generally, for any fixed $\mathrm{U}\in\mathbb{R}_{0}^{+}$, the same characteristic equation determines all eigenvalues of the fiber lying in the resolvent set $\rho(A_{1,1}(\mathrm{U},k))$ of the operator $A_{1,1}(\mathrm{U},k)$. See Theorem 4.4. Also the associated eigenspaces can be explicitly characterized, thanks to Corollary 4.5. In this context, Corollary 4.6 shows that, for any $h_{b}\in[0,1/2]$ and total quasi-momentum $k\in\mathbb{T}^{2}$, there is at most one eigenvalue of $A(\mathrm{U},k)$ in each connected component of $\rho(A_{1,1}(\mathrm{U},k))\cap\mathbb{R}$.
3.2 Dispersion Relation of Dressed Bound Fermion Pairs
By Theorem 3.1, $\mathrm{E}:\mathbb{T}^{2}\rightarrow\mathbb{R}$ is a continuous family of non-degenerate eigenvalues, generally (at least for $k\neq 0$) associated with exponentially localized eigenvectors. Note that the case $k=0$ is particular when $\hat{\upsilon}(0)=0$, since $\mathrm{E}(0)$ is not an isolated eigenvalue of $A(\mathrm{U},0)$. However, the family $(\mathrm{E}(k))_{k\in\mathbb{T}^{2}}$ is still continuous. The peculiar behavior at $k=0$ leads us to only consider total quasi-momenta in the subset
$\mathbb{S}^{2}\doteq\left(-\pi,\pi\right)^{2}\backslash\left\{0\right\}\subseteq\mathbb{T}^{2}\ ,$ (44)
as, for instance, in Theorem 3.1 (iv).
Because of Proposition 2.1, the family $(\mathrm{E}(k))_{k\in\mathbb{T}^{2}}$ can thus be seen as the effective dispersion relation of dressed bound fermion pairs. It is expected to determine transport properties of the quantum system at low temperatures. We now define in mathematical terms what a dispersion relation is.
[…]
Keeping in mind that all our quantities are parametrized by the on-site repulsion $\mathrm{U}\in\left[0,\infty\right]$, we define a family of dispersion relations associated with the fiber Hamiltonians $A(\mathrm{U},k)$ as follows:
Definition 3.7 (Family of dispersion relations)
A function $\varkappa:\left[0,\infty\right]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ is said to be a family of dispersion relations $\varkappa(\mathrm{U},\cdot)$ if the following properties are satisfied for all $\mathrm{U}\in\left[0,\infty\right]$:
  1. i.)
    For any $k\in\mathbb{T}^{2}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$, $\varkappa(\mathrm{U},k)$ is an eigenvalue of $A(\mathrm{U},k)$ and
    $\varkappa\left(\infty,k\right)=\lim\limits_{\mathrm{U}\rightarrow\infty}\varkappa\left(\mathrm{U},k\right)\ .$
  2. ii.)
    For all $\mathrm{U}\in\left[0,\infty\right]$, $\varkappa(\mathrm{U},\cdot)\in C(\mathbb{T}^{2})$ and is of class $C^{2}$ on the open set $\mathbb{S}^{2}\subseteq\mathbb{R}^{2}$.
The first property is a very natural property, having in mind Proposition 2.1 and the theory of direct integrals (Theorem 5.3). The second property of Definition 3.7 is needed to define group velocities and mass tensors.
To explain these two concepts, we need the Hessian of functions $f\in C^{2}(\mathbb{S}^{2})$ at fixed $k$, which is denoted by
$\mathrm{Hess}\left(f\right)\left(k\right)\doteq\left(\begin{array}[]{cc}\partial_{k_{1}}^{2}f&\partial_{k_{1}}\partial_{k_{2}}f\\ \partial_{k_{2}}\partial_{k_{1}}f&\partial_{k_{2}}^{2}f\end{array}\right)\left(k\right)\in\mathsf{M}_{2}\left(\mathbb{R}\right)\ ,\qquad k\in\mathbb{S}^{2}\ ,$ (45)
where $\mathsf{M}_{2}(\mathbb{R})$ is the set of $2\times 2$ matrices with real coefficients. It is a straightforward consequence of the regularity of $f\in C^{2}(\mathbb{S}^{2})$ that
$\mathrm{Hess}\left(f\right):\mathbb{S}^{2}\longrightarrow\mathsf{M}_{2}\left(\mathbb{R}\right)$
is continuous. For any $f\in C^{2}(\mathbb{S}^{2})$, we consider the set
$\mathfrak{M}_{f}\doteq\left\{k\in\mathbb{S}^{2}\,:\,\mathrm{Hess}\left(f\right)\left(k\right)\in\mathsf{GL}_{2}\left(\mathbb{R}\right)\right\}\subseteq\mathbb{S}^{2}$ (46)
with $\mathsf{GL}_{2}\left(\mathbb{R}\right)\subseteq\mathsf{M}_{2}\left(\mathbb{R}\right)$ being the set of invertible $2\times 2$ matrices with real coefficients. As $\mathsf{GL}_{2}(\mathbb{R})\subseteq\mathsf{M}_{2}(\mathbb{R})$ is an open set (see [44, Theorem 1.4]), it then follows that
$\mathfrak{M}_{f}=\mathrm{Hess}\left(f\right)^{-1}\left(\mathsf{GL}_{2}\left(\mathbb{R}\right)\right)$
is also an open set.
We are now in a position to define group velocities and mass tensors of a family of dispersion relations.
Definition 3.8 (Group velocities and mass tensors)
At any $\mathrm{U}\in\left[0,\infty\right]$, the group velocity $\mathbf{v}_{\varkappa,\mathrm{U}}:\mathbb{S}^{2}\rightarrow\mathbb{R}$ and the mass tensor $\mathbf{m}_{\varkappa,\mathrm{U}}:\mathfrak{M}_{\varkappa\left(\mathrm{U},\cdot\right)}\rightarrow\mathsf{M}_{2}\left(\mathbb{R}\right)$ associated with a family $\varkappa:\left[0,\infty\right]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ of dispersion relations are respectively defined by
$\mathbf{v}_{\varkappa,\mathrm{U}}\left(k\right)\doteq\vec{\nabla}_{k}\varkappa\left(\mathrm{U},k\right)\qquad\text{and}\qquad\mathbf{m}_{\varkappa,\mathrm{U}}\left(k\right)\doteq\mathrm{Hess}\left(\varkappa\left(\mathrm{U},\cdot\right)\right)\left(k\right)^{-1}\ .$
We deduce from Theorem 4.8 that $\mathrm{E}$ is a dispersion relation when the function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ is at least $2$ times continuously differentiable and, in this case, we can even compute the group velocity via the characteristic equation (Theorem 3.6).
Theorem 3.9 (Dispersion relations of dressed bound fermion pairs)
Fix $\epsilon\in\mathbb{R}_{0}^{+}$ and $h_{b}\in[0,1/2]$. Assume that $\hat{\upsilon}\in C^{2}(\mathbb{S}^{2})$.
  1. i.)
    Then, $\mathrm{E}:\left[0,\infty\right]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$ given by Theorems 3.1 and 3.5 is a family of dispersion relations.
  2. ii.)
    The associated group velocities are equal to
    $\mathbf{v}_{\mathrm{E},\mathrm{U}}\left(k\right)=\left.\left(\hat{\upsilon}\left(k\right)^{2}\partial_{x}\mathfrak{T}\left(\mathrm{U},k,x\right)+1\right)^{-1}\vec{\nabla}\left(\hat{\upsilon}\left(k\right)^{2}\mathfrak{T}\left(\mathrm{U},k,x\right)-\mathfrak{b}\left(k\right)\right)\right|_{x={\mathrm{E}\left(\mathrm{U},k\right)}}$
    for any $\mathrm{U}\in\left[0,\infty\right]$ and $k\in\mathbb{S}^{2}$, with
    $\mathbf{v}_{\mathrm{E},\infty}\left(k\right)=\lim_{\mathrm{U}\rightarrow\infty}\mathbf{v}_{\mathrm{E},\mathrm{U}}\left(k\right)\ ,\qquad k\in\mathbb{S}^{2}\ .$
  3. iii.)
    If $\hat{\upsilon}$ is real analytic on $\mathbb{S}^{2}$, then, for any $\mathrm{U}\in\left[0,\infty\right]$, either $\mathfrak{M}_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$ has full measure or $\mathfrak{M}_{\mathrm{E}}=\emptyset$. In particular, the tensor masses $\mathbf{m}_{\mathrm{E},\mathrm{U}}$ are either defined almost everywhere in $\mathbb{S}^{2}$ or not defined at all.
Proof. Use Corollaries 4.9 and 4.16.
Similar to Remark 3.2, if for some strictly positive constant $\gamma>0$,
$\sum_{x\in\mathbb{Z}^{2}}\mathrm{e}^{\gamma\left|x\right|}\left|\upsilon\left(x\right)\right|<\infty$
then the Fourier transform $\hat{\upsilon}$ of the function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$, as defined via (26), is real analytic on the whole torus $\mathbb{T}^{2}$. It is very natural to expect a local interaction between fermion pairs and bosons in (10), meaning here that the function $\upsilon$ should even have finite support. In particular, all conditions of Theorem 3.9, including the ones of the third assertion, should hold true in the application to superconducting cuprates.
In fact, as shown in [21], the dispersion relation of Theorem 3.9 yields the formation of $d$-wave pairs when one adjusts the parameters of the model (with $\mathrm{u}=0$) to fit those of cuprate superconductors, in particular the ones of the cuprate $\mathrm{La}_{2-x}\mathrm{Sr}_{x}\mathrm{CuO}_{4}$ (LaSr 214) near optimal doping. When considering the usual Hubbard model, that is, the case where there is no other fermionic repulsion than the on-site one (i.e., $\mathrm{u}=0$) and no fermion-boson exchange (i.e., $\upsilon=0$), $\mathrm{E}$ turns out to be the function $\mathfrak{b}:\mathbb{T}^{2}\rightarrow\mathbb{R}$, defined by (27), which is nothing else than the dispersion relation
$\mathfrak{b}(k)\doteq h_{b}\epsilon\left(2-\cos\left(k\right)\right)\ ,\qquad k\in\mathbb{T}^{2}\ ,$
of free bosons (bipolarons for cuprates). By turning on the fermion-boson-exchange interaction, the dispersion relation of dressed bound fermion pairs with lowest energy can strongly deviate from $\mathfrak{b}$, the unperturbed one. Recall for instance that $\mathfrak{b}$ describes bosons with a very large mass as compared to the effective mass of electrons or holes in cuprates. However, as shown in [21], for typical parameters of the cuprate LaSr 214, the effective mass of the bound pair (with dispersion relation $\mathrm{E}$) is comparable to the mass of electrons or holes. This solves the so-called “large mass paradox” of the microscopic theory of cuprate superconductors, based on some kind mechanism involving bipolarons. For more details, see [21] and references therein.
3.3 Quantum Scattering
Scattering in quantum mechanics constitutes a well-established mathematical theory aiming at analyzing the behavior of quantum systems at large times. To this end, a reference (or free) Hamiltonian $Y$ is chosen and the dynamics $(\mathrm{e}^{itX})_{t\in\mathbb{R}}$ of the quantum system driven by the (full) Hamiltonian $X$ is compared at large (negative and positive) times to $(\mathrm{e}^{itY})_{t\in\mathbb{R}}$. In fact, scattering theory can be viewed as a kind of perturbation theory for the absolutely continuous spectrum of $X$. See, e.g., [34, Chapter X]. For standard textbooks explaining in detail the scattering theory, we recommend [35, 36, 37]. Below, for the reader’s convenience, we shortly recall definitions that are relevant here.
Take two bounded self-adjoint operators $X$ and $Y$ acting on two Hilbert spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively. Let $P_{\mathrm{ac}}(Y)$ be the orthogonal projection onto the absolutely continuous space of $Y$, which is defined as follows:
$\mathrm{ran}\left(P_{\mathrm{ac}}(Y)\right)$ $\doteq$ $\big{\{}\psi\in\mathcal{Y}:\left\langle\psi,\chi_{(\cdot)}(Y)\psi\right\rangle_{\mathcal{Y}}\text{ is absolutely continuous}$ (47)
[…]
$W^{\pm}\left(X,Y;J\right)\doteq s-{\lim\limits_{t\rightarrow\pm\infty}}\mathrm{e}^{itX}J\mathrm{e}^{-itY}P_{\mathrm{ac}}\left(Y\right)\text{ },$ (48)
when it exists. See for instance [37, Definition 1.3]. When $\mathcal{Y}=\mathcal{X}$ and $J=\mathfrak{1}$, like in [37, Definition 1.1], we use the shorter notation
$W^{\pm}\left(X,Y\right)\equiv W^{\pm}\left(X,Y;\mathfrak{1}\right)\doteq s-{\lim\limits_{t\rightarrow\pm\infty}}\mathrm{e}^{itX}\mathrm{e}^{-itY}P_{\mathrm{ac}}\left(Y\right)\text{ }.$ (49)
In case the above wave operators exist, they are partial isometries [35, Proposition 1, Sect. XI.3]. They are said to be complete when
$\mathrm{ran}\left(W^{+}\left(X,Y\right)\right)=\mathrm{ran}\left(W^{-}\left(X,Y\right)\right)=\mathrm{ran}\left(P_{\mathrm{ac}}(X)\right)\ .$
See [35, p. 19, Sect. XI.3]. Similarly, in the general case, $W^{\pm}\left(X,Y;J\right)$ are said to be complete whenever
$\overline{\mathrm{ran}\left(W^{+}\left(X,Y;J\right)\right)}=\overline{\mathrm{ran}\left(W^{-}\left(X,Y;J\right)\right)}=\mathrm{ran}\left(P_{\mathrm{ac}}(X)\right)\ .$
See [35, p. 35, Sect. XI.3]. The corresponding scattering operator is equal to
$S\left(X,Y;J\right)\doteq W^{+}\left(X,Y;J\right)^{\ast}W^{-}\left(X,Y;J\right)\in\mathcal{B}\left(\mathcal{Y}\right)\ .$ (50)
It leads to the scattering matrix (or simply $S$-matrix) in a representation where $Y$ is diagonal, because the scattering operator commutes with $Y$. See [37, Equation (1.12)].
Remark 3.10
For two bounded self-adjoint operators $X$ and $Y$ acting on two Hilbert spaces $\mathcal{X}$ and $\mathcal{Y}$, note [^8] To show this property, take any Borel set $\Omega\subseteq\mathbb{R}$ and observe that $\langle(\varphi,\psi),\chi_{\Omega}(X\oplus Y)(\varphi,\psi)\rangle_{\mathcal{X}\oplus\mathcal{Y}}=\langle\varphi,\chi_{\Omega}(X)\varphi\rangle_{\mathcal{X}}+\langle\psi,\chi_{\Omega}(Y)\psi\rangle_{\mathcal{Y}}.$ Since $\chi_{\Omega}(X)$ and $\chi_{\Omega}(Y)$ are positive operators, the left-hand side of the last expression is zero iff each term in the right-hand side is zero. that
$\mathrm{ran}\left(P_{\mathrm{ac}}(X\oplus Y)\right)=\mathrm{ran}\left(P_{\mathrm{ac}}(X)\right)\oplus\mathrm{ran}\left(P_{\mathrm{ac}}(Y)\right)\ .$
This is an elementary observation used to study the scattering channels in Section 4.5.
In our framework, the Hamiltonian $X$ is the bounded self-adjoint operator
$UHU^{\ast}={\int_{\mathbb{T}^{2}}^{\oplus}}A(k)\,\nu(\mathrm{d}k)$
of Proposition 2.1, which acts on the Hilbert space $L^{2}(\mathbb{T}^{2},\mathcal{H})$. Below, two different (reference) Hamiltonians $Y$ are taken into account, corresponding to two scattering channels: the unbound and bound pair channels. For cuprates, the first channel should be associated with the high temperature regime, while the second one is related to sufficiently low temperatures.
3.3.1 Unbound pair scattering channel
Far apart from each other, two fermions only experience a very weak repulsion force due to the extended Hubbard interaction while the probability that they bind together to form a boson is also very small. Thus, in this situation, one expects that the dynamics of such a pair is governed by the fermionic part, and even by the hopping term alone. During intermediate times they could of course interact, as they may get close to each other, and they could even be bound together via the effective attraction caused by fermion-boson exchange processes. The lifetime of bound fermions should however be finite in this situation and they are expected to be released at some point and behave again as two free fermions that go far apart from each other for large times. We show below that this heuristics can be put in precise mathematical terms.
To this end, define the Hilbert space
$\mathfrak{H}_{f}\doteq L^{2}\left(\mathbb{T}^{2},L^{2}\left(\mathbb{T}^{2}\right),\nu\right)\doteq\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2},\mathbb{C},\nu\right)\,\nu\left(\mathrm{d}k\right)$ (51)
as well as the Hamiltonian
$\mathrm{H}_{f}\equiv\mathrm{H}_{f}\left(\mathrm{V},\mathrm{v}\right)\doteq{\int_{\mathbb{T}^{2}}^{\oplus}}\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\,\nu(\mathrm{d}k)\in\mathcal{B}\left(\mathfrak{H}_{f}\right)$ (52)
for any $\mathrm{V}\in\mathbb{R}_{0}^{+}$ and absolutely summable function $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$, where
$R\left(\mathrm{V},\mathrm{v}\right)\doteq{\sum\limits_{x\in\mathbb{Z}^{2}}}\,\mathrm{v}\left(x\right)P_{x}+\mathrm{V}P_{0}\in\mathcal{B}\left(L^{2}(\mathbb{T}^{2})\right)\ ,$ (53)
$M_{\mathfrak{f}\left(k\right)}$ being the fiber Hamiltonian defined as the multiplication operator by $\mathfrak{f}(k)\in C(\mathbb{T}^{2})$ (see (28)) while $P_{x}$ is the orthogonal projection onto the one-dimensional subspace $\mathbb{C}\mathfrak{\hat{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. Compare (52) with Equations (40)–(31) and Proposition 2.1. By Lemma 4.20, note that $P_{\mathrm{ac}}(\mathrm{H}_{f})=\mathfrak{1}$.
Let us consider the identification operator $\mathfrak{U}:\mathfrak{H}_{f}\rightarrow L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$ defined for any purely fermionic state $\psi\in\mathfrak{H}_{f}$ by
$\begin{array}[]{cccl}\mathfrak{U}\psi:&\mathbb{T}^{2}&\rightarrow&\mathcal{H}\doteq L^{2}\left(\mathbb{T}^{2}\right)\oplus\mathbb{C}\\ &k&\mapsto&\left(\psi\left(k\right),0\right)\end{array}\ .$ (54)
See Equation (23). Note that $\mathfrak{U}$ is an isometry, i.e., a norm preserving linear transformation. In fact, it is the canonical fiberwise inclusion of $\mathfrak{H}_{f}$ into $L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$. Recall from Proposition 2.1 that $A\left(\mathrm{U},\cdot\right)$, defined by (31)–(40), is the fiber decomposition of the operator $UHU^{\ast}$. Then, we obtain wave and scattering operators with respect to fermionic parts:
Theorem 3.11 (Unbound pair (scattering) channel)
Let $\mathrm{V}\in\mathbb{R}_{0}^{+}$ and $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ be any absolutely summable function.
  1. i.)
    Wave operators (see (48)):
    $W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)=\left({\int_{\mathbb{T}^{2}}^{\oplus}}W^{\pm}\left(A\left(k\right),\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}$
    with range equal to
    $\mathrm{ran}\left(W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\right)=\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\,\nu\left(\mathrm{d}k\right)\ .$
  2. ii.)
    Scattering operator (see (50)):
    $S\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)=\mathfrak{U}^{\ast}\left({\int_{\mathbb{T}^{2}}^{\oplus}}S\left(A\left(k\right),\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}\ .$
Proof. Observe that the operator difference $(UHU^{\ast}-\mathrm{H}_{f})$ is not trace-class (it is not even compact) and, thus, the existence of this scattering channel is not a direct consequence of the well-known Kato-Rosenblum theorem [35, Theorem XI.8]. In fact, one of the main steps of the proof is to show that this difference is the direct integral of a strongly measurable family of trace-class operators. By this means we are then able to apply the Kato-Rosenblum theorem “fiberwise” to deduce the first assertion. See Section 4.5.1 for more details, in particular Theorem 4.22. Assertion (ii) is a direct consequence of Assertion (i) together with the theory of direct integrals.
Theorem 3.11 refers to the unbound pair (scattering) channel. The subspace $\mathfrak{H}_{f}\subseteq\mathfrak{H}$ corresponds to the “incoming” ($+$) and “outcoming” (--) scattering states of the quantum system, in this particular scattering channel. Physically, this theorem shows, among other things, that the bosonic component of eitH vanishes on this channel, as $t\rightarrow\pm\infty$.
Remark 3.12
If one would like to go back to the original Hilbert space $\mathfrak{h}_{0}$ (17) for fermion pairs with opposite spins, that is, if one wishes to use space coordinates, instead of the quasi-momenta, then one employs Theorem 3.11, along with the observation that
$W^{\pm}\left(H,U_{f}^{\ast}\mathrm{H}_{f}U_{f};U^{\ast}\mathfrak{U}U_{f}\right)=U^{\ast}W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)U_{f}\ ,$
where $U_{f}$ is defined by (25). See also Equation (24), Proposition 2.1 and Section 5.1.
Note that, in any given fiber, the wave operators, as well as the scattering operators, can be computed even more explicitly via infinite series (perturbative expansions), thanks to Corollary 5.2. Below we give an example of such a computation by taking $\mathrm{U}=\mathrm{V}\in\mathbb{R}_{0}^{+}$ and $\mathrm{v}=\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ in Theorem 3.11. With this particular choice, the unbound pair channel allows one to isolate the fermion-boson exchange mechanism, which in terms of Hamiltonians refers to the use of off-diagonal operators
$B^{(t)}\left(k\right)\doteq\begin{pmatrix}0&B_{1,2}^{(t)}\left(k\right)\\[5.0pt] B_{2,1}^{(t)}\left(k\right)&0\end{pmatrix}\in\mathcal{B}\left(\mathcal{H}\right)\ ,\qquad t\in\mathbb{R},\ k\in\mathbb{T}^{2}\ ,$ (55)
in the fibers, where, for any $t\in\mathbb{R}$ and $k\in\mathbb{T}^{2}$,
$B_{1,2}^{(t)}\left(k\right)\doteq\mathrm{e}^{itA_{1,1}\left(\mathrm{U},k\right)}A_{1,2}\left(k\right)\mathrm{e}^{-itA_{2,2}\left(k\right)}\qquad\text{and}\qquad B_{2,1}^{(t)}\left(k\right)\doteq\mathrm{e}^{itA_{2,2}\left(k\right)}A_{2,1}\left(k\right)\mathrm{e}^{-itA_{1,1}\left(\mathrm{U},k\right)}\ .$ (56)
Recall that, for $m,n\in\{1,2\}$, $A_{m,n}(k)$ is defined by (31)–(39). Below, $B_{m,n}(k)$, $m\neq n$, stands for the norm-continuous family of operators $(B_{m,n}^{(t)}(k))_{t\in\mathbb{R}}$.
To shorten the notation, for any $s,t\in\mathbb{R}$, as well as two norm-continuous families $X\equiv(X_{t})_{t\in\mathbb{R}}$ and $Y\equiv(Y_{t})_{t\in\mathbb{R}}$ of bounded operators $X_{t}:\mathcal{X}\rightarrow\mathcal{Y}$ and $Y_{t}:\mathcal{Y}\rightarrow\mathcal{X}$ on two Hilbert spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively, we define the bounded operators:
$\cos_{\succ}\left(XY;s,t\right)$ $\doteq$ $\mathbf{1}+\sum_{p=1}^{\infty}\left(-1\right)^{p}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{2p-1}}\mathrm{d}\tau_{2p}(X_{\tau_{1}}Y_{\tau_{2}})\cdots(X_{\tau_{2p-1}}Y_{\tau_{2p}})\ ,$ (57)
$\sin_{\succ}\left(XY;s,t\right)$ $\doteq$ $\int_{s}^{t}\mathrm{d}\tau X_{\tau}+\sum_{p=1}^{\infty}\left(-1\right)^{p}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{2p}}\mathrm{d}\tau_{2p+1}X_{\tau_{1}}\left((Y_{\tau_{2}}X_{\tau_{3}})\cdots(Y_{\tau_{2p}}X_{\tau_{2p+1}})\right)\ .$
The integrals above are Riemann ones, noting that $(X_{t})_{t\in\mathbb{R}}$ and $(Y_{t})_{t\in\mathbb{R}}$ are continuous families in Banach spaces, namely $\mathcal{B}\left(\mathcal{X};\mathcal{Y}\right)$ and $\mathcal{B}\left(\mathcal{Y};\mathcal{X}\right)$, respectively. Note that $\cos_{\succ}\left(XY;s,t\right)\in\mathcal{B}\left(\mathcal{Y}\right)$ and $\sin_{\succ}\left(XY;s,t\right)\in\mathcal{B}\left(\mathcal{X},\mathcal{Y}\right)$ are always absolutely summable series in the operator norm. Then, we obtain the following results:
Theorem 3.13 (Scattering operators as pertubative series)
Let $\varepsilon\in\mathbb{R}^{+}$ and $\mathrm{H}_{f}\equiv\mathrm{H}_{f}\left(\mathrm{U},\mathrm{u}\right)$. Then, for any $\varphi\in\mathfrak{H}_{f}$, there is $T>0$ such that
$T$ $<$ $t\implies\left\|\left(W^{+}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)-V_{0,t}\mathfrak{U}\right)\varphi\right\|_{\mathcal{X}}\leq\varepsilon\ ,$
$t$ $<$ $-T\implies\left\|\left(W^{-}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)-V_{0,t}\mathfrak{U}\right)\varphi\right\|_{\mathcal{X}}\leq\varepsilon\ ,$
Moreover, for any $\varphi,\psi\in\mathfrak{H}_{f}$, there is $T>0$ such that
$s<-T<T<t\implies\left\langle\psi,S\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\varphi\right\rangle_{\mathcal{X}}=\left\langle\mathfrak{U}\psi,V_{t,s}\mathfrak{U}\varphi\right\rangle_{\mathcal{X}}+\mathcal{O}\left(\varepsilon\right)\ ,$
where, for all $s,t\in\mathbb{R}$,
$V_{t,s}\doteq{\int_{\mathbb{T}^{2}}^{\oplus}}\left(\begin{array}[]{cc}\cos_{\succ}\left(B_{1,2}\left(k\right)B_{2,1}\left(k\right);s,t\right)&-i\sin_{\succ}\left(B_{1,2}\left(k\right)B_{2,1}\left(k\right);s,t\right)\\ -i\sin_{\succ}\left(B_{2,1}\left(k\right)B_{1,2}\left(k\right);s,t\right)&\cos_{\succ}\left(B_{2,1}\left(k\right)B_{1,2}\left(k\right);s,t\right)\end{array}\right)\nu\left(\mathrm{d}k\right)\ .$
Proof. It suffices to combine Lemma 4.23 with Equation (129) and Theorem 3.11, similar to Corollary 5.2.
Theorem 3.13 provides a way to approximate the scattering matrix associated with the fermion-boson-exchange interaction. Note for instance from (31)–(39) that the operator $B_{1,2}^{(t)}\left(k\right)B_{2,1}^{(s)}\left(k\right)$ and $B_{2,1}^{(t)}\left(k\right)B_{1,2}^{(s)}\left(k\right)$ have a relatively simple form for any $s,t\in\mathbb{R}$ and $k\in\mathbb{T}^{2}$:
$B_{1,2}^{(t)}\left(k\right)B_{2,1}^{(s)}\left(k\right)$ $=$ $\left(\hat{\upsilon}\left(k\right)^{2}\mathrm{e}^{i\left(s-t\right)\mathfrak{b}\left(k\right)}\right)\mathrm{e}^{itA_{1,1}\left(\mathrm{U},k\right)}P_{\mathfrak{d}\left(k\right)}\mathrm{e}^{-isA_{1,1}\left(\mathrm{U},k\right)}\ ,$
$B_{2,1}^{(t)}\left(k\right)B_{1,2}^{(s)}\left(k\right)$ $=$ $\left(\hat{\upsilon}\left(k\right)^{2}\mathrm{e}^{i\left(t-s\right)\mathfrak{b}\left(k\right)}\right)\left\langle\mathfrak{d}\left(k\right),\mathrm{e}^{i(s-t)A_{1,1}\left(\mathrm{U},k\right)}\mathfrak{d}\left(k\right)\right\rangle\ ,$
where $P_{\mathfrak{d}\left(k\right)}$ is the orthogonal projection onto the one-dimensional subspace $\mathbb{C}\mathfrak{d}\left(k\right)\subseteq L^{2}(\mathbb{T}^{2})$. Similar computations can be done for other choices of $\mathrm{V}\in\mathbb{R}_{0}^{+}$ and $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ in Theorem 3.11, like $\mathrm{V}=0=\mathrm{v}$ (non-interacting fermion systems). See again Corollary 5.2.
3.3.2 Bound pair scattering channel
Similarly, we also prove the existence of a scattering channel for dressed bound pairs. As dressed bound pairs are space-localized objects (see, e.g., Theorem 3.1 (iii)), the fermions forming the pair efficiently exchange a boson, at a non-negligible rate, via the terms $W_{\mathrm{b\rightarrow f}}$ (9) and $W_{\mathrm{f\rightarrow b}}$ (15) in the Hamiltonian $H$. In particular, such quantum states must have some non-negligible bosonic component representing the exchanged boson that “glues” the two fermions together. This dressed bound pair is however expected to move like a free (quantum spinless) particle in the real space. We translate this physical heuristics in precise mathematical terms by considering the effective dispersion relations
$\mathrm{E}:\left[0,\infty\right]\times\mathbb{T}^{2}\rightarrow\mathbb{R}$
given by Theorems 3.1, 3.5 and 3.9.
For any $\mathrm{U}\in\mathbb{R}_{0}^{+}\cup\{\infty\}$, we consider the identification operator
$\mathfrak{P}_{\mathrm{U}}:L^{2}\left(\mathbb{T}^{2}\right)\rightarrow L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$
defined for any $\varphi\in L^{2}\left(\mathbb{T}^{2}\right)$ by [^9] Despite the fact that $\mathrm{E}(0)$ might not be in the resolvent set of $A_{1,1}(\mathrm{U},0)$, we can simply ignore it for $\{k=0\}$ has null Lebesgue measure.
$\begin{array}[]{cccl}\mathfrak{P}_{\mathrm{U}}\varphi:&\mathbb{T}^{2}\backslash\{0\}&\rightarrow&\mathcal{H}\doteq L^{2}\left(\mathbb{T}^{2}\right)\oplus\mathbb{C}\\ &k&\mapsto&\varphi\left(k\right)\left\|\Psi\left(\mathrm{U},k\right)\right\|^{-1}\Psi\left(\mathrm{U},k\right)\end{array}\ ,$ (59)
where $\Psi\left(\mathrm{U},k\right)$ is the eigenvector associated with the (non-degenerate) eigenvalue $\mathrm{E}\left(\mathrm{U},k\right)$, as given by Theorems 3.1, 3.5 and 3.9. Note from Theorem 3.1 that the mapping
$k\mapsto\left\|\Psi\left(\mathrm{U},k\right)\right\|^{-1}\Psi\left(\mathrm{U},k\right)$
is continuous on $\mathbb{T}^{2}$ for any $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and its pointwise limit
$k\mapsto\left\|\Psi\left(\infty,k\right)\right\|^{-1}\Psi\left(\infty,k\right)$
(cf. Theorem 3.5) is therefore measurable. In particular, the linear transformation $\mathfrak{P}_{\mathrm{U}}$ is well defined for any $\mathrm{U}\in\mathbb{R}_{0}^{+}\cup\{\infty\}$. Moreover, one checks that it is norm-preserving.
Since $\mathrm{E}\left(\mathrm{U},\cdot\right)\in C(\mathbb{T}^{2};\mathbb{R})\subseteq L^{\infty}(\mathbb{T}^{2},\nu)$ (see Theorems 3.1 (iv) and 3.5), we can consider the multiplication operator by $\mathrm{E}\left(\mathrm{U},\cdot\right)$ on $L^{2}(\mathbb{T}^{2})$, which is denoted by
$M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\doteq{\int_{\mathbb{T}^{2}}^{\oplus}}\mathrm{E}\left(\mathrm{U},k\right)\,\nu\left(\mathrm{d}k\right)\ ,\qquad\mathrm{U}\in\mathbb{R}_{0}^{+}\cup\{\infty\}\ .$ (60)
Remark also from Theorems 3.1 (iv) and 3.5 together with Corollary 5.5 that $P_{\mathrm{ac}}(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)})=\mathfrak{1}$ whenever $\hat{\upsilon}$ is real analytic on $\mathbb{S}^{2}$. We then study now the (dressed) bound pair scattering channel, which is much simpler than in the unbound pair channel:
Theorem 3.14 (Bound pair (scattering) channel)
Let $h_{b}\in[0,1/2]$. Then the following assertions hold true:
  1. i.)
    Dynamics and wave operators at finite $\mathrm{U}\in\mathbb{R}_{0}^{+}$:
    $\mathrm{e}^{itUHU^{\ast}}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}\mathrm{e}^{itM_{\mathrm{E}\left(\mathrm{U},\cdot\right)}}\ ,\qquad t\in\mathbb{R}\ .$
  2. ii.)
    Dynamics in the hard-core limit $\mathrm{U}\rightarrow\infty$:
    $s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\infty}\qquad\text{and}\qquad s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,\mathrm{e}^{itUHU^{\ast}}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\infty}\mathrm{e}^{itM_{\mathrm{E}\left(\mathrm{\infty},\cdot\right)}}\ ,\qquad t\in\mathbb{R}\ .$
Proof. Assertion (i) is Proposition 4.24. Assertion (ii) results from Proposition 4.25.
From Theorem 3.14 (i) the scattering channel is time independent. For instance, for any $\mathrm{U}\in\mathbb{R}_{0}^{+}$, one trivially checks that
$W^{\pm}\left(UHU^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=\mathfrak{P}_{\mathrm{U}}P_{\mathrm{a}\mathrm{c}}\left(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\right)$
and, since $\mathfrak{P}_{\mathrm{U}}^{\ast}\mathfrak{P}_{\mathrm{U}}=\mathfrak{1}$ and $P_{\mathrm{ac}}$ is a projection, its scattering operator is equal to
$S\left(UHU^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=P_{\mathrm{a}\mathrm{c}}\left(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\right)\ .$
If $\hat{\upsilon}$ is additionally real analytic on $\mathbb{S}^{2}$, then $P_{\mathrm{ac}}(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)})=\mathfrak{1}$, thanks to Theorem 3.1 and Corollary 5.5. In this case, the wave and scattering operators are given by
$W^{\pm}\left(UHU^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=\mathfrak{P}_{\mathrm{U}}\qquad\text{and}\qquad S\left(UHU^{\ast},M_{\mathrm{E}\left(\mathrm{U},\cdot\right)};\mathfrak{P}_{\mathrm{U}}\right)=\mathfrak{1}$
for any $\mathrm{U}\in\mathbb{R}_{0}^{+}$. Their hard-core limit are then also trivial, thanks to Theorem 3.14 (ii).
This scattering channel is therefore easy to study. In particular, similar to Remark 3.12, we can easily go back to the original Hilbert space $\mathfrak{h}_{0}$ (17), referring to space coordinates instead of the quasi-momenta. With this aim, we first observe that, for any $\mathrm{U}\in\mathbb{R}_{0}^{+}$,
$\mathrm{e}^{itH}\mathcal{P}_{\mathrm{U}}=\mathcal{P}_{\mathrm{U}}\mathrm{e}^{itU_{f}^{\ast}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}U_{f}}\ ,\qquad t\in\mathbb{R}\ ,$ (61)
where $\mathcal{P}_{\mathrm{U}}\in\mathcal{B}\left(\mathfrak{h}_{0},\mathfrak{H}\right)$ is the new identification operator
$\mathcal{P}_{\mathrm{U}}\doteq U^{\ast}\mathfrak{P}_{\mathrm{U}}U_{f}\ ,\qquad\mathrm{U}\in\mathbb{R}_{0}^{+}\ ,$
and $U_{f}\doteq U_{2}U_{1}$. See Equations (24)–(25), Proposition 2.1 and Section 5.1.
On the one hand, Equation (61) together with Theorem 3.9 shows that $\mathrm{E}\left(\mathrm{U},\cdot\right)$ defines (a family of) dispersion relations, in the sense of Definition 3.7. The Fourier transform of $\mathrm{E}\left(\mathrm{U},\cdot\right)$ is the (effective) hopping strength for the (spatially localized) dressed bound pairs. On the other hand, the new identification operator $\mathcal{P}_{\mathrm{U}}$ is translation invariant, i.e.,
$\mathcal{P}_{\mathrm{U}}\theta_{x}=\Theta_{x}\mathcal{P}_{\mathrm{U}}\ ,\qquad x\in\mathbb{Z}^{2}\ ,$
where, for any fixed $x\in\mathbb{Z}^{2}$, $\theta_{x}\in\mathcal{B}(\mathfrak{h}_{0})$ and $\Theta_{x}\in\mathcal{B}(\mathfrak{H})$ are the unique unitary operators respectively satisfying
$\theta_{x}\left(\mathfrak{e}_{(y,\uparrow)}\wedge\mathfrak{e}_{(z,\downarrow)}\right)=\mathfrak{e}_{(y+x,\uparrow)}\wedge\mathfrak{e}_{(z+x,\downarrow)}\ ,\qquad y,z\in\mathbb{Z}^{2}\ ,$
(see Equations (17)) and
$\Theta_{x}\left(\mathfrak{\psi}\oplus\varphi\right)=\left(\theta_{x}\mathfrak{\psi}\right)\oplus\varphi\left(x+\cdot\right)\ ,\qquad\mathfrak{\psi\in h}_{0},\ \varphi\in\ell^{2}\left(\mathbb{Z}^{2}\right)\ .$
See Equations (18). In addition, when $\hat{\upsilon}(0)\neq 0$, Theorem 3.1 (iii) shows that the (dressed) fermion pair in the bound pair channel is exponentially localized in space, that is, the associate fermion-fermion correlation function decays exponentially fast with respect to the distance between the fermions, uniformly in time. Note that it is not required that the range of $\mathcal{P}_{\mathrm{U}}$ has a vanishing bosonic component, because of the expected presence of “gluing bosons” in the dressed bound fermionic pair.
As a consequence, the bound channel describes an effective system of free localized, spinless quasi-particles which minimize the energy at any fixed total quasi-momentum. In particular, such quasi-particles of lowest energy, or dressed fermion pairs, are stable in time, that is, they cannot decay into an (even only asymptotically) unbound pair of fermions. Conversely, we also show in Section 3.3.1 that a pair of fermions that is asymptotically unbound far in the past is not able to bind together to form a stable bound pair in the distant future.
Nevertheless, these quasi-particles should only be stable with respect to external perturbations as soon as their states are related to quasi-momenta $k$ such that $\mathrm{E}\left(\mathrm{U},k\right)<0$. If the (dressed) quasi-particle is in a state whose support contains fibers $k$ such that $\mathrm{E}\left(\mathrm{U},k\right)\geq 0$, it is not in the most energetically favorable state, since
$\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},0\right)\right)=0$
(see Theorem 3.1). In fact, if the component corresponding to quasi-momenta $k$ such that $\mathrm{E}\left(\mathrm{U},k\right)\geq 0$ has non-vanishing Lebesgue measure, then the quasi-particle should be instable with respect to external perturbations, by possibly creating unbounded fermions with small quasi-momenta to decrease its total energy. This situation clearly appears for quasi-momenta $k\in\mathbb{T}^{2}$ such that $\hat{\upsilon}(k)=0$ or sufficiently small $\left|\hat{\upsilon}(k)\right|\ll 1$ when $k\neq 0$, since in these two cases, either $\mathrm{E}\left(\mathrm{U},k\right)=\mathfrak{b}\left(k\right)$ ($\hat{\upsilon}(k)=0$) or $\mathrm{E}\left(\mathrm{U},k\right)\simeq\mathfrak{b}\left(k\right)$ ($\left|\hat{\upsilon}(k)\right|\ll 1$) with $\mathfrak{b}\left(k\right)\doteq h_{b}\epsilon\left(2-\cos\left(k\right)\right)$ (see (27)). If such a decay process really occurs, one should see critical quasi-momenta, like in physical superconductors.
To prevent from this situation, one needs sufficiently strong $\left|\hat{\upsilon}\left(k\right)\right|\gg 1$ to have $\mathrm{E}\left(\mathrm{U},k\right)<0$ for all $k\in\mathbb{T}^{2}$. In the position space, this means that the exchange strength between the two fermions and the boson, represented by the function $\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ appearing in (9)–(15), has to be sufficiently strong and localized, like in Remark 3.2, in order to get a sufficiently strong “gluing effect” for dressed pairs, at all quasi-momenta. Recall also that the boson should be heavier than the two fermions, i.e., $h_{b}\in[0,1/2]$.
Last but not least, all this discussion can be extended to the hard-core limit $\mathrm{U}\rightarrow\infty$, in view of Theorems 3.9 and 3.14 (ii).
4 Technical Results
4.1 Notation
The purpose of this section is to fix (or recall) the notation and terminology that is used throughout the rest of the article. Let $\mathcal{X}$ be any complex Hilbert space. We denote its scalar product by $\langle\cdot,\cdot\rangle_{\mathcal{X}}$, with the convention that it is antilinear in the first argument and linear in the second one. The norm of $\mathcal{X}$ is thus
$\left\|\varphi\right\|_{\mathcal{X}}\doteq\sqrt{\left\langle\varphi,\varphi\right\rangle_{\mathcal{X}}}\ ,\qquad\varphi\in\mathcal{X}\ .$
[…]
Recall that $\mathcal{B}(\mathcal{X})$ denotes the set of bounded (linear) operators on $\mathcal{X}$. $\mathfrak{1}\equiv\mathfrak{1}_{\mathcal{X}}\in\mathcal{B}(\mathcal{X})$ is the identity operator. Given $T\in\mathcal{B}(\mathcal{X})$, $T^{\ast}$ denotes its adjoint operator. The (full) spectrum, essential spectrum and resolvent set of any $T\in\mathcal{B}(\mathcal{X})$ are denoted by $\sigma(T)$, $\sigma_{\mathrm{ess}}(T)$ and $\rho(T)$, respectively. The operator norm of $\mathcal{B}(\mathcal{X})$ is
$\|T\|_{\mathrm{op}}\doteq\sup\left\{\left\|T\varphi\right\|_{\mathcal{X}}:\varphi\in\mathcal{X}\ \text{with }\left\|\varphi\right\|_{\mathcal{X}}=1\right\}.$
Given $T\in\mathcal{B}(\mathcal{X})$ , $\mathcal{E}_{T}(\lambda)$ stands for the eigenspace associated with an eigenvalue $\lambda\in\sigma(T)$ of $T$.
In all the Section 4, we study properties of the Hamiltonian $H\in\mathcal{B}(\mathfrak{H})$ defined by (19). As one can see from (16)–(19) combined with (6), (8)–(11) and (15), it depends on several parameters. More precisely, $\epsilon,\mathrm{U},h_{b}\in\mathbb{R}_{0}^{+}$ and $\alpha_{0}\in\mathbb{R}^{+}$, while
$\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}\ ,\quad\mathrm{e}^{\alpha_{0}\left|\cdot\right|}\mathfrak{p}:\mathbb{Z}^{2}\rightarrow\mathbb{R}\ ,\quad\text{and}\quad\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$
are all absolutely summable functions that are invariant with respect to $90^{\circ}$-rotations. See Equations (7), (12) and (14). The parameters of the operator $H$ are always fixed and arbitrary, unless we need to specify them to clarify some particular statement. Recall that the invariance under $90^{\circ}$-rotation is not that important here. In fact, here, the only important point concerning this symmetry is that it implies that the Fourier transforms $\hat{\upsilon}$ and $\hat{\mathfrak{p}}$ are real-valued, because
$\upsilon(-x)=\upsilon(x)=\overline{\upsilon(x)}\text{ \ \ and \ \ }\mathfrak{p}(-x)=\mathfrak{p}(x)=\overline{\mathfrak{p}(x)}\text{ },$
i.e., the real valued functions $\upsilon$ and $\mathfrak{p}$ are reflection invariant, as a consequence of their $90^{\circ}$-rotation invariance. Apart of this technical point, it is mainly relevant for the study of unconventional pairings, which is not done in the present work.
Note additionally that the on-site repulsion $\mathrm{U}\in\mathbb{R}_{0}^{+}$ appears explicitly in all the quantities defined in Sections 23. However, in Section 4, this parameter is only important for the Subsections 4.3 and 4.5. Therefore, unless the parameter $\mathrm{U}$ is important for our discussions or statements, we omit it in order to shorten the notation, by writing
$f\left(k\right)\equiv f\left(\mathrm{U},k\right)$
for any function $f(\mathrm{U},k)$ of the parameters $\mathrm{U}$ and $k$.
4.2 Spectrum of the Fiber Hamiltonians
We start with the study of the essential spectrum of fiber Hamiltonians (40) at any total quasi-momentum $k\in\mathbb{T}^{2}$, before considering afterwards the discrete one in Section 4.2.2. Then, in Section 4.2.3 we study the bottom of the spectrum ($k\in\mathbb{T}^{2}$ being fixed). The whole study leads to important spectral properties of $H$, as previously explained, via Proposition 2.1 combined with Theorem 5.3.
4.2.1 Essential spectrum
The essential spectrum $\sigma_{\mathrm{ess}}(A(k))$ of the fiber Hamiltonian
$A\left(k\right)\equiv A\left(\mathrm{U},k\right)\ ,$
defined by (40) at fixed $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and total quasi-momentum $k\in\mathbb{T}^{2}$, is completely determined by the following proposition:
Proposition 4.1 (Essential spectrum of fiber Hamiltonians)
For any $k\in\mathbb{T}^{2}$ and $h_{b},\epsilon,\mathrm{U}\in\mathbb{R}_{0}^{+}$, one has
$\sigma_{\mathrm{ess}}\left(A\left(k\right)\right)=\sigma_{\mathrm{ess}}\left(A_{1,1}\left(k\right)\right)=\sigma_{\mathrm{ess}}\left(B_{1,1}\left(k\right)\right)=\sigma\left(M_{\mathfrak{f}\left(k\right)}\right)=2\epsilon\cos\left(k/2\right)\left[-1,1\right]+4\epsilon\text{ },$
where $M_{\mathfrak{f}\left(k\right)}$ stands for the multiplication operator associated with the function $\mathfrak{f}(k)$ (28), while $B_{1,1}(k)$ and $A_{1,1}(k)\equiv A_{1,1}(\mathrm{U},k)$ are defined by (31) and (32), respectively.
Proof. Recall that $\nu$ is the normalized Haar measure defined by (21) on $\mathbb{T}^{2}$. Fix $k\in\mathbb{T}^{2}$. If $\lambda$ is an eigenvalue of $M_{\mathfrak{f}\left(k\right)}$ with associated eigenvector $\varphi\in L^{2}(\mathbb{T}^{2})$, then
$M_{\mathfrak{f}\left(k\right)}\varphi\left(p\right)\doteq\mathfrak{f}\left(k\right)\left(p\right)\varphi\left(p\right)=\lambda\varphi\left(p\right)$
for almost every $p\in\mathbb{T}^{2}$. As $\varphi\neq 0$, there exists $\Omega\subseteq\mathbb{T}^{2}$ with strictly positive measure $\nu(\Omega)>0$ such that the above equality holds true with $\varphi(p)\neq 0$ for every $p\in\Omega$. Thus, $\mathfrak{f}(k)(p)=\lambda$ for all $p\in\Omega$. Because
$\nu([-\pi,\pi)^{2}\backslash(-\pi,\pi)^{2})=0\ ,$
we can assume without loss of generality that $\Omega\subseteq(-\pi,\pi)^{2}$. Since $\mathfrak{f}(k)-\lambda$ is real analytic on the open domain $(-\pi,\pi)^{2}$ in $\mathbb{R}^{2}$ and the zeros of any non-constant real analytic function have null Lebesgue measure [43], we would have $\nu(\Omega)=0$, which contradicts our choice of the set $\Omega$. Recall indeed that $\nu$ is the Lebesgue measure, up to a normalization constant (see (21)). Hence, $M_{\mathfrak{f}\left(k\right)}$ has no eigenvalues and, thus,
$\sigma_{\mathrm{ess}}\left(M_{\mathfrak{f}\left(k\right)}\right)=\sigma\left(M_{\mathfrak{f}\left(k\right)}\right)=\,\mathfrak{f}\left(k\right)(\mathbb{T}^{2})\text{ }.$
The last equality holds true, for $\mathfrak{f}$ is a continuous function on a compact domain, namely the torus $\mathbb{T}^{2}$. Clearly,
$\mathfrak{f}\left(k\right)(\mathbb{T}^{2})=2\epsilon\cos\left(k/2\right)\left[-1,1\right]+4\epsilon\text{ }.$
Observing that $A_{1,2}(k)$, $A_{2,1}(k)$, $A_{2,2}(k)$ and $P_{x}$ are all rank-one linear transformations, we can apply the stability of the essential spectrum under compact perturbations (see [49, Corollary 8.16]) to conclude that
$\sigma_{\mathrm{ess}}\left(A\left(k\right)\right)=\sigma_{\mathrm{ess}}\left(M_{\mathfrak{f}\left(k\right)}\right)=\,\mathfrak{f}\left(k\right)(\mathbb{T}^{2})\text{ }.$ (62)
In fact, from the absolute summability of the function $\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ (see (7)), along with the closedness of the subspace of compact operators in the Banach space of all bounded operators, the operator defined by the infinite sum
${\sum\limits_{x\in\mathbb{Z}^{2}}}\,\mathrm{u}\left(x\right)P_{x}$
is not only bounded, but even compact on the Hilbert space $L^{2}(\mathbb{T}^{2})$. Recall that $P_{x}$ denotes the orthogonal projection onto the one-dimensional subspace $\mathbb{C}\hat{\mathfrak{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$ for any $x\in\mathbb{Z}^{2}$. For the same reasons,
$\sigma_{\mathrm{ess}}\left(A_{1,1}\left(k\right)\right)=\sigma_{\mathrm{ess}}\left(B_{1,1}\left(k\right)\right)=\sigma_{\mathrm{ess}}\left(M_{\mathfrak{f}\left(k\right)}+\mathrm{U}P_{0}\right)=\sigma_{\mathrm{ess}}\left(M_{\mathfrak{f}\left(k\right)}\right)\ .$
Corollary 4.2 (Bottom of the spectrum of $A_{1,1}(k)$ and $B_{1,1}(k)$)
For any $k\in\mathbb{T}^{2}$, one has that
$\min\sigma\left(A_{1,1}\left(k\right)\right)=\min\sigma\left(B_{1,1}\left(k\right)\right)=\min\sigma\left(M_{\mathfrak{f}\left(k\right)}\right)=4\epsilon-2\epsilon\cos\left(k/2\right)\doteq\mathfrak{z}\left(k\right)\ .$
Proof. Fix $k\in\mathbb{T}^{2}$. Since $\mathrm{U}\in\mathbb{R}_{0}^{+}$, one has the operator inequalities
$M_{\mathfrak{f}\left(k\right)}\leq M_{\mathfrak{f}\left(k\right)}+{\sum\limits_{x\in\mathbb{Z}^{2}}}\,\mathrm{u}\left(x\right)P_{x}\doteq B_{1,1}\left(k\right)\leq B_{1,1}\left(k\right)+\mathrm{U}P_{0}\doteq A_{1,1}(k)\ ,$
for the set of positive operators on a Hilbert space forms a norm-closed convex cone. By combining the last inequalities with Proposition 4.1, one arrives at the assertion.
4.2.2 Discrete spectrum
In the following it is technically convenient to assume that $\mathfrak{b}(k)$, which is the kinetic energy of a boson with quasi-momentum $k$, is below the bottom of the spectrum of $A_{1,1}(k)$, that is, the minimum energy of the fermion pair for the same total quasi-momentum. In other words, we assume from now on that
$\mathfrak{b}(k)\doteq h_{b}\epsilon\left(2-\cos\left(k\right)\right)\leq\mathfrak{z}(k)\doteq 4\epsilon-2\epsilon\cos(k/2)$ (63)
for all $k\in\mathbb{T}^{2}$, with equality only at $k=0$. See Equation (27) and Corollary 4.2. By direct computations101010(63) is clearly true for $h_{b}=0$. Take $h_{b}>0$. Using $\cos\left(\theta\right)=2\cos^{2}\left(\theta/2\right)-1$ and (30), one verifies that (63) is equivalent to $h_{b}\left(4-2x^{2}-2y^{2}\right)\leq 4-2\left(x+y\right)$ for $x,y\in\left[0,1\right]$. Since $\inf_{x\in\left[0,1\right]}\left\{h_{b}x^{2}-x\right\}=-1/(4h_{b})$ for $h_{b}\geq 1/2$ and $\inf_{x\in\left[0,1\right]}\left\{h_{b}x^{2}-x\right\}=h_{b}-1$ for $h_{b}\in(0,1/2]$, we deduce that (63) holds true iff $h_{b}\in[0,1/2]$., one verifies that this amounts to take $h_{b}$ in the interval $[0,1/2]$. This means that we consider a regime where the boson mass is at least the mass of the two fermions, as physically expected for cuprate superconductors, see [21, Section 3.1].
Proposition 4.3 (Eigenvalues of fiber Hamiltonians – I)
Take any $k\in\mathbb{T}^{2}$ and $h_{b}\in[0,1/2]$.
  1. i.)
    $\lambda\neq\mathfrak{b}(k)$ is an eigenvalue of $A(k)$ iff there is a non-zero vector $\varphi\in L^{2}(\mathbb{T}^{2})$ in the kernel of the bounded operator
    $A_{1,1}\left(k\right)-\lambda\mathfrak{1}-\left(\mathfrak{b}\left(k\right)-\lambda\right)^{-1}A_{1,2}\left(k\right)A_{2,1}\left(k\right)\in\mathcal{B}\left(L^{2}(\mathbb{T}^{2})\right)\text{ }.$
    In this case, $\lambda$ is an eigenvalue of $A(k)$ with associated eigenvector
    $\left(\varphi,-\left(\mathfrak{b}\left(k\right)-\lambda\right)^{-1}A_{2,1}\left(k\right)\varphi\right)\in\mathcal{H}\doteq L^{2}\left(\mathbb{T}^{2}\right)\oplus\mathbb{C}\text{ }.$
  2. ii.)
    $\mathfrak{b}(k)$ is an eigenvalue of $A(k)$ iff $\hat{\upsilon}(k)=0$.
Proof. Fix $k\in\mathbb{T}^{2}$ and $h_{b}\in[0,1/2]$. We start with the proof of Assertion (i): If $\lambda\neq\mathfrak{b}(k)$ is an eigenvalue of $A(k)$ with associated eigenvector $\0(\varphi,z)\in\mathcal{H}\backslash\{0\}$, then we directly deduce from (40) that
[…]
By combining these two equations, we obtain
$z=-\left(\mathfrak{b}\left(k\right)-\lambda\right)^{-1}A_{2,1}\left(k\right)\varphi$ (66)
and, thus,
$\left[A_{1,1}\left(k\right)-\lambda\mathfrak{1}-\left(\mathfrak{b}\left(k\right)-\lambda\right)^{-1}A_{1,2}\left(k\right)A_{2,1}\left(k\right)\right]\varphi=0\text{ }.$
We have that $\varphi\neq 0$, for otherwise $z$ would also be zero, by (66), and this would contradict the fact that $(\varphi,z)$ is a non-zero vector. The converse is obvious and Assertion (i) holds true.
We now prove Assertion (ii): It is easy to check from (40) that $\hat{\upsilon}(k)=0$ implies that
$A\left(k\right)\left(0,1\right)=\mathfrak{b}\left(k\right)\left(0,1\right)\text{ }.$ (67)
Conversely, suppose that $\mathfrak{b}(k)$ is an eigenvalue of $A(k)$ with associated eigenvector $\0(\varphi,z)\in\mathcal{H}\backslash\{0\}$, but $\hat{\upsilon}(k)\neq 0$. Then, by (35) and (40),
$\left(A_{1,1}\left(k\right)-\mathfrak{b}\left(k\right)\mathfrak{1}\right)\varphi+A_{1,2}\left(k\right)z=0\qquad\text{and}\qquad A_{2,1}\left(k\right)\varphi\doteq\hat{\upsilon}\left(k\right)\left\langle\mathfrak{d}\left(k\right),\varphi\right\rangle=0\text{ }.$ (68)
Remark that the second equality says that $\varphi\,\bot\,\mathfrak{d}(k)$, since we assume $\hat{\upsilon}(k)\neq 0$. Considering the scalar product of $\varphi$ with both sides of the first equation, we then get that
$\left\langle\varphi,\left(A_{1,1}\left(k\right)-\mathfrak{b}\left(k\right)\mathfrak{1}\right)\varphi\right\rangle+z\hat{\upsilon}\left(k\right)\left\langle\varphi,\mathfrak{d}\left(k\right)\right\rangle=\left\langle\varphi,\left(A_{1,1}\left(k\right)-\mathfrak{b}\left(k\right)\mathfrak{1}\right)\varphi\right\rangle=0\text{ },$ (69)
see (38). Because $h_{b}\in[0,1/2]$, if $k\neq 0$ then (63) holds true with a strict inequality and therefore,
$\mathfrak{b}\left(k\right)<4\epsilon-2\epsilon\cos\left(k/2\right)=\min\sigma\left(A_{1,1}\left(k\right)\right)\ ,$
thanks to Corollary 4.2. Hence,
$A_{1,1}\left(k\right)-\mathfrak{b}\left(k\right)\mathfrak{1}\geq c\mathfrak{1}\ ,$
for some constant $c>0$, which, combined with (69), in turn implies that $\varphi=0$. If now $k=0$, then $\mathfrak{b}(0)=0$ (see (27)) and we obtain from (69) that
${\int_{\mathbb{T}^{2}}}\left|\varphi\left(p\right)\right|^{2}\mathfrak{f}\left(0\right)\left(p\right)\,\nu\left(\mathrm{d}p\right)=\langle\varphi,M_{\mathfrak{f}\left(0\right)}\varphi\rangle\leq\langle\varphi,A_{1,1}\left(0\right)\varphi\rangle=0\text{ },$ (70)
since $M_{\mathfrak{f}\left(0\right)}\leq A_{1,1}(\mathrm{U},0)$ (see (31)–(32)). As
$\mathfrak{f}(0)(p)\doteq\epsilon\{4-2\cos(p)\}\ ,\qquad p\in\mathbb{T}^{2}\ ,$
(see (28) and (30)) defines a positive and continuous function that vanishes at $p=0$ only, one deduces from (70) that $\varphi=0$ also when $k=0$. Therefore, for any fixed $k\in\mathbb{T}^{2}$, $\varphi=0$ and so, (68) combined with (38) yields in this case
$A_{1,2}(k)z\doteq\hat{\upsilon}(k)\mathfrak{d}(k)z=0\text{ },$
i.e., $z=0$. Thus, we arrive at $(\varphi,z)=(0,0)$, which contradicts the fact that $(\varphi,z)$ is a non-zero vector. Therefore, if $\mathfrak{b}(k)$ is an eigenvalue of $A(k)$ then we must have $\hat{\upsilon}(k)=0$.
The Birman-Schwinger principle (Theorem 5.10) allows us to transform the eigenvalue problem for the fiber Hamiltonian $A(k)$ into a non-linear equation on the resolvent set $\rho(A_{1,1}(\mathrm{U},k))$ of the operator $A_{1,1}(\mathrm{U},k)$, which is the resolvent set for a fermion pair with total quasi-momentum $k\in\mathbb{T}^{2}$:
Theorem 4.4 (Characteristic equation for eigenvalues)
Fix $h_{b}\in[0,1/2]$ and $k\in\mathbb{T}^{2}$. Then, $\lambda\in\rho(A_{1,1}(k))$ is an eigenvalue of $A(k)$ iff it is a solution to the equation
$\hat{\upsilon}\left(k\right)^{2}\mathfrak{T}\left(k,z\right)+z-\mathfrak{b}\left(k\right)=0,\qquad z\in\rho\left(A_{1,1}\left(k\right)\right)\text{ },$ (71)
where $\mathfrak{T}$ is the function defined by (43), that is,
$\mathfrak{T}\left(k,z\right)\equiv\mathfrak{T}\left(\mathrm{U},k,z\right)\doteq\left\langle\mathfrak{d}\left(k\right),\left(A_{1,1}\left(k\right)-z\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right\rangle\ ,$ (72)
Proof. Fix $h_{b}\in[0,1/2]$ and $k\in\mathbb{T}^{2}$. We divide the proof in several cases:
Case 1: We first consider the case $\hat{\upsilon}(k)=0$ and $k\neq 0$. In that situation, $\mathfrak{b}(k)$ is trivially the only solution to (71). On the other hand, we already know from Proposition 4.3 (ii) that $\mathfrak{b}(k)$ is an eigenvalue of $A(k)$. We must therefore prove that there is no other eigenvalue $\lambda$ of $A(k)$ in $\rho(A_{1,1}(k))$ but $\mathfrak{b}(k)$. In fact, if such a $\lambda\in\rho(A_{1,1}(k))$ exists then, by Proposition 4.3 (i) with $\hat{\upsilon}(k)=0$, $A_{1,1}(k)-\lambda\mathfrak{1}$ would have a non-trivial kernel, which is not possible, for $\lambda$ is in the resolvent set of $A_{1,1}(k)$.
Case 2: Suppose that $k=0$ and $\hat{\upsilon}(0)=0$. We observe that (71) has no solution because
$0\in[0,8\epsilon]=\sigma_{\mathrm{ess}}(A_{1,1}(0))\ ,$
thanks to Proposition 4.1. In addition, by applying Proposition 4.3 (i) and noting that $\mathfrak{b}(0)=0$, we see that $A(0)$ has no eigenvalues in $\rho(A_{1,1}(0))$.
Case 3: Finally, assume that $\hat{\upsilon}(k)\neq 0$ and take $\lambda\in\rho(A_{1,1}(k))$. Observe from Proposition 4.3 (ii) that $\mathfrak{b}(k)$ cannot be an eigenvalue of $A(k)$. Additionally, $\mathfrak{b}(k)$ cannot be a solution to Equation (71). This last observation is proven as follows: When $k=0$ this is clear because $\mathfrak{b}(0)=0$ is not even in the domain of the equation to be solved in (71). For $k\neq 0$, if $\mathfrak{b}(k)$ is a solution to (71), then $\mathfrak{T}(k,\mathfrak{b}(k))=0$, but we know from Corollary 4.2 and $h_{b}\in[0,1/2]$ that
$A_{1,1}(k)-\mathfrak{b}(k)\mathfrak{1}\geq c\mathfrak{1}$
for some constant $c>0$. Therefore, $\mathfrak{T}(k,\mathfrak{b}(k))=0$ would yield $\mathfrak{d}(k)=0$, which is obviously wrong, by (29). Therefore, in all cases, $\mathfrak{b}(k)$ cannot be a solution to Equation (71) and we can assume that $\lambda\neq\mathfrak{b}(k)$. Now, the remaining part of the proof is essentially the same as the one of [22, Proposition 10], but we reproduce it for completeness. By (35)–(38), the orthogonal projection $S$ onto the subspace $\mathbb{C}\mathfrak{d}(k)\subseteq L^{2}(\mathbb{T}^{2})$ can be written as
$S\varphi=\|\mathfrak{d}(k)\|^{-2}\langle\mathfrak{d}(k),\varphi\rangle\mathfrak{d}(k)=\hat{\upsilon}(k)^{-2}\|\mathfrak{d}(k)\|^{-2}A_{1,2}(k)A_{2,1}(k)\varphi,\qquad\varphi\in L^{2}\left(\mathbb{T}^{2}\right)\ .$
Then, observe from Proposition 4.3 (i) that $\lambda$ is an eigenvalue of $A(k)$ iff $\lambda$ is an eigenvalue of $T-V^{2}$ with
$V$ $\doteq$ $\hat{\upsilon}(k)(\mathfrak{b}(k)-\lambda)^{-1/2}\|\mathfrak{d}(k)\|S\ ,$ (73)
$T$ $\doteq$ $A_{1,1}(k)\text{ },$ (74)
Thus, by applying Theorem 5.10, we deduce that $\lambda$ is an eigenvalue of $A(k)$ iff $1$ is an eigenvalue of the corresponding Birman-Schwinger operator, which, with the above operators $T$ and $V$, is equal to
$\mathrm{B}(\lambda)=\hat{\upsilon}(k)^{2}(\mathfrak{b}(k)-\lambda)^{-1}\|\mathfrak{d}(k)\|^{2}S(A_{1,1}(k)-\lambda\mathfrak{1})^{-1}S\text{ }.$
Remark in this case that
$\mathcal{E}_{\mathrm{B}(\lambda)}(1)=\mathbb{C}\mathfrak{d}(k)\qquad\text{and}\qquad\dim\mathcal{E}_{T-V^{2}}(\lambda)=\dim\mathcal{E}_{\mathrm{B}(\lambda)}(1)=1\text{ },$ (75)
since, obviously,
$\mathrm{B}(\lambda)L^{2}\left(\mathbb{T}^{2}\right)\subseteq SL^{2}\left(\mathbb{T}^{2}\right)=\mathbb{C}\mathfrak{d}(k)\text{ }.$
We thus conclude that $\lambda$ is an eigenvalue of $A(k)$ iff
$\mathrm{B}(\lambda)\mathfrak{d}(k)=\mathfrak{d}(k)$ $\Leftrightarrow\langle\mathfrak{d}(k),\mathrm{B}(\lambda)\mathfrak{d}(k)-\mathfrak{d}(k)\rangle=0$
$\Leftrightarrow\langle\mathfrak{d}(k),\mathrm{B}(\lambda)\mathfrak{d}(k)\rangle=\|\mathfrak{d}(k)\|^{2}$
$\Leftrightarrow\hat{\upsilon}(k)^{2}(\mathfrak{b}(k)-\lambda)^{-1}\|\mathfrak{d}(k)\|^{2}\langle\mathfrak{d}(k),S(A_{1,1}(k)-\lambda\mathfrak{1})^{-1}S\mathfrak{d}(k)\rangle=\|\mathfrak{d}(k)\|^{2}$
$\Leftrightarrow\hat{\upsilon}(k)^{2}\mathfrak{T}(k,\lambda)=\mathfrak{b}(k)-\lambda\text{ }.$
This completes the proof.
Corollary 4.5 (Eigenspaces of fiber Hamiltonians)
Fix $h_{b}\in[0,1/2]$ and $k\in\mathbb{T}^{2}$. If $\lambda\in\rho(A_{1,1}(k))$ is an eigenvalue of the fiber Hamiltonian $A(k)$ then the associated eigenspace is
$\mathcal{E}_{A\left(k\right)}\left(\lambda\right)=\mathbb{C}g\left(k,\lambda\right)\text{ },$
where
$g\left(k,\lambda\right)\doteq\left(\hat{\upsilon}\left(k\right)\left(A_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right),-1\right)\in\mathcal{H}\ .$
In particular, $\lambda$ is a non-degenerated eigenvalue of $A(k)$.
Proof. Assume that $\hat{\upsilon}(k)\neq 0$. Recall from the proof of Theorem 4.4 that in this case $\mathfrak{b}(k)$ is not an eigenvalue of $A(k)$ and so, we assume without loss of generality that $\lambda\neq\mathfrak{b}(k)$. In this case, observe that we have (75) and a close look at the proof of Theorem 5.10, in particular Lemma 5.9, leads us to
$\mathcal{E}_{T-V^{2}}(\lambda)\doteq\ker(T-V^{2}-\lambda\mathfrak{1})=\mathbb{C}\varphi_{0}\ ,$
where
$\varphi_{0}\doteq(T-\lambda\mathfrak{1})^{-1}V\mathfrak{d}(k)=\hat{\upsilon}(k)(\mathfrak{b}(k)-\lambda)^{-1/2}\|\mathfrak{d}(k)\|(A_{1,1}(k)-\lambda\mathfrak{1})^{-1}\mathfrak{d}(k)\ ,$ (76)
by (73) and (74). From Proposition 4.3 (i) ($\hat{\upsilon}(k)\neq 0$), one then obtains that
$\mathcal{E}_{A(k)}(\lambda)$ $=$ $\left\{(\varphi,-(\mathfrak{b}(k)-\lambda)^{-1}A_{2,1}(k)\varphi)\in\mathcal{H}\,:\,\varphi\in\ker(T-V^{2}-\lambda\mathfrak{1})\right\}$
$=$ $\mathbb{C}\ (\varphi_{0},-(\mathfrak{b}(k)-\lambda)^{-1}A_{2,1}(k)\varphi_{0})\ .$
In view of Equation (35), (76) and Theorem 4.4, the last vector can be rewritten as follows:
$(\varphi_{0},-(\mathfrak{b}(k)-\lambda)^{-1}A_{2,1}(k)\varphi_{0})$ $=(\varphi_{0},-(\mathfrak{b}(k)-\lambda)^{-1}\hat{\upsilon}(k)\langle\mathfrak{d}(k),\varphi_{0}\rangle)$
$=(\varphi_{0},-(\mathfrak{b}(k)-\lambda)^{-3/2}\hat{\upsilon}(k)^{2}\|\mathfrak{d}(k)\|\mathfrak{T}(k,\lambda))$
$=(\varphi_{0},-(\mathfrak{b}(k)-\lambda)^{-1/2}\|\mathfrak{d}(k)\|)$
$=(\mathfrak{b}(k)-\lambda)^{-1/2}\|\mathfrak{d}(k)\|g(k,\lambda)\ ,$
whenever $\hat{\upsilon}(k)\neq 0$. Finally, if $\hat{\upsilon}(k)=0$ and $\lambda\in\rho(A_{1,1}(k))$ is an eigenvalue of $A(k)$ then it is straightforward to check that $\lambda=\mathfrak{b}(k)$ with eigenspace generated by the vector $(0,1)\in\mathcal{H}$, see for instance (64)–(65) and (67).
Corollary 4.6 (Eigenvalues of fiber Hamiltonians – II)
Fix $h_{b}\in[0,1/2]$ and $k\in\mathbb{T}^{2}$. There is at most one eigenvalue of $A(k)$ in each connected component of $\rho(A_{1,1}(k))\cap\mathbb{R}$.
Proof. In view of Theorem 4.4, it suffices to show that the derivative of the mapping
$\rho(A_{1,1}(k))\cap\mathbb{R}\ni x\longmapsto\hat{\upsilon}(k)^{2}\mathfrak{T}(k,x)+x-\mathfrak{b}(k)\in\mathbb{R}$
is strictly positive. For any $x_{0}\in\rho(A_{1,1}(k))\cap\mathbb{R}$, we have that
$\left.\partial_{x}\left\{\hat{\upsilon}(k)^{2}\mathfrak{T}(k,x)+x-\mathfrak{b}(k)\right\}\right|_{x=x_{0}}=\hat{\upsilon}(k)^{2}\|(A_{1,1}(k)-x_{0}\mathfrak{1})^{-1}\mathfrak{d}(k)\|^{2}+1>1\ .$ (77)
4.2.3 Bottom of the spectrum
As is well-known, physical properties of quantum systems at very low temperatures are essentially determined by the bottom of the spectrum of the corresponding Hamiltonian. In our case, having in mind the application to superconductivity in cuprates, we would like to study the bottom of the spectrum of the Hamiltonian $H\in\mathcal{B}(\mathfrak{H})$ defined by (19). By Proposition 2.1 and Theorem 5.3, we thus study the bottom of the spectrum of the fiber Hamiltonian $A(k)$ (40) at fixed total quasi-momentum $k\in\mathbb{T}^{2}$, similar to [22, 21].
Theorem 4.7 (Bottom of the spectrum of $A(k)$)
Fix $h_{b}\in[0,1/2]$. If $k\neq 0$ then there is exactly one eigenvalue $\mathrm{E}\left(k\right)\equiv\mathrm{E}\left(\mathrm{U},k\right)$ of the Hamiltonian $A(k)$ strictly below $\sigma_{\mathrm{ess}}\left(A\left(k\right)\right)$. In this case, the eigenvalue is non-degenerated and $\mathrm{E}(k)<\mathfrak{b}(k)$ when $\hat{\upsilon}(k)\neq 0$, whereas $\mathrm{E}(k)=\mathfrak{b}(k)$ if $\hat{\upsilon}(k)=0$. This statement remains valid for $k=0$ provided $\hat{\upsilon}(0)\neq 0$.
Proof. Assume that $k\neq 0$. Recall that $\mathfrak{z}(k)$ is defined in Corollary 4.2. From Corollaries 4.2 and 4.6, the interval $(-\infty,\mathfrak{z}(k))$ contains at most one eigenvalue of $A(k)$. By Corollary 4.5, the eigenvalue is non-degenerate, if it exists. If $\hat{\upsilon}(k)=0$, we know from Proposition 4.3 (ii) that $\mathfrak{b}(k)$ is such an eigenvalue. Recall that for $k\neq 0$ one has that $\mathfrak{b}(k)<\mathfrak{z}(k)$, because $h_{b}\in[0,1/2]$. Now, suppose that $\hat{\upsilon}(k)\neq 0$. When $\mathfrak{b}(k)\leq x<\mathfrak{z}(k)$, we have
$\left(A_{1,1}\left(k\right)-x\mathfrak{1}\right)^{-1}\geq c\mathfrak{1}$
for some constant $c>0$ and hence, $\mathfrak{T}(k,x)>0$, see (72). Consequently,
$\hat{\upsilon}(k)^{2}\mathfrak{T}(k,x)+x-\mathfrak{b}(k)>0\ ,$
which means that $A(k)$ has no eigenvalues in the interval $[\mathfrak{b}(k),\mathfrak{z}(k))$, by Theorem 4.4. We shall now look for an eigenvalue in the interval $(-\infty,\mathfrak{b}(k))$. On the one hand, using Corollary 4.2, observe that
$\|\mathfrak{d}(k)\|^{-2}|\mathfrak{T}(k,x)|\leq\|(A_{1,1}(k)-x\mathfrak{1})^{-1}\|_{\mathrm{o}\mathrm{p}}=(\mathfrak{z}(k)-x)^{-1}\leq(\mathfrak{b}(k)-x)^{-1},$
whenever $x<\mathfrak{b}(k)$. Taking $x\rightarrow-\infty$, $\mathfrak{T}(k,x)$ tends to zero and, hence,
${\lim\limits_{x\rightarrow-\infty}}\left\{\hat{\upsilon}(k)^{2}\mathfrak{T}(k,x)+x-\mathfrak{b}(k)\right\}=-\infty\text{ }.$
On the other hand, the continuity of the mapping
$\mathfrak{T}(k,\cdot):\rho(A_{1,1}(k))\rightarrow\mathbb{R}$
on $(-\infty,\mathfrak{b}(k)]$ gives us
${\lim\limits_{x\rightarrow\mathfrak{b}(k)}}\left\{\hat{\upsilon}(k)^{2}\mathfrak{T}(k,x)+x-\mathfrak{b}(k)\right\}=\hat{\upsilon}(k)^{2}\mathfrak{T}(k,\mathfrak{b}(k))>0\text{ }.$
By the intermediate value theorem, there is $\mathrm{E}(k)\in(-\infty,\mathfrak{b}(k))$ such that
$\hat{\upsilon}(k)^{2}\mathfrak{T}(k,\mathrm{E}(k))+\mathrm{E}(k)-\mathfrak{b}(k)=0\text{ }.$
By Theorem 4.4, $\mathrm{E}(k)$ must be an eigenvalue of $A(k)$.
The proof for $k=0$ is done in a similar way. Basically, the only difference is that, in this case, $\mathfrak{b}(0)=\mathfrak{z}(0)=0$ and
${\lim\limits_{x\rightarrow 0^{-}}}\left\{\hat{\upsilon}(0)^{2}\mathfrak{T}(0,x)+x\right\}\in(0,\infty]$
occurs due to other reasons. Indeed, from Corollary 4.2 we deduce that $\mathfrak{T}(0,\cdot)$ is strictly positive on the interval $(-\infty,0)$. Because of (77) we also have that $\partial_{x}\mathfrak{T}(0,x)|_{x=x_{0}}\geq 0$ whenever $x_{0}<0$. Thus the limit of $\mathfrak{T}(0,x)$ as $x\to 0^{-}$ exists, being possibly infinite.
If $\hat{\upsilon}(0)=0$ then, by Theorem 4.4, the fiber Hamiltonian $A(0)$ has no negative eigenvalues. In this case we set $\mathrm{E}(0)=0$, which is obviously an eigenvalue of $A(0)$ with associated eigenvector $(0,1)$. (Note that $\sigma_{\mathrm{ess}}(A(0))=[0,8\epsilon]$, by Proposition 4.1.) With this definition, observe that, for all $k\in\mathbb{T}^{2}$, $\mathrm{E}(k)$ is the minimum spectral value of $A(k)$:
$\mathrm{E}\left(k\right)=\min\sigma\left(A\left(k\right)\right)\leq\mathfrak{z}\left(k\right)=\min\sigma_{\mathrm{ess}}\left(A\left(k\right)\right)\ .$ (78)
The lowest eigenvalue $\mathrm{E}(k)$ of $A(k)$, when $\mathrm{E}(k)<0$, is related to the formation of dressed bound fermion pairs of fermions with total quasi-momentum $k\in\mathbb{T}^{2}$. In Theorem 4.19, we make this claim more precise and prove the spatial localization of such bounded pairs. Before doing that we study the regularity of the real-valued function
$\mathrm{E}\equiv\mathrm{E}\left(\mathrm{U},\cdot\right):\mathbb{T}^{2}\rightarrow\mathbb{R}$
on the two-dimensional torus.
To this end, we rewrite the characteristic equation given by Theorem 4.4 via the function $\Phi:\mathcal{O}\rightarrow\mathbb{R}$ defined by
$\Phi\left(k,x\right)\equiv\Phi\left(\mathrm{U},k,x\right)\doteq\hat{\upsilon}\left(k\right)^{2}\mathfrak{T}\left(k,x\right)+x-\mathfrak{b}\left(k\right)\ ,$ (79)
where $\mathcal{O}$ is the open set
$\mathcal{O}\doteq\left\{\left(k,x\right)\in\mathbb{S}^{2}\times\mathbb{R}:x<\mathfrak{z}\left(k\right)\right\}\subseteq\mathbb{R}^{3}\ .$ (80)
Observe from Equation (77) that $\partial_{x}\Phi>0$ over the whole domain of $\Phi$.
We now study the continuity of the function $\mathrm{E}:\mathbb{T}^{2}\rightarrow\mathbb{R}$ and give a sufficient condition for $\mathrm{E}$ to be of class $C^{d}$ on $\mathbb{S}^{2}$ for every $d\in\mathbb{N}\cup\{\omega,a\}$, where $C^{d}(\Omega)$, $d\in\mathbb{N}$, stands for the space of $d$times continuously differentiable functions on $\Omega$, while $C^{\omega}(\Omega)$ and $C^{a}(\Omega)$ refer to the space of smooth and real analytic functions on $\Omega$, respectively.
Theorem 4.8 (Regularity of the function $\mathrm{E}$)
Let $h_{b}\in[0,1/2]$.
  1. i.)
    The family $\{\mathrm{E}(\mathrm{U},\cdot)\}_{\mathrm{U}\in\mathbb{R}_{0}^{+}}$ of real-valued functions on $\mathbb{T}^{2}$ is equicontinuous with respect to the metric [^11] Note that as $d_{\mathbb{T}^{2}}$ is smaller than the Euclidean metric for $\mathbb{T}^{2}$ as a subset of $\mathbb{R}^{2}$, the equicontinuity also holds true for the Euclidean metric. $d_{\mathbb{T}^{2}}$.
  2. ii.)
    If $\hat{\upsilon}\in C^{d}(\mathbb{S}^{2})$ ($\mathbb{S}^{2}\subseteq\mathbb{R}^{2}$) for some $d\in\mathbb{N}\cup\{\omega,a\}$, then $\mathrm{E}\equiv\mathrm{E}(\mathrm{U},\cdot)\in C^{d}(\mathbb{S}^{2})$ for all $\mathrm{U}\in\mathbb{R}_{0}^{+}$. In this case,
    $\partial_{k_{j}}\mathrm{E}\left(k\right)=-\left(\partial_{x}\Phi\left({k,\mathrm{E}\left(k\right)}\right)\right)^{-1}\partial_{k_{j}}\Phi\left({k,\mathrm{E}\left(k\right)}\right)\ ,\qquad k\in\mathbb{S}^{2},\ j\in\{1,2\}\ .$ (81)
Proof. By the spectral theorem, we deduce from (78) that, for any $\mathrm{U}\in\mathbb{R}_{0}^{+}$,
$\mathrm{E}\left(\mathrm{U},k\right)=\min\sigma\left(A\left(\mathrm{U},k\right)\right)={\inf_{\psi\in\mathcal{H},\|\psi\|=1}}\left\langle\psi,A\left(\mathrm{U},k\right)\psi\right\rangle,\qquad k\in\mathbb{T}^{2}\ .$ (82)
Given any $\varepsilon>0$ and $k_{0}\in\mathbb{T}^{2}$, by the (operator norm) continuity of the mapping $A\left(0,\cdot\right):\mathbb{T}^{2}\rightarrow\mathcal{B}(\mathcal{H})$ at the point $k_{0}$ we can find $\delta>0$ such that
$\sup_{\mathrm{U}\in\mathbb{R}_{0}^{+}}{\sup_{\psi\in\mathcal{H},\|\psi\|=1}}\left|\left\langle\psi,A\left(\mathrm{U},k\right)\psi\right\rangle-\left\langle\psi,A\left(\mathrm{U},k_{0}\right)\psi\right\rangle\right|$ $\leq\sup_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\left\|A\left(\mathrm{U},k\right)-A\left(\mathrm{U},k_{0}\right)\right\|_{\mathrm{op}}$
$=\sup_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\left\|A\left(0,k\right)-A\left(0,k_{0}\right)\right\|_{\mathrm{op}}<\varepsilon$
for every $k\in\mathbb{T}^{2}$ with $d_{\mathbb{T}^{2}}(k,k_{0})<\delta$. Recall that $\mathcal{H}$ stands for the (fiber) Hilbert space $L^{2}(\mathbb{T}^{2})\oplus\mathbb{C}$, see (23), while $d_{\mathbb{T}^{2}}$ is the metric (20) on the torus $\mathbb{T}^{2}$. Therefore, $\mathrm{E}:\mathbb{T}^{2}\rightarrow\mathbb{R}$ can be expressed as the infimum over the equicontinuous family $\{\langle\psi,A(\mathrm{U},\cdot)\psi\rangle\}_{\mathrm{U}\in\mathbb{R}_{0}^{+},\psi\in\mathcal{H},\|\psi\|=1}$ of (continuous) functions and Assertion (i) follows.
Take again some $k_{0}\in\mathbb{S}^{2}$. Assume that $\hat{\upsilon}$ is of class $C^{d}$ on $\mathbb{S}^{2}\subseteq\mathbb{R}^{2}$ with $d\in\mathbb{N}\cup\{\omega,a\}$. Let $\vartheta=(k_{0},\mathrm{E}(k_{0}))\in\mathcal{O}$. Using Theorems 4.4 and 4.7 as well as Equation (77) and the (operator norm) continuity of the mapping
$A_{1,1}\left(\cdot\right):\mathbb{T}^{2}\rightarrow\mathcal{B}\left(L^{2}(\mathbb{T}^{2})\right)\ ,$ (83)
one checks that $\Phi\in C^{d}(\mathcal{O})$ with $d\geq 1$, $\Phi(\vartheta)=0$ and $\partial_{x}\Phi(\vartheta)\neq 0$. See for instance (77). We can thus apply the implicit function theorem (see, e.g., [45] for an ordinary version and [46] for an analytic version) to obtain open subsets $U\subseteq\mathbb{R}^{2}$ and $J\subseteq\mathbb{R}$ such that $\vartheta\in U\times J\subseteq\mathcal{O}$ and, for each $k\in U$, there is a unique real number $\xi(k)\in J$ satisfying $\Phi(k,\xi(k))=0$. Moreover, the mapping $\xi:U\rightarrow J$ defined in this way is of class $C^{d}$ and its partial derivatives are given by
${\partial}_{k_{j}}{\xi\left({k}\right)}=-\left(\partial_{x}\Phi{\left({k,\xi\left(k\right)}\right)}\right)^{-1}{{\partial}_{k_{j}}\Phi\left({k,\xi\left(k\right)}\right)}\ ,\qquad k\in U,\ j\in\{1,2\}\ .$ (84)
As $\xi(k_{0})=\mathrm{E}(k_{0})<\mathfrak{z}(k_{0})$ (see (78)), by continuity there exists a neighborhood $V\subseteq U$ of $k_{0}$ such that $\xi(k)<\mathfrak{z}(k)$ for every $k\in V$. It follows that, for all $k\in V$, $\xi(k)$ and $\mathrm{E}(k)$ are in the same connected component $(-\infty,\mathfrak{z}(k))$ and from $\partial_{x}\Phi>0$ we conclude that $\mathrm{E}\upharpoonright V=\xi\upharpoonright V$. So, $\mathrm{E}$ is of class $C^{d}$ near $k_{0}$ and (84) yields (81) for any $k\in V$, in particular for $k=k_{0}$. As $k_{0}$ is arbitrary, Assertion (ii) follows.
We can now deduce from Theorem 4.8 that $\mathrm{E}$ is a dispersion relation (see Definition 3.7) when the function $\hat{\upsilon}:\mathbb{S}^{2}\rightarrow\mathbb{R}$ is at least $2$ times continuously differentiable and in this case, we can even compute the group velocity. To see this, recall that, for any $f\in C^{2}(\mathbb{S}^{2})$, we define in (46) the subset
$\mathfrak{M}_{f}\doteq\left\{k\in\mathbb{S}^{2}\,:\,\mathrm{Hess}\left(f\right)\left(k\right)\in\mathsf{GL}_{2}\left(\mathbb{R}\right)\right\}\subseteq\mathbb{S}^{2}$
with $\mathsf{GL}_{2}\left(\mathbb{R}\right)$ being the set of invertible $2\times 2$ matrices with real coefficients.
Corollary 4.9 ($\mathrm{E}$ as a dispersion relation and group velocity)
Let $h_{b}\in[0,1/2]$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$. Then, $\mathrm{E}\equiv\mathrm{E}(\mathrm{U},\cdot)\in C(\mathbb{T}^{2})$ and is of class $C^{2}$ on the open set $\mathbb{S}^{2}\subseteq\mathbb{R}^{2}$ whenever $\hat{\upsilon}$ is of class $C^{2}$ on $\mathbb{S}^{2}$. In this case, the corresponding group velocity is
$\mathbf{v}_{\mathrm{E}}\left(k\right)\doteq\vec{\nabla}_{k}\mathrm{E}\left(k\right)=-\left(\partial_{x}\Phi\left({k,\mathrm{E}\left(k\right)}\right)\right)^{-1}\vec{\nabla}_{k}\Phi\left({k,\mathrm{E}\left(k\right)}\right)\ ,\qquad k\in\mathbb{S}^{2}\ .$
Moreover, if $\hat{\upsilon}$ is real analytic (i.e., of class $C^{a}$ in the above terminology) on $\mathbb{S}^{2}$, then either $\mathfrak{M}_{\mathrm{E}}$ has full measure or is empty.
Proof. The first part of the assertion is a direct application of Theorem 4.8. It remains to study the set $\mathfrak{M}_{\mathrm{E}}$. If $\hat{\upsilon}$ is real analytic then, from Equation (45) and Theorem 4.8, the function $f:\mathbb{S}^{2}\longrightarrow\mathbb{R}$ defined by
$f(k)\doteq\det\left(\mathrm{Hess}{\left(\mathrm{E}\right)\left(k\right)}\right)\ ,\qquad k\in\mathbb{S}^{2}\ ,$
is real analytic and satisfies $f^{-1}(\{0\})=\mathbb{S}^{2}\backslash\mathfrak{M}_{\mathrm{E}}$. Since the zeros of any non-constant real analytic function have null Lebesgue measure (see, e.g., [43]), either $\mathfrak{M}_{\mathrm{E}}$ has full measure or is empty.
It is natural to derive now the ground state energy of the Hamiltonian $H\in\mathcal{B}(\mathfrak{H})$ defined by (19), which is related to the ground state energy of fiber Hamiltonians, thanks to Proposition 2.1 and Theorem 5.3. As expected, one has the following equality for the ground state energy
$E{\left(\mathrm{U}\right)}\doteq\min\sigma\left(H\right)=\min\mathrm{E}\left(\mathbb{T}^{2}\right)\text{ }.$
A proof can be done like [22, Lemma 8] by using Kato’s perturbation theory. In the sequel, we provide an alternative way of proving the equality, which is much more direct.
Proposition 4.10 (Bottom of the spectrum of $H$)
We have that
$E{\left(\mathrm{U}\right)}={\min\limits_{k\in\mathbb{T}^{2}}}\,\min\sigma\left(A\left(k\right)\right)=\min\mathrm{E}\left(\mathbb{T}^{2}\right)\leq 0\ .$
Proof. We first remark that the union
$\mathcal{K}\doteq{\bigcup}\left\{\sigma\left(A\left(k\right)\right)\,:\,k\in\mathbb{T}^{2}\right\}\subseteq\mathbb{R}$
of the spectra of all fiber Hamiltonians is closed. To see this, let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence of real numbers converging to $\lambda\in\mathbb{R}$ with $\lambda_{n}\in\sigma(A(k_{n}))$ for some $k_{n}\in\mathbb{T}^{2}$ at $n\in\mathbb{N}$. By compactness of $\mathbb{T}^{2}$, we can assume without loss of generality that $(k_{n})_{n\in\mathbb{N}}$ converges to some point $k_{0}\in\mathbb{T}^{2}$. By the (operator norm) continuity of the mapping $A:\mathbb{T}^{2}\rightarrow\mathcal{B}(\mathcal{H})$, it follows that
${\lim\limits_{n\rightarrow\infty}}\,(A(k_{n})-\lambda_{n}\mathfrak{1)}=A(k_{0})-\lambda\mathfrak{1}$
(in operator norm). Hence, $\lambda\in\sigma(A(k_{0}))$, for otherwise $A(k_{n})-\lambda_{n}\mathfrak{1}$ would be invertible [^12] Recall that the set of invertible operators on a Banach space $X$ is an open subset of $\mathcal{B}(X)$, with respect to the operator norm. See, e.g., [44, Theorem 1.4]. for sufficiently large $n$. Thus, $\mathcal{K}$ is a closed set and, as a consequence, for any $s\notin\mathcal{K}$, there is $\varepsilon>0$ such that
$\left(s-\varepsilon,s+\varepsilon\right)\cap\sigma\left(A\left(k\right)\right)=\emptyset\text{ },\qquad k\in\mathbb{T}^{2}\ .$
With this property, we infer from Proposition 2.1 and Theorem 5.3 that $\sigma(H)\subseteq\mathcal{K}$, which yields the inequality
$E{\left(\mathrm{U}\right)}\geq{\min\limits_{k\in\mathbb{T}^{2}}}\,\min\sigma\left(A\left(k\right)\right)=\min\mathrm{E}\left(\mathbb{T}^{2}\right)\text{ }.$ (85)
Note that the last equality results from (78). By Theorem 4.8 (i), $\mathrm{E}:\mathbb{T}^{2}\rightarrow\mathbb{R}$ is continuous. From the compactness of the torus $\mathbb{T}^{2}$ and the Weierstrass extreme value theorem, $\mathrm{E}$ has a minimizer in $\mathbb{T}^{2}$, say $k_{0}\in\mathbb{T}^{2}$. The continuity of $\mathrm{E}$ at $k_{0}$ implies that, for every $\varepsilon>0$, there is $\delta>0$ such that, for all $k\in\mathbb{T}^{2}$ satisfying $d_{\mathbb{T}^{2}}(k,k_{0})<\delta$,
$\mathrm{E}\left(k\right)\in\left(\mathrm{E}\left(k_{0}\right)-\varepsilon,\mathrm{E}\left(k_{0}\right)+\varepsilon\right)\ ,$
and, as a consequence,
$\nu\left(\left\{k\in\mathbb{T}^{2}:\sigma\left(A\left(k\right)\right)\cap\left(\mathrm{E}\left(k_{0}\right)-\varepsilon,\mathrm{E}\left(k_{0}\right)+\varepsilon\right)\neq\emptyset\right\}\right)\geq\nu\left(\mathbb{T}^{2}\cap B_{\delta}\left(k_{0}\right)\right)>0\text{ },$
where $B_{\delta}(k_{0})$ is the open ball (for the metric $d_{\mathbb{T}^{2}}$) centered at $k_{0}\in\mathbb{T}^{2}$ of radius $\delta\in\mathbb{R}^{+}$. By Theorem 5.3, this implies that $\mathrm{E}(k_{0})\in\sigma(H)$, which, combined with (85), yields the equalities
$E{\left(\mathrm{U}\right)}={\min\limits_{k\in\mathbb{T}^{2}}}\,\min\sigma\left(A\left(k\right)\right)=\min\mathrm{E}\left(\mathbb{T}^{2}\right)\text{ }.$
Using Equations (63) and (78), we note that
$E{\left(\mathrm{U}\right)}=\min\mathrm{E}\left(\mathbb{T}^{2}\right)\leq{\min\limits_{k\in\mathbb{T}^{2}}\ }\mathfrak{z}(k)=4\epsilon-2\epsilon\max_{k\in\mathbb{T}^{2}}\cos(k/2)=\mathfrak{z}(0)=0\ .$
4.3 Spectral Properties in the Hard-Core Limit
We now study the spectral properties of fiber Hamiltonians $A(k)$ (40) in the hard-core limit. It refers to the limit $\mathrm{U}\rightarrow\infty$. In fact, a very strong on-site repulsion $\mathrm{U}$ (see Equation (6)) prevents two fermions of opposite spins from occupying the same lattice site. We study in particular the continuous function $\mathrm{E}:\mathbb{T}^{2}\rightarrow\mathbb{R}$ defined by Theorem 4.7, which corresponds to the continuous family of non-degenerate eigenvalues at lowest energies in each fiber, in this limit.
An important result in this context is the characterization of such eigenvalues via the Birman-Schwinger principle, given by Theorem 4.4. In particular, we need first to study the hard-core limit of the characteristic equation, which amounts to determine the limit $\mathrm{U}\rightarrow\infty$ of the quantity (72), i.e.,
$\mathfrak{T}\left(\mathrm{U},k,\lambda\right)\equiv\mathfrak{T}\left(k,\lambda\right)\doteq\left\langle\mathfrak{d}\left(k\right),\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right\rangle\text{ },\qquad\lambda\in\rho\left(A_{1,1}\left(\mathrm{U},k\right)\right)\ ,$ (86)
for $h_{b}\in[0,1/2]$ and $k\in\mathbb{T}^{2}$. We start with this point, which allows us to study afterwards the limit of the lowest eigenvalues and all the derived quantities, like for instance the group velocity.
4.3.1 The characteristic equation in the hard-core limit
Let $\mathfrak{s}=\mathfrak{\hat{e}}_{0}\in L^{2}(\mathbb{T}^{2})$ denote the constant function 11\ 11 on the torus $\mathbb{T}^{2}$. For any fixed $k\in\mathbb{T}^{2}$ and $\lambda\in\rho(B_{1,1}(k))$, define the following four constants:
$R_{\mathfrak{s},\mathfrak{s}}$ $\equiv R_{\mathfrak{s},\mathfrak{s}}\left(k,\lambda\right)\doteq\left\langle\mathfrak{s},\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{s}\right\rangle\ ,$ (87)
$R_{\mathfrak{s},\mathfrak{d}}$ $\equiv R_{\mathfrak{s},\mathfrak{d}}\left(k,\lambda\right)\doteq\left\langle\mathfrak{s},\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right\rangle\ ,$ (88)
$R_{\mathfrak{d},\mathfrak{s}}$ $\equiv R_{\mathfrak{d},\mathfrak{s}}\left(k,\lambda\right)\doteq\left\langle\mathfrak{d}\left(k\right),\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{s}\right\rangle\ ,$ (89)
$R_{\mathfrak{d},\mathfrak{d}}$ $\equiv R_{\mathfrak{d},\mathfrak{d}}\left(k,\lambda\right)\doteq\left\langle\mathfrak{d}\left(k\right),\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right\rangle\ ,$ (90)
where we recall that $B_{1,1}\left(k\right)$ is defined by (31). In the following lemma we write $\mathfrak{T}(\mathrm{U},k,\lambda)$, which is defined by (86), in terms of these four quantities.
In the following it is technically convenient to assume that $\mathfrak{d}(k)\notin\mathbb{C}\mathfrak{s}$ for all $k\in\mathbb{T}^{2}$. Notice that this holds true iff $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$, i.e., $r_{\mathfrak{p}}>0$. See discussions around Equation (13).
Lemma 4.11
Let $k\in\mathbb{T}^{2}$, $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and $\lambda<\mathfrak{z}(k)$, with $\mathfrak{z}(k)\in\mathbb{R}$ defined in Corollary 4.2. Then,
$\mathfrak{T}\left(\mathrm{U},k,\lambda\right)=\frac{R_{\mathfrak{d},\mathfrak{d}}}{\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1}+\mathrm{U}\frac{R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}-\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}}{\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1}\text{ },$
with $R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}-\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}\geq 0$. Moreover, if $r_{\mathfrak{p}}>0$ (see (13)), then the inequality is strict.
Proof. The proof of the first part is a slightly more complicated version of the one of [22, Lemma 14]. Fix $k\in\mathbb{T}^{2}$, $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and $\lambda<\mathfrak{z}(k)$. Define the complex numbers
$Q_{\mathfrak{s},\mathfrak{s}}$ $\doteq$ $\left\langle\mathfrak{s},\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{s}\right\rangle\ ,$
$Q_{\mathfrak{s},\mathfrak{d}}$ $\doteq$ $\left\langle\mathfrak{s},\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right\rangle\ ,$
$Q_{\mathfrak{d},\mathfrak{s}}$ $\doteq$ $\left\langle\mathfrak{d}\left(k\right),\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{s}\right\rangle\ ,$
$Q_{\mathfrak{d},\mathfrak{d}}$ $\doteq$ $\left\langle\mathfrak{d}\left(k\right),\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right\rangle=\mathfrak{T}\left(\mathrm{U},k,\lambda\right)\ ,$
where we recall that
$A_{1,1}\left(\mathrm{U},k\right)\doteq B_{1,1}\left(k\right)+\mathrm{U}P_{0}\geq B_{1,1}\left(k\right)\ ,$ (91)
by Equation (32). Using Corollary 4.2 note at this point that
$\mathfrak{z}\left(k\right)=\min\sigma\left(A_{1,1}\left(\mathrm{U},k\right)\right)=\min\sigma\left(B_{1,1}\left(k\right)\right)\geq 0\ .$ (92)
In particular,
$\lambda\in\left(-\infty,\mathfrak{z}\left(k\right)\right)\subseteq\rho\left(A_{1,1}\left(\mathrm{U},k\right)\right)\cap\rho\left(B_{1,1}\left(k\right)\right)$ (93)
and the resolvent operators $(A_{1,1}(\mathrm{U},k)-\lambda\mathfrak{1})^{-1}$ and $(B_{1,1}(k)-\lambda\mathfrak{1})^{-1}$ are strictly positive. By using the second resolvent identity together with Equation (91), we compute that
$\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}=\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}-\mathrm{U}\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}P_{0}\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\ .$
Recalling that $P_{0}$ is the orthogonal projection onto $\mathbb{C}\mathfrak{s=\mathbb{C}\hat{e}}_{0}$, we have that
$Q_{\mathfrak{s},\mathfrak{s}}$ $=R_{\mathfrak{s},\mathfrak{s}}-\mathrm{U}\left\langle\mathfrak{s},\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}P_{0}\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{s}\right\rangle$
$=R_{\mathfrak{s},\mathfrak{s}}-\mathrm{U}Q_{\mathfrak{s},\mathfrak{s}}R_{\mathfrak{s},\mathfrak{s}}\ ,$
$Q_{\mathfrak{d},\mathfrak{d}}$ $=R_{\mathfrak{d},\mathfrak{d}}-\mathrm{U}\left\langle\mathfrak{d}\left(k\right),\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}P_{0}\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right\rangle$
$=R_{\mathfrak{d},\mathfrak{d}}-\mathrm{U}Q_{\mathfrak{d},\mathfrak{s}}R_{\mathfrak{s},\mathfrak{d}}\ ,$
$Q_{\mathfrak{s},\mathfrak{d}}$ $=R_{\mathfrak{s},\mathfrak{d}}-\mathrm{U}\left\langle\mathfrak{s},\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}P_{0}\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right\rangle$
$=R_{\mathfrak{s},\mathfrak{d}}-\mathrm{U}Q_{\mathfrak{s},\mathfrak{s}}R_{\mathfrak{s},\mathfrak{d}}\ ,$
$Q_{\mathfrak{d},\mathfrak{s}}$ $=R_{\mathfrak{d},\mathfrak{s}}-\mathrm{U}\left\langle\mathfrak{d}\left(k\right),\left(A_{1,1}\left(\mathrm{U},k\right)-\lambda\mathfrak{1}\right)^{-1}P_{0}\left(B_{1,1}\left(k\right)-\lambda\mathfrak{1}\right)^{-1}\mathfrak{s}\right\rangle$
$=R_{\mathfrak{d},\mathfrak{s}}-\mathrm{U}Q_{\mathfrak{d},\mathfrak{s}}R_{\mathfrak{s},\mathfrak{s}}\ .$
In matrix notation, the above equations can be rewritten as
$\begin{pmatrix}R_{\mathfrak{s},\mathfrak{s}}&R_{\mathfrak{d},\mathfrak{s}}\\[5.0pt] R_{\mathfrak{s},\mathfrak{d}}&R_{\mathfrak{d},\mathfrak{d}}\end{pmatrix}=\mathrm{U}\begin{pmatrix}Q_{\mathfrak{s},\mathfrak{s}}R_{\mathfrak{s},\mathfrak{s}}&Q_{\mathfrak{d},\mathfrak{s}}R_{\mathfrak{s},\mathfrak{s}}\\ Q_{\mathfrak{s},\mathfrak{s}}R_{\mathfrak{s},\mathfrak{d}}&Q_{\mathfrak{d},\mathfrak{s}}R_{\mathfrak{s},\mathfrak{d}}\end{pmatrix}+\begin{pmatrix}Q_{\mathfrak{s},\mathfrak{s}}&Q_{\mathfrak{d},\mathfrak{s}}\\ Q_{\mathfrak{s},\mathfrak{d}}&Q_{\mathfrak{d},\mathfrak{d}}\end{pmatrix}=\begin{pmatrix}\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1&0\\[5.0pt] \mathrm{U}R_{\mathfrak{s},\mathfrak{d}}&1\end{pmatrix}\begin{pmatrix}Q_{\mathfrak{s},\mathfrak{s}}&Q_{\mathfrak{d},\mathfrak{s}}\\ Q_{\mathfrak{s},\mathfrak{d}}&Q_{\mathfrak{d},\mathfrak{d}}\end{pmatrix}\ .$
As $(B_{1,1}(k)-\lambda\mathfrak{1})^{-1}\geq 0$ (because of (93)) and $\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}\geq 0$,
$\det\begin{pmatrix}\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1&0\\[5.0pt] \mathrm{U}R_{\mathfrak{s},\mathfrak{d}}&1\end{pmatrix}=\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1>0\ ,$
which means that the matrix appearing in the above determinant is invertible. From this we conclude that
$\begin{pmatrix}Q_{\mathfrak{s},\mathfrak{s}}&Q_{\mathfrak{d},\mathfrak{s}}\\ Q_{\mathfrak{s},\mathfrak{d}}&Q_{\mathfrak{d},\mathfrak{d}}\end{pmatrix}=\frac{1}{\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1}\begin{pmatrix}1&0\\ -\mathrm{U}R_{\mathfrak{s},\mathfrak{d}}&\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1\end{pmatrix}\begin{pmatrix}R_{\mathfrak{s},\mathfrak{s}}&R_{\mathfrak{d},\mathfrak{s}}\\[5.0pt] R_{\mathfrak{s},\mathfrak{d}}&R_{\mathfrak{d},\mathfrak{d}}\end{pmatrix}\ .$
In particular, since $R_{\mathfrak{s},\mathfrak{d}}=\overline{R_{\mathfrak{d},\mathfrak{s}}}$,
$\mathfrak{T}\left(\mathrm{U},k,\lambda\right)=Q_{\mathfrak{d},\mathfrak{d}}=R_{\mathfrak{d},\mathfrak{d}}-\frac{\mathrm{U}}{\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1}\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}=\frac{R_{\mathfrak{d},\mathfrak{d}}}{\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1}+\mathrm{U}\frac{R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}-\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}}{\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}+1}\ .$
Because $\lambda<\mathfrak{z}(k)\leq 0$ and $(B_{1,1}(k)-\lambda\mathfrak{1})^{-1}\geq\left|\mathfrak{z}(k)-\lambda\right|^{-1}\mathfrak{1}$ (see (93)), the sesquilinear form
$\left(\varphi,\psi\right)\mapsto\langle\varphi,(B_{1,1}(k)-\lambda\mathfrak{1})^{-1}\psi\rangle$
is a scalar product and using the Cauchy-Schwarz inequality [^13] Recall that the Cauchy-Schwarz inequality applied to a scalar product is an equality iff the vectors are linearly dependent., we deduce that
$R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}-\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}\geq 0\ .$
When $r_{\mathfrak{p}}>0$, the set $\{\mathfrak{d}(k),\mathfrak{s}\}$ is linearly independent for every $k\in\mathbb{T}^{2}$.
The last lemma is useful to deduce the behavior of the quantity $\mathfrak{T}(\mathrm{U},k,\lambda)$ at large Hubbard coupling constant $\mathrm{U}\gg 1$:
Corollary 4.12 ($\mathfrak{T}(\mathrm{U}_{0},k,\lambda)$ at large on-site repulsions)
Let $k\in\mathbb{T}^{2}$ and $\lambda<\mathfrak{z}(k)$, with $\mathfrak{z}(k)\in\mathbb{R}$ defined in Corollary 4.2. Then, for all $\mathrm{U}\in\mathbb{R}_{0}^{+}$,
$0\leq\mathfrak{T}\left(\mathrm{U},k,\lambda\right)-\mathfrak{T}\left(\infty,k,\lambda\right)\leq\frac{R_{\mathfrak{d},\mathfrak{d}}}{1+\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}}\ ,$
where
$\mathfrak{T}\left(\infty,k,\lambda\right)\doteq R_{\mathfrak{s},\mathfrak{s}}^{-1}\left({R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}-}\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}\right)\geq 0{\ .}$ (94)
Proof. By Lemma 4.11, we have $R_{\mathfrak{s},\mathfrak{s}}>0$ and $R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}\geq\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}$ while
$\mathfrak{T}(\mathrm{U},k,\lambda)-{\frac{R_{\mathfrak{d},\mathfrak{d}}R_{\mathfrak{s},\mathfrak{s}}-\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}}{R_{\mathfrak{s},\mathfrak{s}}}}=\frac{\left|R_{\mathfrak{s},\mathfrak{d}}\right|^{2}}{\left(1+\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}\right)R_{\mathfrak{s},\mathfrak{s}}}\leq\frac{R_{\mathfrak{d},\mathfrak{d}}}{1+\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}}\ .$
4.3.2 Hard-core dispersion relation of bound pairs of lowest energy
We are now in a position to study the spectral properties of the model $H\in\mathcal{B}(\mathfrak{H})$ defined by (19) in the hard-core limit. We study in particular its ground state energy $E(\mathrm{U})$ (see Corollary 3.4) and the limit of the continuous function $\mathrm{E}:\mathbb{T}^{2}\rightarrow\mathbb{R}$ defined by Theorem 4.7, which corresponds to the continuous family of non-degenerate eigenvalues at lowest energies in the fibers.
We start with the hard-core ground state energy, which is a well-defined quantity that even stays negative:
Lemma 4.13 (Existence of the hard-core ground state energy)
In the hard-core limit $\mathrm{U}\rightarrow\infty$, the hard-core ground state energy (42) is well-defined and is equal to
$E\left(\infty\right)={\sup\limits_{\mathrm{U}\in\mathbb{R}_{0}^{+}}}E(\mathrm{U})\leq 0.$
Proof. When $0\leq\mathrm{U}\leq\mathrm{V}$, one obviously has
$A(\mathrm{V},k)-A(\mathrm{U},k)=(\mathrm{V}-\mathrm{U})\begin{pmatrix}P_{0}&0\\ 0&0\end{pmatrix}\geq 0\ ,$
see Equation (40). In other words, $A(\mathrm{U},k)$, $\mathrm{U}\in\mathbb{R}_{0}^{+}$, defines an increasing family of bounded operators and by Proposition 4.1 and Equation (63),
$\min\sigma\left(A\left(\mathrm{U},k\right)\right)\leq\min\sigma\left(A\left(\mathrm{V},k\right)\right)\leq\mathfrak{z}\left(k\right)$ (95)
whenever $0\leq\mathrm{U}\leq\mathrm{V}$. In particular, by taking the minimum over $k\in\mathbb{T}^{2}$, if $0\leq\mathrm{U}\leq\mathrm{V}$ then
$E(\mathrm{U})\leq E(\mathrm{V})\leq 0\ .$
See Proposition 4.10. This shows that $E$ is an increasing function of $\mathrm{U}\in\mathbb{R}_{0}^{+}$, which is bounded from above by 00. This yields the assertion, thanks to the monotone convergence theorem.
We give in the next theorem a hard-core limit version of Theorems 4.4, 4.7 and 4.8. To this end, recall that $\mathrm{E}\left(k\right)$, $k\in\mathbb{T}^{2}$, are given by Theorem 4.7 as a family of non-degenerate eigenvalues. This family depends upon the parameter $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and we thus use here the notation $\mathrm{E}(\mathrm{U},k)\equiv\mathrm{E}(k)$. This defines a function $\mathrm{E}:\mathbb{R}_{0}^{+}\times\mathbb{T}^{2}\rightarrow\mathbb{R}$.
Theorem 4.14 (Dispersion relation in the hard-core limit)
Let $h_{b}\in[0,1/2]$. Recall that $\mathfrak{T}(\infty,k,\lambda)\geq 0$ is defined by (94).
  1. i.)
    For every $k\in\mathbb{T}^{2}$, the following limit exists:
    $\mathrm{E}\left(\infty,k\right)={\lim\limits_{\mathrm{U}\rightarrow\infty}}\mathrm{E}\left(\mathrm{U},k\right)={\sup\limits_{\mathrm{U}\in\mathbb{R}_{0}^{+}}}\mathrm{E}\left(\mathrm{U},k\right)\ ;$
  2. ii.)
    $\mathrm{E}\left(\infty,\cdot\right):\mathbb{T}^{2}\rightarrow\mathbb{R}$ is a continuous function;
  3. iii.)
    For $k\neq 0$, $\mathrm{E}(\infty,k)$ is the unique solution to the equation
    $\hat{\upsilon}\left(k\right)^{2}\mathfrak{T}\left(\infty,k,z\right)+z-\mathfrak{b}\left(k\right)=0\text{ },\qquad z<\mathfrak{z}(k)\ .$ (96)
  4. iv.)
    If $\hat{\upsilon}$ is of class $C^{d}$ on $\mathbb{S}^{2}\subseteq\mathbb{R}^{2}$ with $d\in\mathbb{N}\cup\{\omega,a\}$ then so does $\mathrm{E}(\infty,\cdot)$.
If, in addition, $r_{\mathfrak{p}}>0$ then:
  1. v.)
    For every $k\in\mathbb{T}^{2}$, $\mathrm{E}(\infty,k)\leq\mathfrak{b}(k)$ with equality iff $\hat{\upsilon}(k)=0$.
Proof. Fix $h_{b}\in[0,1/2]$. By (78) and (95) together with Corollary 4.2 and Theorem 4.7,
$\mathrm{E}\left(\mathrm{U},k\right)\leq\mathrm{E}\left(\mathrm{V},k\right)\leq\mathfrak{b}\left(k\right)\leq\mathfrak{z}\left(k\right)=\min\sigma\left(B_{1,1}\left(k\right)\right)\ ,\qquad k\in\mathbb{T}^{2}\ ,$ (97)
whenever $0\leq\mathrm{U}\leq\mathrm{V}$. This shows that, at any fixed $k\in\mathbb{T}^{2}$, the function $\mathrm{U}\mapsto\mathrm{E}\left(\mathrm{U},k\right)$ from $\mathbb{R}_{0}^{+}$ to $\mathbb{R}$ is increasing and bounded. Therefore, for any $k\in\mathbb{T}^{2}$,
$\mathrm{E}\left(\infty,k\right)={\lim\limits_{\mathrm{U}\rightarrow\infty}}\mathrm{E}\left(\mathrm{U},k\right)={\sup\limits_{\mathrm{U}\in\mathbb{R}_{0}^{+}}}\mathrm{E}\left(\mathrm{U},k\right)\leq\mathfrak{b}\left(k\right)\ ,$ (98)
thanks to the monotone convergence theorem. In particular, Assertion (i) holds true.
By Combining Equation (98) with (82), note that
$\mathrm{E}\left(\infty,k\right)={\sup\limits_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\inf_{\psi\in\mathcal{H},\|\psi\|=1}}\left\langle\psi,A\left(\mathrm{U},k\right)\psi\right\rangle\ ,\qquad k\in\mathbb{T}^{2}\ .$
Meanwhile, for any $\varepsilon>0$ and $k_{0}\in\mathbb{T}^{2}$, there is $\delta>0$ such that
$\left\|A\left(\mathrm{U},k\right)-A\left(\mathrm{U},k_{0}\right)\right\|_{\mathrm{op}}=\left\|A\left(0,k\right)-A\left(0,k_{0}\right)\right\|_{\mathrm{op}}\leq\delta\ .$
As a consequence, similar to what is done after (82), the set
$\{\left\langle\psi,A(\mathrm{U},\cdot)\psi\right\rangle:\mathrm{U}\in\mathbb{R}_{0}^{+},\ \psi\in\mathcal{H}\text{ with }\left\|\psi\right\|=1\}$
is a family of equicontinuous functions on $\mathbb{T}^{2}$. Therefore, $\mathrm{E}(\infty,\cdot)$ is continuous and Assertion (ii) holds true.
Fix $k\neq 0$. If $\hat{\upsilon}(k)\neq 0$ and $\mathrm{U}\in\mathbb{R}^{+}$, then we deduce from Corollary 4.2, Theorem 4.4 and Equations (97)–(98) that
$\left\{\mathrm{E}\left(\mathrm{U},k\right):\mathrm{U}\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}\right\}\subseteq\rho\left(B_{1,1}\left(k\right)\right)$
and
$\hat{\upsilon}\left(k\right)^{-2}\left(\mathfrak{b}\left(k\right)-\mathrm{E}\left(\mathrm{U},k\right)\right)-\mathfrak{T}\left(\infty,k,\mathrm{E}\left(\mathrm{U},k\right)\right)=\mathfrak{T}\left(k,\mathrm{E}\left(\mathrm{U},k\right)\right)-\mathfrak{T}\left(\infty,k,\mathrm{E}\left(\mathrm{U},k\right)\right)\ .$
Invoking next Corollary 4.12, we obtain the inequality
$\left|\mathfrak{b}\left(k\right)-\mathrm{E}\left(\mathrm{U},k\right)-\hat{\upsilon}\left(k\right)^{2}\mathfrak{T}\left(\infty,k,\mathrm{E}\left(\mathrm{U},k\right)\right)\right|<{\frac{\hat{\upsilon}\left(k\right)^{2}R_{\mathfrak{d},\mathfrak{d}}}{\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}}}\text{ }.$
We can take the limit $\mathrm{U}\rightarrow\infty$ in this last inequality by using (98) and the continuity of the mapping
$\mathfrak{T}\left(\infty,k,\cdot\right):\rho\left(B_{1,1}\left(k\right)\right)\rightarrow\mathbb{R}^{+}$
at the point $\mathrm{E}(\infty,k)$, to arrive at the equality
$\hat{\upsilon}(k)^{2}\mathfrak{T}\left(\infty,k,\mathrm{E}\left(\infty,k\right)\right)+\mathrm{E}\left(\infty,k\right)-\mathfrak{b}\left(k\right)=0\ .$ (99)
This proves that $\mathrm{E}(\infty,k)$ is a solution to (96). There is no other solution below $\mathfrak{z}(k)$ because of the following arguments: Given any $\mathrm{U}\in[0,\infty]$, let
$f_{\mathrm{U}}\left(x\right)\doteq\hat{\upsilon}(k)^{2}\mathfrak{T}(\mathrm{U},k,x)+x\ ,\quad x\in(-\infty,\mathfrak{z}(k))\ .$
By Corollary 4.12, $(f_{\mathrm{U}})_{\mathrm{U}\in\mathbb{R}_{0}^{+}}$ converges pointwise to $f_{\infty}$, as $\mathrm{U}\rightarrow\infty$. Since the pointwise limit of monotonically increasing function is again monotonically increasing, it follows that $f_{\infty}$ is monotonically increasing. Given any $x<y<\mathfrak{z}(k)$, take any $r>0$ with $r\geq f_{\infty}(y)-f_{\infty}(x)\geq 0$. Then, for some $\mathrm{U}_{0}\in\mathbb{R}_{0}^{+}$ sufficiently large, one has
$-r<f_{\mathrm{U}_{0}}(y)-f_{\infty}(y)<r\qquad\text{and}\qquad-r<f_{\mathrm{U}_{0}}(x)-f_{\infty}(x)<r,$
so that
$2r$ $>(f_{\mathrm{U}_{0}}(y)-f_{\infty}(y))-(f_{\mathrm{U}_{0}}(x)-f_{\infty}(x))$
$=(f_{\mathrm{U}_{0}}(y)-f_{\mathrm{U}_{0}}(x))-(f_{\infty}(y)-f_{\infty}(x))$
$\geq f_{\mathrm{U}_{0}}(y)-f_{\mathrm{U}_{0}}(x)-r.$
Then, the mean value theorem combined with (77) implies that
$3r\geq f_{\mathrm{U}_{0}}(y)-f_{\mathrm{U}_{0}}(x)=f_{\mathrm{U}_{0}}^{\prime}(c)(y-x)\geq y-x>0$ (100)
for some $c\in(x,y)$. This implies that the function $f_{\infty}$ is strictly increasing on $(-\infty,\mathfrak{z}(k))$ and, hence, there is a unique solution, $\mathrm{E}\left(\infty,k\right)$, to (96). Meanwhile, if $\hat{\upsilon}(k)=0$ then Theorem 4.7 implies that $\mathrm{E}(\mathrm{U},k)=\mathfrak{b}(k)$ for all $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and, obviously, $\mathrm{E}(\infty,k)=\mathfrak{b}(k)$ is the unique solution to (96).
Consider now the open set $\mathcal{O}$ defined by (80). Let $\Phi(\infty,\cdot):\mathcal{O}\rightarrow\mathbb{R}$ be defined by
$\Phi\left(\infty,k,x\right)\doteq\hat{\upsilon}\left(k\right)^{2}\mathfrak{T}\left(\infty,k,x\right)+x-\mathfrak{b}\left(k\right)\ ,\qquad\left(k,x\right)\in\mathcal{O}\ .$ (101)
By Corollary 4.12, note that $\Phi(\infty,\cdot):\mathcal{O}\rightarrow\mathbb{R}$ is nothing else than the pointwise limit of the function $\Phi(\mathrm{U},\cdot):\mathcal{O}\rightarrow\mathbb{R}$ defined by (79):
$\lim_{\mathrm{U}\rightarrow\infty}\Phi\left(\mathrm{U},k,x\right)=\Phi\left(\infty,k,x\right)\ .$ (102)
The function $\Phi(\infty,\cdot,\cdot)$ is a continuously differentiable function satisfying
$\partial_{x}\Phi\left(\infty,k,x\right)\geq 1/3>0\ ,\qquad\left(k,x\right)\in\mathcal{O}\ ,$
thanks to Inequality (100). Observe also that $\Phi(\infty,\cdot)\in C^{d}(\mathcal{O})$ if $\hat{\upsilon}$ is of class $C^{d}$ on $\mathbb{S}^{2}\subseteq\mathbb{R}^{2}$ with $d\in\mathbb{N}\cup\{\omega,a\}$. Therefore, by repeating essentially the same argument used in the proof of Theorem 4.8, one concludes that $\mathrm{E}(\infty,\cdot)\in C^{d}(\mathbb{S}^{2})$ whenever $\hat{\upsilon}$ is of class $C^{d}$ on $\mathbb{S}^{2}\subseteq\mathbb{R}^{2}$ with $d\in\mathbb{N}\cup\{\omega,a\}$.
Finally, assume that $r_{\mathfrak{p}}>0$, that is, $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$. For $k\neq 0$, we deduce from (99) and Lemma 4.11 that
$\mathrm{E}(\infty,k)=\mathfrak{b}(k)\Leftrightarrow\hat{\upsilon}(k)^{2}\mathfrak{T}\left(\infty,k,\mathrm{E}\left(\infty,k\right)\right)=0\Leftrightarrow\hat{\upsilon}(k)=0\ .$
To conclude the proof of Assertion (v), it remains to show that $\hat{\upsilon}(0)\neq 0$ implies $\mathrm{E}(\infty,0)<\mathfrak{b}(0)$. Assume on the contrary that
$\lim_{\mathrm{U}\rightarrow\infty}\mathrm{E}(\mathrm{U},0)=\mathfrak{b}\left(0\right)\ ,$ (103)
keeping in mind that $\mathfrak{b}\left(0\right)=\mathfrak{z}(0)=0$. Then, we infer from Theorems 4.4 and 4.7 that the following equality must be true:
$\lim_{\mathrm{U}\rightarrow\infty}\mathfrak{T}\left(\mathrm{U},0,\mathrm{E}(\mathrm{U},0)\right)=\lim_{\mathrm{U}\rightarrow\infty}\hat{\upsilon}(0)^{-2}\left(\mathfrak{b}\left(0\right)-\mathrm{E}(\mathrm{U},0)\right)=0\ .$
By Lemma 4.11, it follows that
$\lim_{\mathrm{U}\rightarrow\infty}\frac{R_{\mathfrak{d},\mathfrak{d}}\left(\mathrm{U}\right)}{\mathrm{U}R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)+1}=0\ ,$ (104)
where, by a slight abuse of notation,
$R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)$ $\doteq$ $R_{\mathfrak{s},\mathfrak{s}}\left(0,\mathrm{E}(\mathrm{U},0)\right)\ ,$
$R_{\mathfrak{d},\mathfrak{d}}\left(\mathrm{U}\right)$ $\doteq$ $R_{\mathfrak{d},\mathfrak{d}}\left(0,\mathrm{E}(\mathrm{U},0)\right)\ ,$
$R_{\mathfrak{s},\mathfrak{d}}\left(\mathrm{U}\right)$ $\doteq$ $R_{\mathfrak{s},\mathfrak{d}}\left(0,\mathrm{E}(\mathrm{U},0)\right)\ .$
Therefore, we deduce from Corollary 4.12 that
$\lim_{\mathrm{U}\rightarrow\infty}\frac{R_{\mathfrak{d},\mathfrak{d}}\left(\mathrm{U}\right)R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)-\left|R_{\mathfrak{s},\mathfrak{d}}\left(\mathrm{U}\right)\right|^{2}}{R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)}=\lim_{\mathrm{U}\rightarrow\infty}\mathfrak{T}(\infty,0,\mathrm{E}(\mathrm{U},0))=0\ .$ (105)
Now, observe that
$\frac{R_{\mathfrak{d},\mathfrak{d}}\left(\mathrm{U}\right)R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)-\left|R_{\mathfrak{s},\mathfrak{d}}\left(\mathrm{U}\right)\right|^{2}}{R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)}$ (106)
$=$ $\frac{1}{R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)}\left\langle\left(R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)\mathfrak{d}\left(0\right)-R_{\mathfrak{s},\mathfrak{d}}\left(\mathrm{U}\right)\mathfrak{s}\right),(B_{1,1}(0)-\mathrm{E}(\mathrm{U},0)\mathfrak{1})^{-1}\left(R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)\mathfrak{d}\left(0\right)-R_{\mathfrak{s},\mathfrak{d}}\left(\mathrm{U}\right)\mathfrak{s}\right)\right\rangle$
$=$ $\left\langle\mathfrak{d}\left(0\right)-\mathfrak{\alpha\left(\mathrm{U}\right)s},(B_{1,1}(0)-\mathrm{E}(\mathrm{U},0)\mathfrak{1})^{-1}\left(\mathfrak{d}\left(0\right)-\mathfrak{\alpha\left(\mathrm{U}\right)s}\right)\right\rangle$
where
$\alpha\left(\mathrm{U}\right)\doteq R_{\mathfrak{s},\mathfrak{s}}\left(\mathrm{U}\right)^{-1}R_{\mathfrak{s},\mathfrak{d}}\left(\mathrm{U}\right)\ .$
Let $\varphi_{\mathrm{U}}=\mathfrak{d}(0)-\alpha(\mathrm{U})\mathfrak{s}\neq 0$ and $W_{\mathrm{U}}=B_{1,1}(0)-\mathrm{E}(\mathrm{U},0)\mathfrak{1}$. Since $W_{\mathrm{U}}$ is strictly positive, it has an unique square root, which is also strictly positive. Then
$||\varphi_{\mathrm{U}}||^{4}=|\langle W_{\mathrm{U}}^{-1/2}\varphi_{\mathrm{U}},W_{\mathrm{U}}^{1/2}\varphi_{\mathrm{U}}\rangle|^{2}\leq||W_{\mathrm{U}}^{-1/2}\varphi_{\mathrm{U}}||^{2}||W_{\mathrm{U}}^{1/2}\varphi_{\mathrm{U}}||^{2}\leq\\[10.00002pt] \leq||W_{\mathrm{U}}||_{\mathrm{op}}||\varphi_{\mathrm{U}}||^{2}\langle\varphi_{\mathrm{U}},W_{\mathrm{U}}^{-1}\varphi_{\mathrm{U}}\rangle\leq||\varphi_{\mathrm{U}}||^{2}\langle\varphi_{\mathrm{U}},W_{\mathrm{U}}^{-1}\varphi_{\mathrm{U}}\rangle\big{(}||B_{1,1}(0)||_{\mathrm{op}}+|\mathrm{E}(0,0)|\big{)}.$
As $\langle\varphi_{\mathrm{U}},W_{\mathrm{U}}^{-1}\varphi_{\mathrm{U}}\rangle$ tends to 00 when $\mathrm{U}\rightarrow\infty$, it follows that also $\varphi_{\mathrm{U}}$ vanishes in this limit, that is,
$\lim_{\mathrm{U}\rightarrow\infty}\mathfrak{\alpha\left(\mathrm{U}\right)s}=\mathfrak{d}\left(0\right)\ .$ (107)
As $\mathbb{C}\mathfrak{s}$ is a closed subspace of $L^{2}(\mathbb{T}^{2})$, we get that $\mathfrak{d}(0)\in\mathbb{C}\mathfrak{s}$, which is a contradiction. Therefore, Assertion (v) holds true.
Remark 4.15
Under the additional conditions that $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$, Assertion (iv) remains valid with $(-\pi,\pi)^{2}$ instead of $\mathbb{S}^{2}$, i.e., by including the zero quasi-momentum case $k=0$. Also, Assertion (iii) holds true for $k=0$. In fact, under these further conditions, we know from Assertion (v) that $\mathrm{E}(\infty,0)<\mathfrak{z}(0)=0$.
Corollary 4.16 (Hard-core dispersion relation)
Let $h_{b}\in[0,1/2]$ and $\hat{\upsilon}$ be of class $C^{2}$ on $\mathbb{S}^{2}$. Then, $\mathrm{E}(\infty,\cdot)\in C(\mathbb{T}^{2})$ and is of class $C^{2}$ on $\mathbb{S}^{2}$. In this case, for any $k\in\mathbb{S}^{2}$,
$\mathbf{v}_{\mathrm{E},\infty}\left(k\right)\doteq\vec{\nabla}_{k}\mathrm{E}\left(\infty,k\right)=-\left(\partial_{x}\Phi\left(\infty,{k,\mathrm{E}\left(\infty,k\right)}\right)\right)^{-1}\vec{\nabla}_{k}\Phi\left(\infty,{k,\mathrm{E}\left(\infty,k\right)}\right)={\lim_{\mathrm{U}\rightarrow\infty}}\mathbf{v}_{\mathrm{E},\mathrm{U}}\left(k\right)\ .$
Moreover, if $\hat{\upsilon}$ is real analytic (of class $C^{a}$, in our terminology) on $\mathbb{S}^{2}$, then either $\mathfrak{M}_{\mathrm{E}\left(\infty,\cdot\right)}$ has full measure or is empty.
Proof. Recall that $\mathcal{O}$ is the open set (80) and $\Phi(\cdot,\cdot):\mathbb{R}_{0}^{+}\times\mathcal{O}\rightarrow\mathbb{R}$ is the real-valued function defined by (79). Note that, for any $\mathrm{U}\in\mathbb{R}_{0}^{+}$, the function $\Phi(\mathrm{U},\cdot)$ is a smooth function on $\mathcal{O}$ and we estimate its derivatives with respect to the parameter $k$, at fixed $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and $x$, where $(k,x)\in\mathcal{O}$: For any $(k,x)\in\mathcal{O}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$, Equations (31)–(32) together with the second resolvent formula yield the derivative
$\partial_{k_{j}}\left(A_{1,1}\left(\mathrm{U},k\right)-x\mathfrak{1}\right)^{-1}=-\left(A_{1,1}\left(\mathrm{U},k\right)-x\mathfrak{1}\right)^{-1}\left\{\partial_{k_{j}}M_{\mathfrak{f}\left(k\right)}\right\}\left(A_{1,1}\left(\mathrm{U},k\right)-x\mathfrak{1}\right)^{-1}$
for any $j\in\{1,2\}$, where $k=(k_{1},k_{2})\in\mathbb{S}^{2}$. Therefore, by (28) and (79), for any $(k,x)\in\mathcal{O}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$,
$\left|\partial_{k_{j}}\Phi\left(\mathrm{U},k,x\right)\right|\leq 8\epsilon\hat{\upsilon}\left(k\right)^{2}\left|\mathfrak{z}\left(k\right)-x\right|^{-2}$ (108)
for any $j\in\{1,2\}$, where $k=(k_{1},k_{2})\in\mathbb{S}^{2}$. Taking the second derivative, one can easily check that
$\left|\partial_{k_{j}}^{2}\Phi\left(\mathrm{U},k,x\right)\right|\leq 8\epsilon\hat{\upsilon}\left(k\right)^{2}\left|\mathfrak{z}\left(k\right)-x\right|^{-3}$ (109)
for $j\in\{1,2\}$, where $k=(k_{1},k_{2})\in\mathbb{S}^{2}$. In the same way, we deduce from the identities
$\partial_{x}\Phi\left(\mathrm{U},k,x\right)$ $=$ $\hat{\upsilon}\left(k\right)^{2}\left\langle\mathfrak{d}\left(k\right),\left(A_{1,1}\left(\mathrm{U},k\right)-x\mathfrak{1}\right)^{-2}\mathfrak{d}\left(k\right)\right\rangle+1$
$\partial_{x}^{2}\Phi\left(\mathrm{U},k,x\right)$ $=$ $2\hat{\upsilon}\left(k\right)^{2}\left\langle\mathfrak{d}\left(k\right),\left(A_{1,1}\left(\mathrm{U},k\right)-x\mathfrak{1}\right)^{-3}\mathfrak{d}\left(k\right)\right\rangle$
the following inequalities:
$\left|\partial_{x}\Phi\left(\mathrm{U},k,x\right)\right|$$\leq$ $\hat{\upsilon}\left(k\right)^{2}\left|\mathfrak{z}\left(k\right)-x\right|^{-2}+1\ ,$ (110)
$\left|\partial_{x}^{2}\Phi\left(\mathrm{U},k,x\right)\right|$$\leq$ $2\hat{\upsilon}\left(k\right)^{2}\left|\mathfrak{z}\left(k\right)-x\right|^{-3}\ ,$ (111)
for any $(k,x)\in\mathcal{O}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$.
Now, fix $(k_{0},x_{0})\in\mathcal{O}$. The function $\mathfrak{z}$ of Corollary 4.2 is continuous with respect to $k_{1},k_{2}$, where $k=(k_{1},k_{2})\in\mathbb{S}^{2}$. Then, there is a closed cube centered at $(k_{0},x_{0})$ with side length $\delta\in\mathbb{R}^{+}$ contained in $\mathcal{O}$. Suppose that $\partial_{x}\Phi(\mathrm{U},k_{0},x_{0})$ does not converge to $\partial_{x}\Phi(\infty,k_{0},x_{0})$. Then we can find $r_{0}\in\mathbb{R}^{+}$ and a sequence $(\mathrm{U}_{n})_{n\in\mathbb{N}}$ of positive numbers such that $\mathrm{U}_{n}\rightarrow\infty$ as $n\rightarrow\infty$ and, for every $n\in\mathbb{N}$,
$|\partial_{x}\Phi(\mathrm{U}_{n},k_{0},x_{0})-\partial_{x}\Phi(\infty,k_{0},x_{0})|\geq r_{0}\ .$
Let
$\mathfrak{F}=\{\partial_{x}\Phi(\mathrm{U}_{n},k_{0},\cdot)\upharpoonright[x_{0}-\delta,x_{0}+\delta]\,:\,n\in\mathbb{N}\}\ .$
By combining Equation (111) with the mean value theorem, we see that $\mathfrak{F}$ is Lipschitz equicontinuous. In particular this family of functions is equicontinuous. Moreover, it follows from (110) that $\mathfrak{F}$ is bounded in the supremum norm. We can hence apply the (Arzelà-) Ascoli theorem [38, Theorem A5], according to which $\partial_{x}\Phi(\mathrm{U}_{n},k_{0},\cdot)$, when restricted to the compact interval $[x_{0}-\delta,x_{0}+\delta]$, converges uniformly along some subsequence. Assume, for simplicity and without loss of generality, that the (full) sequence of functions converges itself. By Equation (102), recall that $\Phi(\mathrm{U},\cdot,\cdot)$ converges pointwise to the function $\Phi(\infty,\cdot,\cdot)$ on $\mathcal{O}$, which is is the real-valued function defined by (101). Then, by [47, Theorem 7.17],
$\partial_{x}\Phi\left(\infty,k_{0},x\right)=\lim_{n\rightarrow\infty}\partial_{x}\Phi\left(\mathrm{U}_{n},k_{0},x\right)\ ,\qquad x\in[x_{0}-\delta,x_{0}+\delta]\ .$
For $x=x_{0}$, this lead us to a contradiction. Thus
$\partial_{x}\Phi\left(\infty,k,x\right)=\lim_{\mathrm{U}\rightarrow\infty}\partial_{x}\Phi\left(\mathrm{U},k,x\right)$
for every $(k,x)\in\mathcal{O}$. In the same way, we invoke Equations (108) and (109) together with the mean value theorem and the (Arzelà-) Ascoli theorem [38, Theorem A5] to deduce that
$\partial_{k_{j}}\Phi\left(\infty,k,x\right)=\lim_{\mathrm{U}\rightarrow\infty}\partial_{k_{j}}\Phi\left(\mathrm{U},k,x\right)\ ,\qquad j=1,2\ .$
for every $(k,x)\in\mathcal{O}$. To prove the corollary, we eventually use these observations together with Theorem 4.14 (i), Corollary 4.9 and the equicontinuity of
$\{\partial_{\mu}\Phi\left(\mathrm{U},k_{0},\cdot\right)\upharpoonright(-\infty,\mathfrak{b}(k_{0})]\,:\,\mathrm{U}\in\mathbb{R}^{+}\}\ ,$
with $\mu$ standing for the variables $k_{1}$, $k_{2}$ or $x$. Note that the last assertion concerning $\mathfrak{M}_{\mathrm{E}\left(\infty,\cdot\right)}$ is a direct consequence of the fact that the zeros of any non-constant real analytic function have null Lebesgue measure [43], see the proof of the same assertion for $\mathrm{U}<\infty$ in Corollary 4.9.
Note that no additional condition is required for $\mathrm{E}$ to have a well-defined hard-core limit. Compare Corollary 4.9 with Theorem 4.14. Moreover, by Corollary 4.16, $\mathrm{E}(\infty,\cdot)$ can be viewed as the (effective) dispersion relation of the dressed bound fermion pairs, with lowest energy, in the hard-core limit.
We close this section by showing the convergence of the low-energy eigenvector of $A(\mathrm{U},k)$ for large Hubbard couplings. This refers to the (hard-core) limit $\mathrm{U}\rightarrow\infty$ of the vector
$\Psi\left(\mathrm{U},k\right)\equiv\Psi\left(k\right)\doteq\left(\hat{\upsilon}\left(k\right)\left(A_{1,1}\left(\mathrm{U},k\right)-\mathrm{E}\left(\mathrm{U},k\right)\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right),-1\right)\in\mathcal{H}\ ,$
see Equation (118).
Proposition 4.17 (Hard-core limit of eigenvectors)
Let $h_{b}\in[0,1/2]$. Fix $k\in\mathbb{T}^{2}\backslash\{0\}$. The following limit exits:
$\0\Psi\left(\infty,k\right)\doteq{\lim_{\mathrm{U}\rightarrow\infty}}\,\Psi\left(\mathrm{U},k\right)\in\mathcal{H}\backslash\{0\}\ .$
This statement remains valid for $k=0$ provided that $\hat{\upsilon}(0)\neq 0$ and $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$).
Proof. Fix $k\in\mathbb{T}^{2}$ with $k\neq 0$. By using the first resolvent formula together with Theorem 4.7, Theorem 4.14 (i) and (v), we find that
$\|(A_{1,1}(\mathrm{U},k)-\mathrm{E}\left(\infty,k\right)\mathfrak{1})^{-1}-(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\|_{\mathrm{o}\mathrm{p}}\leq{\frac{|\mathrm{E}(\mathrm{U},k)-\mathrm{E}(\infty,k)|}{|\mathfrak{z}(k)-\mathrm{E}(\infty,k)|^{2}}}\rightarrow 0\ .$
Note that $\mathrm{E}(\infty,k)\leq\mathfrak{b}\left(k\right)<\mathfrak{z}\left(k\right)$ for any $k\neq 0$. (When $\hat{\upsilon}(0)\neq 0$ and $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$, we also have that $\mathrm{E}(\infty,0)<\mathfrak{b}(0)=\mathfrak{z}(0)$.) On the other hand, by Proposition 5.14, $\{(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\infty,k)\mathfrak{1})^{-1}\}_{\mathrm{U}\geq 0}$ is a decreasing family of positive operators and, by Proposition 5.15, it converges strongly as $\mathrm{U}\rightarrow\infty$. Consequently, $(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}$ also converges strongly.
4.4 Spectral Gap and Anderson Localization
By Equation (78), the spectral gap of fiber Hamiltonians is equal to
$\mathfrak{g}\left(\mathrm{U},k\right)\doteq\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},k\right)\right)-\mathrm{E}\left(\mathrm{U},k\right)\geq 0,\qquad k\in\mathbb{T}^{2}\ ,$ (112)
for any Hubbard coupling constant $\mathrm{U}\in\mathbb{R}_{0}^{+}$. When $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$, this quantity turns out to be strictly positive, uniformly with respect to the parameter $\mathrm{U}$:
Proposition 4.18 (Uniform spectral gap of fiber Hamiltonians)
Fix $h_{b}\in[0,1/2]$. If $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$ then
$\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)>0\ .$
Proof. The family $\{\mathrm{E}(\mathrm{U},\cdot)\}_{\mathrm{U}\in\mathbb{R}_{0}^{+}}$ of real-valued functions on $\mathbb{T}^{2}$ is equicontinuous, thanks to Theorem 4.8 (i). Since Proposition 4.1 says that, for any $k\in\mathbb{T}^{2}$,
$\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},k\right)\right)=\mathfrak{z}\left(k\right)\doteq 4\epsilon-2\epsilon\cos\left(k/2\right)\ ,$ (113)
we thus deduce from (112) that the family $\{\mathfrak{g}(\mathrm{U},\cdot)\}_{\mathrm{U}\in\mathbb{R}_{0}^{+}}$ of real-valued functions on $\mathbb{T}^{2}$ is equicontinuous. It follows that the function
$\mathbb{R}_{0}^{+}\ni\mathrm{U}\longmapsto\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)\in\mathbb{R}$ (114)
is continuous. Moreover, from the compactness of $\mathbb{T}^{2}$, $\mathfrak{g}(\mathrm{U},\cdot)$ has a global minimizer, say $k_{\mathrm{U}}\in\mathbb{T}^{2}$ for all $\mathrm{U}\in\mathbb{R}_{0}^{+}$. Since $\hat{\upsilon}(0)\neq 0$, by Theorem 4.7,
$\mathrm{E}\left(\mathrm{U},k_{\mathrm{U}}\right)<\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},k_{\mathrm{U}}\right)\right)=0$
when $k_{\mathrm{U}}=0$, while in the case $k_{\mathrm{U}}\neq 0$,
$\mathrm{E}\left(\mathrm{U},k_{\mathrm{U}}\right)\leq\mathfrak{b}(k_{\mathrm{U}})<\mathfrak{z}\left(k\right)=\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},k_{\mathrm{U}}\right)\right)\ .$
In particular,
$\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)=\mathfrak{g}\left(\mathrm{U},k_{\mathrm{U}}\right)>0\ ,\qquad\mathrm{U}\in\mathbb{R}_{0}^{+}\ .$
Using this together with the continuity of the function (114) we arrive at the inequality
$\inf_{\mathrm{U}\in\left[0,c\right]}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)>0$ (115)
for any positive parameter $c\in\mathbb{R}_{0}^{+}$. Now, we perform the limit $\mathrm{U}\rightarrow\infty$. Since $\mathbb{T}^{2}$ is compact, the net$\ (k_{\mathrm{U}})_{\mathrm{U}\in\mathbb{R}_{0}^{+}}$ converges along subnets (in fact subsequences). Assume without loss of generality that $(k_{\mathrm{U}})_{\mathrm{U}\in\mathbb{R}_{0}^{+}}$ converges to some $k_{\infty}\in\mathbb{T}^{2}$ (otherwise one uses all the following arguments on subsequences). If $k_{\infty}\neq 0$ then
$\lim_{\mathrm{U}\rightarrow\infty}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)\geq\lim_{\mathrm{U}\rightarrow\infty}\mathfrak{z}\left(k_{\mathrm{U}}\right)-\mathfrak{b}(k_{\mathrm{U}})=\mathfrak{z}\left(k_{\infty}\right)-\mathfrak{b}(k_{\infty})>0\ ,$ (116)
thanks to Theorem 4.7 and the continuity of the functions $\mathfrak{z}$ and $\mathfrak{b}$. Assume now that $k_{\infty}=0$. Since, for all $\mathrm{U}\in\mathbb{R}_{0}^{+}$,
$\left|\mathrm{E}\left(\mathrm{U},k_{\mathrm{U}}\right)-\mathrm{E}\left(\infty,0\right)\right|\leq\left|\mathrm{E}\left(\mathrm{U},k_{\mathrm{U}}\right)-\mathrm{E}\left(\mathrm{U},0\right)\right|+\left|\mathrm{E}\left(\mathrm{U},0\right)-\mathrm{E}\left(\infty,0\right)\right|,$
we infer from the equicontinuity of the family $\{\mathrm{E}(\mathrm{U},\cdot)\}_{\mathrm{U}\in\mathbb{R}_{0}^{+}}$ (Theorem 4.8 (i)) and Theorem 4.14 (i) that
$\lim_{\mathrm{U}\rightarrow\infty}\mathrm{E}\left(\mathrm{U},k_{\mathrm{U}}\right)=\mathrm{E}\left(\infty,0\right)\ .$
Combined with Theorem 4.14 (v) and $\hat{\upsilon}(0)\neq 0$, this last limit in turn implies that
$\lim_{\mathrm{U}\rightarrow\infty}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)=\lim_{\mathrm{U}\rightarrow\infty}\left\{\mathfrak{z}(k_{\mathrm{U}})-\mathrm{E}\left(\mathrm{U},k_{\mathrm{U}}\right)\right\}=\mathfrak{z}\left(0\right)-\mathrm{E}\left(\infty,0\right)=-\mathrm{E}\left(\infty,0\right)>0\ .$ (117)
The assertion is therefore a combination of Inequalities (115), (116) and (117).
We study now the space localization of the (dressed) bound pair with total quasi-momentum $k\in\mathbb{T}^{2}$ and energy $\mathrm{E}\left(\mathrm{U},k\right)$. Assume that $\hat{\upsilon}(0)\neq 0$. By Corollary 4.5, for any fixed $k\in\mathbb{T}^{2}$, it corresponds to study the fermionic part of the eigenvector
$\Psi\left(\mathrm{U},k\right)\doteq g\left(k,\mathrm{E}\left(\mathrm{U},k\right)\right)=\left(\hat{\upsilon}\left(k\right)\left(A_{1,1}\left(\mathrm{U},k\right)-\mathrm{E}\left(\mathrm{U},k\right)\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right),-1\right)\in\mathcal{H}\ ,$ (118)
written in the real space $\mathbb{Z}^{2}$ via the inverse Fourier transform $\mathcal{F}^{-1}$ (see (26)). This function is denoted by
$\psi_{\mathrm{U},k}\doteq\mathcal{F}^{-1}\left[\hat{\upsilon}\left(k\right)\left(A_{1,1}\left(\mathrm{U},k\right)-\mathrm{E}\left(\mathrm{U},k\right)\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right]\in\ell^{2}\left(\mathbb{Z}^{2}\right)$ (119)
for any fixed $k\in\mathbb{T}^{2}$. One should not be confused here by the parameter $k$. Recall for instance that, given $k\in\mathbb{T}^{2}$, $\mathfrak{d}(k)\in C\left(\mathbb{T}^{2}\right)$ is itself a function on the torus $\mathbb{T}^{2}$, defined by
$\mathfrak{d}(k)(p)\doteq\mathfrak{\hat{p}}(k+p)\ ,\qquad p\in\mathbb{T}^{2}\ ,$ (120)
see Equation (29). In particular, observe that
$\psi_{\mathrm{U},k}=\hat{\upsilon}\left(k\right)\mathcal{F}^{-1}\left[\left(A_{1,1}\left(\mathrm{U},k\right)-\mathrm{E}\left(\mathrm{U},k\right)\mathfrak{1}\right)^{-1}\mathfrak{d}\left(k\right)\right]\ .$ (121)
We now show that this function is exponentially localized in the real space:
Theorem 4.19 (Exponentially localized dressed bound fermion pairs)
Fix $h_{b}\in[0,1/2]$, $k\in\mathbb{T}^{2}$ and suppose that $r_{\mathfrak{p}}>0$ (i.e., $\mathfrak{p}\notin\mathbb{C}\mathfrak{e}_{0}$) and $\hat{\upsilon}(0)\neq 0$. There exist positive constants $C,\alpha>0$ such that, for all $k\in\mathbb{T}^{2}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$,
$\left|\psi_{\mathrm{U},k}(x)\right|\leq C\mathrm{e}^{-\alpha|x|}\ ,\qquad x\in\mathbb{Z}^{2}\ .$
Proof. By (26) and (120), for any $p,k\in\mathbb{T}^{2}$, note that
$\mathfrak{d}(k)(p)={\sum_{x\in\mathbb{Z}^{2}}}\,\mathrm{e}^{i(k+p)\cdot x}\mathfrak{p}(x)=\mathcal{F}[\mathrm{e}^{ik\cdot(x)}\mathfrak{p}(x)](p)\ ,$
where $\mathrm{e}^{ik\cdot(x)}\mathfrak{p}(x)$ stands for the function $x\mapsto\mathrm{e}^{ik\cdot(x)}\mathfrak{p}(x)$. It follows that
$\mathcal{F}^{-1}\left[(A_{1,1}(\mathrm{U},k)-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\mathfrak{d}(k)\right]$ $=(\mathcal{F}^{-1}A_{1,1}(\mathrm{U},k)\mathcal{F}-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\mathcal{F}^{-1}\mathfrak{d}(k)$
$={\sum_{y\in\mathbb{Z}^{2}}}\,\mathrm{e}^{ik\cdot y}\mathfrak{p}(y)(\mathcal{F}^{-1}A_{1,1}(\mathrm{U},k)\mathcal{F}-\mathrm{E}(\mathrm{U},k)\mathfrak{1})^{-1}\mathfrak{e}_{y}\ .$ (122)
By Equation (121), it suffices to estimate the exponential decay of this particular function. This is done by using the celebrated Combes-Thomas estimates, which correspond here to Theorem 5.13. To this end, several quantities, one of them being related to the spectral gap $\mathfrak{g}(\mathrm{U},k)$ (112), have to be controlled and, as in Section 5.5, we use the notation (146), i.e.,
$\Delta(\lambda;T)\doteq\min\left\{\left|\lambda-a\right|:a\in\sigma(T)\right\}$ (123)
for the distance between a complex number $\lambda\in\mathbb{C}$ and the spectrum $\sigma(T)$ of an operator $T\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}))$, as well as (144), which, in the present case, refers to the quantity
$\mathbf{S}(T,\mu)\doteq{\sup_{x\in\mathbb{Z}^{2}}\,\sum_{y\in\mathbb{Z}^{2}}}\,\left(\mathrm{e}^{\mu\left|x-y\right|}-1\right)\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|\in\left[0,\infty\right]$ (124)
for any $T\in\mathcal{B}(\ell^{2}(\mathbb{Z}^{2}))$ and $\mu\in\mathbb{R}_{0}^{+}$. We do it in three steps: The first one controls the spectral gap $\mathfrak{g}(\mathrm{U},k)$ (112) and a quantity like (123) for $\lambda=\mathrm{E}(\mathrm{U},k)$, while the second step is an analysis of quantities like (124). These two steps allow us to apply, in the last step, Theorem 5.13 in order to get the desired result.
Step 1: Observe from Equation (112) and Proposition 4.18 that we can find $\alpha>0$ such that, for all $k\in\mathbb{T}^{2}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$,
$0<4\epsilon(\mathrm{e}^{\alpha}-1)<\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)\leq\mathfrak{g}\left(\mathrm{U},k\right)\doteq\min\sigma_{\mathrm{ess}}\left(A\left(\mathrm{U},k\right)\right)-\mathrm{E}\left(\mathrm{U},k\right)\ .$
Using now Proposition 4.1 and the fact that
$\min\sigma_{\mathrm{ess}}(A_{1,1}(\mathrm{U},k))=\min\sigma(A_{1,1}(\mathrm{U},k))\ ,$
for all $k\in\mathbb{T}^{2}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$, we deduce from the last inequalities that
$0<4\epsilon(\mathrm{e}^{\alpha}-1)<\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)\leq\Delta\left(\mathrm{E}(\mathrm{U},k);A_{1,1}(\mathrm{U},k)\right)=\min\sigma(A_{1,1}(\mathrm{U},k))-\mathrm{E}(\mathrm{U},k)\ ,$
see also Equation (123). Since $\mathcal{F}$ is a unitary transformation, $A_{1,1}(\mathrm{U},k)$ and $\mathcal{F}^{\ast}A_{1,1}(\mathrm{U},k)\mathcal{F}$ have the same spectrum and it follows that
$0<4\epsilon(\mathrm{e}^{\alpha}-1)<\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)\leq\Delta\left(\mathrm{E}(\mathrm{U},k);\mathcal{F}^{\ast}A_{1,1}(\mathrm{U},k)\mathcal{F}\right)$ (125)
for all $k\in\mathbb{T}^{2}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$.
Step 2: By Equations (5) and (26), one easily checks that
$\mathfrak{\hat{e}}_{y}\left(p\right)\doteq{\sum_{x\in\mathbb{Z}^{2}}}\,\mathrm{e}^{ip\cdot x}\mathfrak{e}_{y}\left(x\right)=\mathrm{e}^{ip\cdot y}\ ,\qquad p\in\mathbb{T}^{2},\ y\in\mathbb{Z}^{2}\ ,$
while, for any fixed $k\in\mathbb{T}^{2}$, the real-valued functions $\mathfrak{f}(k)$, defined by (28) and (30) on the torus $\mathbb{T}^{2}$, can be rewritten as
$\mathfrak{f}\left(k\right)\left(p\right)\doteq\epsilon\left\{4-\cos\left(p+k\right)-\cos\left(p\right)\right\}=4\epsilon-{\frac{\epsilon}{2}\sum_{z\in\mathbb{Z}^{2}\,,\,|z|=1}}\left(\mathrm{e}^{i\left(p+k\right)\cdot z}+\mathrm{e}^{ip\cdot z}\right)\ ,\qquad p\in\mathbb{T}^{2}\ .$
Therefore, since $M_{\mathfrak{f}\left(k\right)}$ stands for the multiplication operator by $\mathfrak{f}(k)\in C(\mathbb{T}^{2})$, for every $p,k\in\mathbb{T}^{2}$ and $y\in\mathbb{Z}^{2}$,
$M_{\mathfrak{f}\left(k\right)}\mathfrak{\hat{e}}_{y}(p)=4\mathrm{e}^{ip\cdot y}\epsilon-{\frac{\epsilon}{2}\sum_{z\in\mathbb{Z}^{2}\,,\,|z|=1}}\left(\mathrm{e}^{ik\cdot z}+1\right)\mathrm{e}^{ip\cdot(y+z)}=4\epsilon\mathfrak{\hat{e}}_{y}(p)-{\frac{\epsilon}{2}\sum_{z\in\mathbb{Z}^{2}\,,\,|z|=1}}\left(\mathrm{e}^{ik\cdot z}+1\right)\mathfrak{\hat{e}}_{y+z}(p)\ ,$
which, by (31)–(32), in turn implies that
$\mathcal{F}^{\ast}A_{1,1}(\mathrm{U},k)\mathcal{F}\mathfrak{e}_{y}$ $=\mathcal{F}^{\ast}\left(M_{\mathfrak{f}\left(k\right)}\mathfrak{\hat{e}}_{y}+\mathrm{U}P_{0}\mathfrak{\hat{e}}_{y}+{\sum_{z\in\mathbb{Z}^{2}}}\,\mathrm{u}(z)P_{z}\mathfrak{\hat{e}}_{y}\right)$
$=\mathcal{F}^{\ast}\left(M_{\mathfrak{f}\left(k\right)}\mathfrak{\hat{e}}_{y}+\left(\mathrm{U}\delta_{y,0}+\mathrm{u}(y)\right)\mathfrak{\hat{e}}_{y}\right)$
$=4\epsilon\mathfrak{e}_{y}-{\frac{\epsilon}{2}\sum_{z\in\mathbb{Z}^{2}\,,\,|z|=1}}\left(\mathrm{e}^{ik\cdot z}+1\right)\mathfrak{e}_{y+z}+\left(\mathrm{U}\delta_{y,0}+\mathrm{u}(y)\right)\mathfrak{e}_{y}\ ,$
keeping in mind that $P_{x}$ is the orthogonal projection onto the one-dimensional subspace $\mathbb{C}\hat{\mathfrak{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. Recall that $\mathcal{F}^{\ast}=\mathcal{F}^{-1}$, the Fourier transform being unitary. Thus, since $\alpha>0$, for each $x\in\mathbb{Z}^{2}$, we obtain that
${\sum_{y\in\mathbb{Z}^{2}}}\,\left|\mathrm{e}^{\alpha\left|x-y\right|}-1\right|\left|\left\langle\mathfrak{e}_{x},\mathcal{F}^{\ast}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F}\mathfrak{e}_{y}\right\rangle\right|=\frac{\epsilon}{2}\left(\mathrm{e}^{\alpha}-1\right){\sum_{y\in\mathbb{Z}^{2},\,|x-y|=1}}\left|\mathrm{e}^{ik\cdot(x-y)}+1\right|\leq 4\epsilon\left(\mathrm{e}^{\alpha}-1\right)\ .$
Hence, taking the supremum over all $x\in\mathbb{Z}^{2}$ in this equation and using the notation given by (124) as well as (125), we arrive at
$\mathbf{S}\left(\mathcal{F}^{\ast}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F},\alpha\right)\leq 4\epsilon(\mathrm{e}^{\alpha}-1)<\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)\leq\Delta\left(\mathrm{E}\left(\mathrm{U},k\right);\mathcal{F}^{\ast}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F}\right)$ (126)
for any fixed $k\in\mathbb{T}^{2}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$.
Step 3: Thanks to (126), we are now in a position to apply Theorem 5.13 for $H=\mathcal{F}^{\ast}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F}$ and $\mu=\alpha$ to obtain that, for any $x,y\in\mathbb{Z}^{2}$,
$\left|\left\langle\mathfrak{e}_{x},\left(\mathcal{F}^{\ast}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F-}\mathrm{E}(\mathrm{U},k)\mathfrak{1}\right)^{-1}\mathfrak{e}_{y}\right\rangle\right|$
$\leq$ $\frac{\mathrm{e}^{-\alpha\left|x-y\right|}}{\Delta\left(\mathrm{E}\left(\mathrm{U},k\right);\mathcal{F}^{\ast}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F}\right)-\mathbf{S}\left(\mathcal{F}^{\ast}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F},\alpha\right)}$
$\leq$ $\frac{\mathrm{e}^{-\alpha\left|x-y\right|}}{\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)-4\epsilon(\mathrm{e}^{\alpha}-1)}\ .$
Combined with Equations (121)–(122) and the triangle inequality as well as the reverse one $|x-y|\geq|x|-|y|$, we then arrive at
$\left|\psi_{\mathrm{U},k}\left(x\right)\right|$ $=$ $\left|\left\langle\mathfrak{e}_{x},\psi_{\mathrm{U},k}\right\rangle\right|\leq\left|\hat{\upsilon}\left(k\right)\right|{\sum_{y\in\mathbb{Z}^{2}}}\left|\mathfrak{p}\left(y\right)\right|\left|\left\langle\mathfrak{e}_{x},\left(\mathcal{F}^{-1}A_{1,1}\left(\mathrm{U},k\right)\mathcal{F}-\mathrm{E}(\mathrm{U},k)\mathfrak{1}\right)^{-1}\mathfrak{e}_{y}\right\rangle\right|$ (127)
$\leq$ $\frac{\left|\hat{\upsilon}\left(k\right)\right|}{\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)-4\epsilon(\mathrm{e}^{\alpha}-1)}{\sum_{y\in\mathbb{Z}^{2}}}\left|\mathfrak{p}\left(y\right)\right|\mathrm{e}^{-\alpha\left|x-y\right|}$
$\leq$ $\frac{\left|\hat{\upsilon}\left(k\right)\right|\mathrm{e}^{-\alpha\left|x\right|}}{\inf_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\min_{k\in\mathbb{T}^{2}}\mathfrak{g}\left(\mathrm{U},k\right)-4\epsilon(\mathrm{e}^{\alpha}-1)}{\sum_{y\in\mathbb{Z}^{2}}}\left|\mathfrak{p}\left(y\right)\right|\mathrm{e}^{\alpha\left|y\right|}$
for all $x\in\mathbb{Z}^{2}$, $k\in\mathbb{T}^{2}$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$. By choosing $\alpha$ sufficiently small (more precisely $\alpha\leq\alpha_{0}$), we can assume, without loss of generality, that the above sum is finite, see (12). This completes the proof, because the Fourier transform $\hat{\upsilon}$ of $\upsilon$ is a continuous function on the torus $\mathbb{T}^{2}$, which is compact, and is consequently bounded.
4.5 Scattering Channels
4.5.1 Unbound pair scattering channel
Recall that $\mathfrak{H}_{f}$ is defined by (51) and $\mathrm{H}_{f}$ is the operator defined by (52) for any $\mathrm{V}\in\mathbb{R}_{0}^{+}$ and absolutely summable function $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$.
For any operator $Y$ acting on a Hilbert space $\mathcal{Y}$, $P_{\mathrm{ac}}(Y)$ denotes the orthogonal projection on the absolutely continuous space of $Y$, defined by (47). In order to show the existence of a unbound pair scattering channel, we need the following technical lemma:
Lemma 4.20 (Absolute continuous space of fermionic Hamiltonians)
For any $\mathrm{V}\in\mathbb{R}_{0}^{+}$ and every absolutely summable function $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$, the orthogonal projection $P_{\mathrm{ac}}(\mathrm{H}_{f})$ on the absolutely continuous space of $\mathrm{H}_{f}$, defined by (47), is equal to $\mathfrak{1}$.
Proof. Take any $\psi\in\mathfrak{H}_{f}$ and observe from Corollary 5.5 that
$\psi\left(k\right)\in\mathrm{ran}\left(P_{\mathrm{ac}}\left(M_{\mathfrak{f}\left(k\right)}\right)\right),\qquad k\in\mathbb{T}^{2}\ ,$
where $M_{\mathfrak{f}\left(k\right)}$ is the fiber Hamiltonian defined as the multiplication operator associated with the continuous function $\mathfrak{f}(k)\in C(\mathbb{T}^{2})$ (see (28)). For any $\mathrm{V}\in\mathbb{R}_{0}^{+}$ and absolutely summable function $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$, the operator defined by (53), that is,
$R\left(\mathrm{V},\mathrm{v}\right)\doteq{\sum\limits_{x\in\mathbb{Z}^{2}}}\,\mathrm{v}\left(x\right)P_{x}+\mathrm{V}P_{0}\in\mathcal{B}\left(L^{2}(\mathbb{T}^{2})\right)\ ,$
is a trace-class operator, where we recall that $P_{x}$ is the orthogonal projection on the one-dimensional subspace $\mathbb{C}\mathfrak{\hat{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. By [34, Theorem 4.4, Chapter X], it follows in this case that
$\psi\left(k\right)\in\mathrm{ran}\left(P_{\mathrm{ac}}\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\right)\ ,\qquad k\in\mathbb{T}^{2}\ .$
Let $B\subseteq\mathbb{R}$ be an arbitrary Borel set with zero Lebesgue measure. By using (52), we deduce that
$\left\langle\psi,\chi_{B}\left(M_{\mathfrak{f}}\right)\psi\right\rangle_{\mathfrak{H}_{f}}$ $=\left\langle\psi,\left({\int_{\mathbb{T}^{2}}^{\oplus}}\chi_{B}\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\,\nu\left(\mathrm{d}k\right)\right)\psi\right\rangle_{\mathfrak{H}_{f}}$
$={\int_{\mathbb{T}^{2}}}\left\langle\psi\left(k\right),\chi_{B}\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\psi\left(k\right)\right\rangle_{L^{2}\left(\mathbb{T}^{2}\right)}\,\nu\left(\mathrm{d}k\right)=0\text{ }.$
For the first equality note that we apply Theorem 5.3 (iii).
The following results imply that the dynamic generated by the Hamiltonian $H$ (i.e., included the exchange interaction and extended Hubbard repulsions) asymptotically far in the past or future approaches the purely fermionic dynamics for two unbound fermions. This is of course physically expected, since all interaction strengths get weak as the distance between the fermions increases. This is a consequence of the next assertions.
To shorten the notation, for any $\mathrm{V}\in\mathbb{R}_{0}^{+}$ and every absolutely summable function $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$, we define the Hamiltonian
$H^{(1)}\equiv H^{(1)}\left(\mathrm{V},\mathrm{v}\right)\doteq{\int_{\mathbb{T}^{2}}^{\oplus}}H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\nu\left(\mathrm{d}k\right)\in\mathcal{B}\left(L^{2}(\mathbb{T}^{2},\mathcal{H})\right)$
with
$H^{(1)}\left(k\right)\doteq M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\in\mathcal{B}\left(\mathcal{H}\right)\ ,\qquad k\in\mathbb{T}^{2}\ .$ (128)
Here, $M_{\mathfrak{f}\left(k\right)}$, $R\left(\mathrm{V},\mathrm{v}\right)$ and $A_{2,2}(k)$ are respectively the multiplication operator associated with the continuous function $\mathfrak{f}(k)\in C(\mathbb{T}^{2})$ (see (28)), the trace-class operator (53) and the operator defined on $\mathbb{C}$ by (39). We start with the unbounded pair scattering channel in each fiber:
Lemma 4.21 (Fiberwise unbound pair (scattering) channel)
For any $\mathrm{V}\in\mathbb{R}_{0}^{+}$, every absolutely summable function $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ and all $k\in\mathbb{T}^{2}$, the wave operators
$W^{\pm}\left(A\left(k\right),H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)=s-{\lim\limits_{t\rightarrow\pm\infty}}\mathrm{e}^{itA\left(k\right)}\mathrm{e}^{-itH^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)}\ ,$
as defined by Equation (49), exist and are complete with
$\mathrm{ran}\left(W^{\pm}\left(A\left(k\right),H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)\right)=\mathrm{ran}\left(P_{\mathrm{ac}}\left(H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)\right)\ .$
Proof. Note that $P_{\mathrm{ac}}(\mathrm{H}_{f})=\mathfrak{1}$, thanks to Lemma 4.20. By (49), it justifies the strong limit given in the lemma. As we discuss in the proof of Proposition 4.1, for any $k\in\mathbb{T}^{2}$, the operator $A(k)$ is the sum of $M_{\mathfrak{f}\left(k\right)}\oplus A_{2,2}(k)$ and a compact operator $T$. In fact, as the function $\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ (defining the fiber Hamiltonian $A(k)$) is, absolutely summable (see (7)), the operator difference $T$ is even trace-class. As explained in Lemma 4.20, $R\left(\mathrm{V},\mathrm{v}\right)$ is also a trace-class operator, because $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ is absolutely summable, again by assumption. By using the Kato-Rosenblum theorem [35, Theorem XI.8], it thus follows that the wave operators
$W^{\pm}\left(A\left(k\right),H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)=W^{\pm}\left(A\left(k\right),\left(M_{\mathfrak{f}\left(k\right)}+R\left(\mathrm{V},\mathrm{v}\right)\right)\oplus A_{2,2}\left(k\right)\right)$
exist and are complete for every $k\in\mathbb{T}^{2}$.
We are now in a position to prove Theorem 3.11. Recall that $\mathfrak{U}:\mathfrak{H}_{f}\rightarrow L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)$ is the operator defined by (54), while $A_{2,2}(k)$ and $H^{(1)}\left(k\right)$ are respectively defined by (39) and (128). The definition of wave operators $W^{\pm}$ are given by Equations (48)–(49).
Theorem 4.22 (Unbound pair (scattering) channel)
For any $\mathrm{V}\in\mathbb{R}_{0}^{+}$ and every absolutely summable function $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$,
$W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)=\left({\int_{\mathbb{T}^{2}}^{\oplus}}W^{\pm}\left(A\left(k\right),H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}$
with
$\mathrm{ran}\left(W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\right)=\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\,\nu\left(\mathrm{d}k\right)\ .$
Proof. For almost every $k\in\mathbb{T}^{2}$ and every $\psi\in\mathfrak{H}_{f}$,
$\left(H^{(1)}\mathfrak{U}\psi\right)\left(k\right)=H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\left(\mathfrak{U}\psi\right)\left(k\right)=(H^{(1)}\left(k\right)\psi\left(k\right),0)=\left(\mathfrak{U}\mathrm{H}_{f}\psi\right)\left(k\right)\ .$
In other words, $\mathfrak{U}$ is an intertwining operator for $\mathrm{H}_{f}$ and $H^{(1)}$ and hence, for their respective complex exponential:
$\mathfrak{U}\mathrm{e}^{-it\mathrm{H}_{f}}=\mathrm{e}^{-itH^{(1)}}\mathfrak{U}\ ,\qquad t\in\mathbb{R}\ .$ (129)
We also observe that, for any $z\in\mathbb{C}$ and any Borel set $B\subseteq\mathbb{R}$ containing $\mathfrak{b}(k)\in\mathbb{R}$ (see (27)),
$\left\langle z,\chi_{B}\left(A_{2,2}\left(k\right)\right)z\right\rangle_{\mathbb{C}}=\left|z\right|^{2}\chi_{B}\left(\mathfrak{b}\left(k\right)\right)\neq 0$
even if the Lebesgue measure of $B$ is zero. This last observation, together with Remark 3.10 and Lemma 4.20, yields
$\mathrm{ran}\left(P_{\mathrm{ac}}\left(H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)\right)=L^{2}\left(\mathbb{T}^{2}\right)\oplus\left\{0\right\}\ ,\qquad k\in\mathbb{T}^{2}\ .$ (130)
In particular,
${\int_{\mathbb{T}^{2}}^{\oplus}}P_{\mathrm{ac}}\left(H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)\mathfrak{U=U}\ .$ (131)
We can then apply Proposition 5.6 together with Lemmata 4.20, 4.21, Theorem 5.3 and Equations (129)–(131) to arrive at
$W^{\pm}\left(UHU^{\ast},M_{\mathfrak{f}};\mathfrak{U}\right)$ $\doteq s-\lim_{t\rightarrow\mp\infty}\mathrm{e}^{itUHU^{\ast}}\mathfrak{U}\mathrm{e}^{-it\mathrm{H}_{f}}P_{\mathrm{ac}}\left(\mathrm{H}_{f}\right)$
$=s-\lim_{t\rightarrow\mp\infty}\mathrm{e}^{itUHU^{\ast}}\mathfrak{U}\mathrm{e}^{-it\mathrm{H}_{f}}$
$=s-\lim_{t\rightarrow\mp\infty}\mathrm{e}^{itUHU^{\ast}}\mathrm{e}^{-itH^{(1)}}\mathfrak{U}$
$=s-\lim_{t\rightarrow\mp\infty}\left({\int_{\mathbb{T}^{2}}^{\oplus}}\mathrm{e}^{itA\left(k\right)}\,\nu\left(\mathrm{d}k\right)\right)\left({\int_{\mathbb{T}^{2}}^{\oplus}}\mathrm{e}^{-it\left(H^{(1)}\left(k\right)\,\oplus\,A_{2,2}\left(k\right)\right)}\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}$
$=s-\lim_{t\rightarrow\mp\infty}\left({\int_{\mathbb{T}^{2}}^{\oplus}}\mathrm{e}^{itA\left(k\right)}\mathrm{e}^{-it(H^{(1)}\left(k\right)\,\oplus\,A_{2,2}\left(k\right))}P_{\mathrm{ac}}\left(H^{(1)}\left(k\right)\,\oplus\,A_{2,2}\left(k\right)\right)\,\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}$
$=\left({\int_{\mathbb{T}^{2}}^{\oplus}}W^{\pm}\left(A\left(k\right),H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}\ .$
Note that Lemma 4.21 combined with (130) implies that
$\mathrm{ran}\left(W^{\pm}\left(A\left(k\right),H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)\right)=L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\ .$
In particular,
$\mathrm{ran}\left(W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\right)=\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\,\nu\left(\mathrm{d}k\right)\ .$
Observe that Lemma 5.1 allows one to write
$\mathrm{e}^{itX}\mathrm{e}^{i\left(s-t\right)\left(X+Y\right)}\mathrm{e}^{-isX}\ ,\qquad s,t\in\mathbb{R}\ ,$
as a Dyson series for all bounded operators $X,Y$. This can be applied to $X=H^{(1)}$ and $Y=UHU^{\ast}-X$, or in each fiber $k\in\mathbb{T}^{2}$ to $X=H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)$ and $Y=A\left(k\right)-X$. When $\mathrm{U}=\mathrm{V}\in\mathbb{R}_{0}^{+}$ and $\mathrm{v}=\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ in $H^{(1)}$, this result is particularly advantageous because the operator family $(Y_{t})_{t\in\mathbb{R}}$ appearing in Lemma 5.1 can be represented in a relatively simple way in this situation:
Lemma 4.23 (Finite-time scattering and wave operators)
For $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and all $s,t\in\mathbb{R}$,
$\mathrm{e}^{itH^{(1)}\left(\mathrm{U},\mathrm{u}\right)}\mathrm{e}^{i\left(s-t\right)UHU^{\ast}}\mathrm{e}^{-isH^{(1)}\left(\mathrm{U},\mathrm{u}\right)}\\ =\int_{\mathbb{T}^{2}}^{\oplus}\left(\begin{array}[]{cc}\cos_{\succ}\left(B_{1,2}\left(k\right)B_{2,1}\left(k\right);s,t\right)&-i\sin_{\succ}\left(B_{1,2}\left(k\right)B_{2,1}\left(k\right);s,t\right)\\ -i\sin_{\succ}\left(B_{2,1}\left(k\right)B_{1,2}\left(k\right);s,t\right)&\cos_{\succ}\left(B_{2,1}\left(k\right)B_{1,2}\left(k\right);s,t\right)\end{array}\right)\nu\left(\mathrm{d}k\right)\ ,$
where $B_{1,2}\left(k\right)$ and $B_{2,1}\left(k\right)$ are the operator families defined by (56) for any $k\in\mathbb{T}^{2}$, while $\cos_{\succ}$ and $\sin_{\succ}$ are respectively defined by (57) and (LABEL:sinus-operator).
Proof. Let $H^{(1)}\equiv H^{(1)}\left(\mathrm{U},\mathrm{u}\right)$. We infer from Lemma 5.1 applied to $X=H^{(1)}$ and $Y=UHU^{\ast}-X$ that
$\mathrm{e}^{itH^{(1)}}\mathrm{e}^{i\left(s-t\right)UHU^{\ast}}\mathrm{e}^{-isH^{(1)}}=V_{t,s}\doteq\mathbf{1}+\sum_{n=1}^{\infty}\left(-i\right)^{n}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{n-1}}\mathrm{d}\tau_{n}B^{(\tau_{1})}\cdots B^{(\tau_{n})}\ ,$
with $(B^{(t)})_{t\in\mathbb{R}}\subseteq\mathcal{B}(L^{2}(\mathbb{T}^{2},\mathcal{H}))$ being the norm-continuous family defined by
$B^{(t)}$ $=$ $\mathrm{e}^{itH^{(1)}}\left(UHU^{\ast}-H^{(1)}\right)\mathrm{e}^{-itH^{(1)}}$
$=$ $\int_{\mathbb{T}^{2}}^{\oplus}\left(\begin{array}[]{cc}\mathrm{e}^{itA_{1,1}\left(\mathrm{U},k\right)}&0\\ 0&\mathrm{e}^{itA_{2,2}\left(k\right)}\end{array}\right)\left(\begin{array}[]{cc}0&A_{1,2}\left(k\right)\\ A_{2,1}\left(k\right)&0\end{array}\right)\left(\begin{array}[]{cc}\mathrm{e}^{-itA_{1,1}\left(\mathrm{U},k\right)}&0\\ 0&\mathrm{e}^{-itA_{2,2}\left(k\right)}\end{array}\right)\nu\left(\mathrm{d}k\right)$
$=$ $\int_{\mathbb{T}^{2}}^{\oplus}\begin{pmatrix}0&B_{1,2}^{(t)}\left(k\right)\\[5.0pt] B_{2,1}^{(t)}\left(k\right)&0\end{pmatrix}\nu\left(\mathrm{d}k\right)\ ,$
the operators $B_{2,1}^{(t)}\left(k\right)$ and $B_{1,2}^{(t)}\left(k\right)$ being defined by (56) for any $t\in\mathbb{R}$ and $k\in\mathbb{T}^{2}$. Then, one combines explicit computations together with Proposition 5.6 and 5.7 to arrive at the assertion. Notice that the above integrals are Riemann ones, $(B^{(t)})_{t\in\mathbb{R}}$ being a continuous family in the Banach space $\mathcal{B}(L^{2}(\mathbb{T}^{2},\mathcal{H}))$.
4.5.2 Bound pair scattering channel
We start by studying the wave operator (48) with respect to the operators $X=UHU^{\ast}$ and $Y=M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$ (60), the identification operator $J$ being $\mathfrak{P}_{\mathrm{U}}$ (59) for any fixed Hubbard coupling constant $\mathrm{U}\in\mathbb{R}_{0}^{+}$.
Proposition 4.24 (Wave operators in the bound pair channel)
Let $h_{b}\in[0,1/2]$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$. Then, $UHU^{\ast}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$ and for every bounded continuous function $f\in C_{b}(\mathbb{R})$,
$\,f\left(UHU^{\ast}\right)\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}{\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\mathrm{U},k\right)\right)\nu\left(\mathrm{d}k\right)\text{ }.$
Proof. Using Proposition 2.1 and Theorem 4.7, we note that, for any $\varphi\in L^{2}(\mathbb{T}^{2})$ and almost every $k\in\mathbb{T}^{2}$,
$\left(UHU^{\ast}\mathfrak{P}_{\mathrm{U}}\varphi\right)\left(k\right)=A\left(k\right)\left(\mathfrak{P}_{\mathrm{U}}\varphi\right)\left(k\right)=\mathrm{E}\left(\mathrm{U},k\right)\varphi\left(k\right)\left\|\Psi\left(\mathrm{U},k\right)\right\|^{-1}\Psi\left(\mathrm{U},k\right)=\left(\mathfrak{P}_{\mathrm{U}}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\varphi\right)\left(k\right)\ ,$
i.e., $UHU^{\ast}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$, keeping in mind Equations (59) and (60). We then obtain the last assertion by using the Stone-Weierstrass theorem and the spectral theorem.
We now study the (dressed) bound pair channel of lowest energy in the hard-core limit. This is a consequence of the following assertion:
Proposition 4.25 (Bound pair channel in the hard-core limit)
Fix $h_{b}\in[0,1/2]$. Then,
$s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\mathfrak{P}(\mathrm{U})=\mathfrak{P}_{\infty}$ (133)
and for every bounded continuous function $f\in C_{b}(\mathbb{R})$,
$s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,f\left(UHU^{\ast}\right)\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\infty}{\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\infty,k\right)\right)\nu\left(\mathrm{d}k\right)\text{ }.$
Proof. Fix $\varphi\in L^{2}(\mathbb{T}^{2})$. For any $\mathrm{U}\in\mathbb{R}_{0}^{+}$ and almost every $k\in\mathbb{T}^{2}$, one has that
$\left\|\mathfrak{P}_{\mathrm{U}}\varphi\left(k\right)-\mathfrak{P}_{\infty}\varphi\left(k\right)\right\|\leq 2\|\Psi\left(\infty,k\right)\|^{-1}\|\Psi\left(\mathrm{U},k\right)-\Psi\left(\infty,k\right)\|\left|\varphi\left(k\right)\right|\ .$
Thus, by Lebesgue’s dominated convergence theorem, we arrive at
${\lim\limits_{\mathrm{U}\rightarrow\infty}}\,\left\|\mathfrak{P}_{\mathrm{U}}\varphi\left(k\right)-\mathfrak{P}_{\infty}\varphi\right\|_{L^{2}\left(\mathbb{T}^{2},\mathcal{H}\right)}^{2}={\lim\limits_{\mathrm{U}\rightarrow\infty}\int_{\mathbb{T}^{2}}}\left\|\mathfrak{P}_{\mathrm{U}}\varphi\left(k\right)-\mathfrak{P}_{\infty}\varphi\left(k\right)\right\|^{2}\,\nu\left(\mathrm{d}k\right)=0\text{ }.$
Take now a bounded continuous function $f\in C_{b}(\mathbb{R})$. In particular, there exists $L\in\mathbb{R}^{+}$ such that
$\sup_{\mathrm{U}\in\mathbb{R}_{0}^{+}}\sup_{k\in\mathbb{T}^{2}}\left|f\left(\mathrm{E}\left(\mathrm{U},k\right)\right)\right|\leq L\ .$
By Theorem 4.14 (i) and continuity of the function $f$, one has that
${\lim_{\mathrm{U}\rightarrow\infty}}\,f\left(\mathrm{E}\left(\mathrm{U},k\right)\right)=f\left(\mathrm{E}\left(\infty,k\right)\right)\ ,\text{$\qquad$}k\in\mathbb{T}^{2}\ .$
Moreover,
$\mathbb{T}^{2}\ni k\longmapsto f\left(\mathrm{E}\left(\mathrm{U},k\right)\right)\in\mathcal{L}(\mathbb{C})(=\mathcal{B}(\mathbb{C}))$
is a composition of continuous functions and is in particular strongly measurable. Using Proposition 5.6 (or Lebesgue’s dominated convergence theorem) we arrive at the limit
$s-{\lim\limits_{\mathrm{U}\rightarrow\infty}\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\mathrm{U},k\right)\right)\nu\left(\mathrm{d}k\right)={\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\infty,k\right)\right)\nu\left(\mathrm{d}k\right)\ .$ (134)
By using Equations (133), (134), Proposition 4.24, Theorem 5.3 (iii) and the fact that $\|\mathfrak{P}(\mathrm{U})\|_{\mathrm{op}}=1$ for all $\mathrm{U}\in\mathbb{R}_{0}^{+}$, we find that
$s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,f\left(UHU^{\ast}\right)\mathfrak{P}_{\mathrm{U}}$ $=$ $s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,\mathfrak{P}_{\mathrm{U}}f\left(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\right)$
$=$ $s-{\lim\limits_{\mathrm{U}\rightarrow\infty}\mathfrak{P}_{\mathrm{U}}\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\mathrm{U},k\right)\right)\nu\left(\mathrm{d}k\right)$
$=$ $\mathfrak{P}_{\infty}{\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\infty,k\right)\right)\nu\left(\mathrm{d}k\right)\ .$
5 Appendix
5.1 The Unitary Transformation
We give here the explicit definition of the unitary transformation
$U\doteq U_{2}U_{1}\oplus\mathcal{F}:\mathfrak{H}\longrightarrow{\int_{\mathbb{T}^{2}}^{\oplus}}L^{2}\left(\mathbb{T}^{2},\mathbb{C},\nu\right)\oplus\mathbb{C}\,\nu(\mathrm{d}k)$
of Proposition 2.1. Here, $\mathcal{F}$ is the Fourier transform (26) on $\ell^{2}(\mathbb{Z}^{2})$, while the fermionic part of $U$ is the composition of the following unitary linear transformations:
$\begin{array}[]{cccccccc}U_{1}:&\mathfrak{h}_{0}&\rightarrow&\ell^{2}\left(\mathbb{Z}^{2}\times\mathbb{Z}^{2}\right)&\rightarrow&\ell^{2}\left(\mathbb{Z}^{2}\times\mathbb{Z}^{2}\right)&\rightarrow&\ell^{2}\left(\mathbb{Z}^{2}\right)\otimes\ell^{2}\left(\mathbb{Z}^{2}\right)\\ &\mathfrak{e}_{(x,\uparrow)}\wedge\mathfrak{e}_{(y,\downarrow)}&\mapsto&\mathfrak{e}_{(x,y)}&\mapsto&\mathfrak{e}_{(x,x-y)}&\mapsto&\mathfrak{e}_{x}\otimes\mathfrak{e}_{x-y}\end{array}$
and
$\begin{array}[]{cccccc}U_{2}:&\ell^{2}\left(\mathbb{Z}^{2}\right)\otimes\ell^{2}\left(\mathbb{Z}^{2}\right)&\rightarrow&L^{2}\left(\mathbb{T}^{2}\right)\otimes L^{2}\left(\mathbb{T}^{2}\right)&\rightarrow&\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\nu\left(\mathrm{d}k\right)\\ &\mathfrak{e}_{x}\otimes\mathfrak{e}_{x-y}&\mapsto&\mathfrak{\hat{e}}_{x}\otimes\mathfrak{\hat{e}}_{x-y}&\mapsto&\mathfrak{\hat{e}}_{x}\left(\cdot\right)\mathfrak{\hat{e}}_{x-y}\end{array}\ .$
Because $\{\mathfrak{e}_{(x,\uparrow)}\wedge\mathfrak{e}_{(y,\downarrow)}\}_{x,y\in\mathbb{Z}^{2}}$, $\{\mathfrak{e}_{(x,y)}\}_{x,y\in\mathbb{Z}^{2}}$ and $\{\mathfrak{e}_{x}\otimes\mathfrak{e}_{y}\}_{x,y\in\mathbb{Z}^{2}}$ are orthonormal bases and $(x,y)\mapsto(x,x-y)$ is a bijection on $\mathbb{Z}^{2}\times\mathbb{Z}^{2}$, $U_{1}$ is well-defined as a composition of three unitary linear transformations. Note also that the last unitary linear transformation defining $U_{2}$ is defined as in Proposition 5.8, while the first one defining $U_{2}$ is the tensor product $\mathcal{F}\otimes\mathcal{F}$ of the Fourier transform $\mathcal{F}$ on $\ell^{2}(\mathbb{Z}^{2})$.
5.2 Non-autonomous Evolution Equations and Scattering Theory
This section collects simple results on wave and scattering operators (48)–(50) for bounded Hamiltonians, related to their approximation by Dyson series. We start with the following elementary lemma, resulting from the theory of non-autonomous evolution equations:
Lemma 5.1 (Finite-time scattering and wave operators)
For any self-adjoint $X,Y\in\mathcal{B}\left(\mathcal{X}\right)$ acting on a Hilbert space $\mathcal{X}$ and all $s,t\in\mathbb{R}$,
$\mathrm{e}^{itX}\mathrm{e}^{i\left(s-t\right)\left(X+Y\right)}\mathrm{e}^{-isX}=\mathbf{1}+\sum_{n=1}^{\infty}\left(-i\right)^{n}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{n-1}}\mathrm{d}\tau_{n}Y_{\tau_{1}}\cdots Y_{\tau_{n}}$ (135)
with $(Y_{t})_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathcal{X})$ being the norm-continuous family
$Y_{t}\doteq\mathrm{e}^{itX}Y\mathrm{e}^{-itX}\ ,\qquad t\in\mathbb{R}\ .$ (136)
Proof. We compute that, for any $s,t\in\mathbb{R}$,
$\partial_{t}\left\{\mathrm{e}^{itX}\mathrm{e}^{i\left(s-t\right)\left(X+Y\right)}\mathrm{e}^{-isX}\right\}=-i\left(\mathrm{e}^{itX}Y\mathrm{e}^{-itX}\right)\left(\mathrm{e}^{itX}\mathrm{e}^{i\left(s-t\right)\left(X+Y\right)}\mathrm{e}^{-isX}\right)$
as well as
$\partial_{s}\left\{\mathrm{e}^{itX}\mathrm{e}^{i\left(s-t\right)\left(X+Y\right)}\mathrm{e}^{-isX}\right\}=\left(\mathrm{e}^{itX}\mathrm{e}^{i\left(s-t\right)\left(X+Y\right)}\mathrm{e}^{-isX}\right)\left(i\mathrm{e}^{isX}Y\mathrm{e}^{-isX}\right)$
both in $\mathcal{B}(\mathcal{X})$. In other words, the family
$V_{t,s}\doteq\mathrm{e}^{itX}\mathrm{e}^{i\left(s-t\right)\left(X+Y\right)}\mathrm{e}^{-isX}\ ,\qquad s,t\in\mathbb{R}\ ,$ (137)
of (uniformly) bounded operators is a norm-continuous two-parameter family of unitary operators solving the non-autonomous evolution equations
$\forall s,t\in\mathbb{R}:\qquad\partial_{t}Z_{t,s}=-iY_{t}Z_{t,s}\ ,\qquad\partial_{s}Z_{t,s}=iZ_{t,s}Y_{s}\ ,\qquad Z_{s,s}=\mathbf{1}\ ,$ (138)
in $\mathcal{B}(\mathcal{X})$, where $(Y_{t})_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathcal{X})$ is the norm-continuous family defined by (136). As is well-known, there is a unique solution to this non-autonomous evolution equations (138), which is given by the Dyson series (135). This series is absolutely summable in $\mathcal{B}(\mathcal{X})$. Notice that the integrals appearing in it are Riemann integrals, for their arguments are continuous functions taking values in a Banach space.
Corollary 5.2 (Approximation of scattering and wave operators)
Let $X,Y\in\mathcal{B}\left(\mathcal{X}\right)$ be two self-adjoint operators acting on a Hilbert space $\mathcal{X}$. Assume that the waves operators
$W^{\pm}\left(X+Y,X\right)\doteq s-{\lim\limits_{t\rightarrow\pm\infty}}\mathrm{e}^{it\left(X+Y\right)}\mathrm{e}^{-itX}P_{\mathrm{ac}}\left(X\right)$
exist. Let $\varepsilon\in\mathbb{R}^{+}$. Then:
  1. i.)
    For any $\varphi\in\mathrm{ran}\left(P_{\mathrm{ac}}\left(X\right)\right)$, there is $T>0$ such that
    $T<t\implies\left\|\left(W^{+}\left(X+Y,X\right)-V_{0,t}\right)\varphi\right\|_{\mathcal{X}}\leq\varepsilon\ ,$
    whereas
    $t<-T\implies\left\|\left(W^{-}\left(X+Y,X\right)-V_{0,t}\right)\varphi\right\|_{\mathcal{X}}\leq\varepsilon\ .$
  2. ii.)
    For any $\varphi,\psi\in\mathrm{ran}\left(P_{\mathrm{ac}}\left(X\right)\right)$, there is $T>0$ such that
    $\left\langle\psi,S\left(X+Y,X\right)\varphi\right\rangle_{\mathcal{X}}=\left\langle\psi,W^{+}\left(X+Y,X\right)^{\ast}W^{-}\left(X+Y,X\right)\varphi\right\rangle_{\mathcal{X}}=\left\langle\psi,V_{t,s}\varphi\right\rangle_{\mathcal{X}}+\mathcal{O}\left(\varepsilon\right)$
    uniformly for $s<-T<T<t$.
Here,
$V_{t,s}\doteq\mathbf{1}+\sum_{n=1}^{\infty}\left(-i\right)^{n}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{n-1}}\mathrm{d}\tau_{n}Y_{\tau_{1}}\cdots Y_{\tau_{n}}\ ,$
the norm-continuous family $(Y_{t})_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathcal{X})$ being defined by (136).
Proof. Assertion (i) is a direct consequence of Lemma 5.1. Concerning the scattering operator, remark in particular that, for any $r,s,t\in\mathbb{R}$, $V_{t,s}V_{s,r}=V_{t,r}$ and $V_{t,s}^{\ast}=V_{s,t}$. Given $\varphi,\psi\in\mathrm{ran}\left(P_{\mathrm{ac}}\left(Y\right)\right)$, we have that
$\left|\left\langle\psi,\left(S\left(X+Y,X\right)-V_{t,s}\right)\varphi\right\rangle_{\mathcal{X}}\right|\\ =\left|\left\langle\psi,\left(W^{+}\left(X+Y,X\right)^{\ast}W^{-}\left(X+Y,X\right)-V_{0,t}^{\ast}V_{0,s}\right)\varphi\right\rangle_{\mathcal{X}}\right|\\ =\left|\left\langle\psi,\left(W^{+}\left(X+Y,X\right)-V_{0,t}\right)^{\ast}W^{-}\left(X+Y,X\right)\varphi\right\rangle_{\mathcal{X}}\right.\\ \left.+\left\langle\psi,V_{0,t}^{\ast}\left(W^{-}\left(X+Y,X\right)-V_{0,s}\right)\varphi\right\rangle_{\mathcal{X}}\right|\\ \leq\left\|\left(W^{+}\left(X+Y,X\right)-V_{0,t}\right)\psi\right\|_{\mathcal{X}}\left\|W^{-}\left(X+Y,X\right)\varphi\right\|_{\mathcal{X}}\\ +\left\|\left(W^{-}\left(X+Y,X\right)-V_{0,s}\right)\varphi\right\|_{\mathcal{X}}\left\|\psi\right\|_{\mathcal{X}}\ .$
Assertion (ii) therefore follows from assertion (i).
5.3 Constant Fiber Direct Integrals
For more details, we refer to [32, Section XIII.16] as well as [39] for the general theory.
Let $(\mathcal{X},\mu)$ be any semifinite measure space and $\mathcal{Y}$ any separable Hilbert space. The constant fiber direct integral of $\mathcal{Y}$ over $\mathcal{X}$ is, by definition, the Hilbert space
$\int_{\mathcal{X}}^{\oplus}\mathcal{Y}\,\mu\left(\mathrm{d}x\right)\equiv L^{2}\left(\mathcal{X},\mathcal{Y},\mu\right)\doteq\left\{F\in\mathcal{Y}^{\mathcal{X}}\,:\,\left\|F\left(\cdot\right)\right\|_{\mathcal{Y}}^{2}\in L^{1}\left(\mathcal{X},\mu\right)\right\}$
of equivalence classes of square-integrable $\mathcal{Y}$–valued functions with scalar product [^14] The scalar product is well-defined, by the polarization identity and the Cauchy-Schwarz inequality. See for instance [40, Section 7.3.2].
$\left\langle\varphi,\psi\right\rangle\equiv\left\langle\varphi,\psi\right\rangle_{L^{2}\left(\mathcal{X},\mathcal{Y},\mu\right)}\doteq\int_{\mathcal{X}}\left\langle\varphi\left(x\right),\psi\left(x\right)\right\rangle_{\mathcal{Y}}\,\mu\left(\mathrm{d}x\right)\ ,\qquad\varphi,\psi\in L^{2}\left(\mathcal{X},\mathcal{Y},\mu\right)\ ,$
and the pointwise vector space operations
$\left(\varphi+\psi\right)\left(x\right)=\varphi\left(x\right)+\psi\left(x\right),\text{$\qquad$}\left(\text{$\alpha$}\varphi\right)\left(x\right)=\text{$\alpha$}\varphi\left(x\right)\ ,\qquad\text{$\alpha\in\mathbb{C}$},\ \varphi,\psi\in L^{2}\left(\mathcal{X},\mathcal{Y},\mu\right)\ .$
If $\mathcal{Y}=\mathbb{C}$ then we use the shorter notation $L^{2}\left(\mathcal{X},\mu\right)\equiv L^{2}\left(\mathcal{X},\mathbb{C},\mu\right)$.
A mapping $A:\mathcal{X}\rightarrow\mathcal{B}(\mathcal{Y})$ is strongly measurable whenever the mapping $x\mapsto\left\langle\varphi,A(x)\psi\right\rangle_{\mathcal{Y}}$ from $\mathcal{X}$ to $\mathbb{C}$ is measurable for all $\varphi,\psi\in\mathcal{Y}$. Let $L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$ be the $C^{\ast}$-algebra of equivalence classes of strongly measurable functions $A:\mathcal{X}\rightarrow\mathcal{B}(\mathcal{Y})$ with
$\left\|A\right\|_{\infty}\doteq\mathrm{ess-}\sup\left\{\left\|A\left(x\right)\right\|_{\mathrm{op}}:x\in\mathcal{X}\right\}<\infty\ .$
Here, $\mathrm{ess-}\sup$ denotes the essential supremum and $\left\|\cdot\right\|_{\mathrm{op}}$ stands for the operator norm. If $\mathcal{Y}=\mathbb{C}$ then we use the shorter notation $L^{\infty}\left(\mathcal{X},\mu\right)\equiv L^{\infty}\left(\mathcal{X},\mathbb{C},\mu\right)$.
A bounded operator $D$ on $L^{2}(\mathcal{X},\mathcal{Y},\mu)$ is decomposable if there is $A\in L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$ such that, for all $\psi\in L^{2}(\mathcal{X},\mathcal{Y},\mu)$,
[…]
If such an $A$ exists, then it is unique. Moreover, the mapping $A\mapsto D$ defined by the above equality is a $\ast$-homomorphism which is isometric. See [32, Theorem XIII.83]. The operators $A(x)\in\mathcal{B}(\mathcal{Y})$, $x\in\mathcal{X}$, are called the fibers of $D$ and we write
$D=\int_{\mathcal{X}}^{\oplus}A\left(x\right)\,\mu\left(\mathrm{d}x\right)\text{ }.$
For the reader’s convenience, we now give three essential properties of decomposable operators used in the paper, referring to [32, Theorem XIII.85 (a), (c) and (d)]. Note that $\sigma(X)$ denotes below the spectrum of any operator $X$ acting on some Hilbert space, as is usual.
Theorem 5.3 (Properties of decomposable operators)
Let $D$ be a decomposable operator on $L^{2}(\mathcal{X},\mathcal{Y},\mu)$, the fibers $A(x)\in\mathcal{B}(\mathcal{Y})$, $x\in\mathcal{X}$, of which are all self-adjoint. Then:
  1. i.)
    $D$ is self-adjoint.
  2. ii.)
    $\lambda\in\sigma(D)$ iff, for all $\varepsilon\in\mathbb{R}^{+}$,
    $\mu\left(\left\{x\in\mathcal{X}:\sigma\left(A\left(x\right)\right)\cap\left(\lambda-\varepsilon,\lambda+\varepsilon\right)\neq\emptyset\right\}\right)>0\text{ }.$
  3. iii.)
    For any bounded Borel function $f$ on $\mathbb{R}$, $f(D)$ is decomposable and has fibers $f(A(x))$, $x\in\mathcal{X}$, that is,
    $f(D)=\int_{\mathcal{X}}^{\oplus}f\left(A\left(x\right)\right)\,\mu\left(\mathrm{d}x\right){\ .}$
The above theorem can be used to elegantly prove the following well-known results about multiplication operators $M_{\varphi}$ by any bounded measurable function $\varphi\in L^{\infty}(\mathcal{X},\mu)$, defined for any $\psi\in L^{2}(\mathcal{X},\mu)$, by
[…]
Corollary 5.4 (Properties of multiplication operators)
The multiplication operators $M_{\varphi}$ by $\varphi\in L^{\infty}(\mathcal{X},\mu)$ has the following properties:
  1. i.)
    For any bounded Borel function $f$ on $\mathbb{R}$, one has $f(M_{\varphi})=M_{f\circ\varphi}$.
  2. ii.)
    $\sigma(M_{\varphi})$ is the essential range $\mathrm{ess-im}(\varphi)$ of $\varphi$.
  3. iii.)
    Its operator norm $\|M_{\varphi}\|_{\mathrm{op}}$ is equal to $\|\varphi\|_{\infty}$.
Proof. Noting that $M_{\varphi}$ is a decomposable operator on
$L^{2}(\mathcal{X},\mu)=\int_{\mathcal{X}}^{\oplus}\mathbb{C}\,\mu\left(\mathrm{d}x\right)$
with $\varphi(x)$, seen as a linear operator on $\mathbb{C}$, being its fibers, we can use Theorem 5.3 (iii) to get the equality
$f(M_{\varphi})=\int_{\mathcal{X}}^{\oplus}f\left(\varphi\left(x\right)\right)\,\mu\left(\mathrm{d}x\right)=M_{f\circ\varphi}\text{ }.$
This proves Assertion (i). For the second one, we use that $\sigma(\varphi(x))=\{\varphi(x)\}$ for all $x\in\mathcal{X}$ and thus infer from Theorem 5.3 (ii) that $\lambda\in\sigma(M_{f})$ iff, for all $\varepsilon\in\mathbb{R}^{+}$,
$\mu\left(\left\{x\in\mathcal{X}:\left|\lambda-\varphi\left(x\right)\right|<\varepsilon\right\}\right)>0\text{ }.$
In other words, one gets Assertion (ii). Assertion (iii) is an elementary application of [32, Theorem XIII.83].
Below we study the special case of multiplication operators on $L^{2}(\mathbb{T}^{2},\nu)$, where $\mathbb{T}^{2}\doteq[-\pi,\pi)^{2}$ is the torus and $\nu$ is the normalized Haar measure (21) on $\mathbb{T}^{2}$. It is again an elementary result, used in the paper. To this end, we recall that, for any self-adjoint operator $Y$ acting on a Hilbert space $\mathcal{Y}$, $P_{\mathrm{ac}}(Y)$ is the orthogonal projection onto the absolutely continuous space of $Y$, which is defined by (47).
Corollary 5.5 (Absolutely continuous space of multiplication operators on $L^{2}(\mathbb{T}^{2},\nu)$)
[…]
$\mathrm{ran}\left(P_{\mathrm{ac}}\left(M_{\varphi}\right)\right)=L^{2}(\mathbb{T}^{2},\nu)\ .$
Proof. Given any Borel set $\Omega\subseteq\mathbb{R}$, we deduce from Corollary 5.4 (i) that
$\chi_{\Omega}\left(M_{\varphi}\right)=M_{\chi_{\Omega}\circ\varphi}=M_{\chi_{\varphi^{-1}\left(\Omega\right)}}\ ,$
which in turn implies that, for any $\psi\in L^{2}(\mathbb{T}^{2},\nu)$,
$\left\langle\psi,\chi_{\Omega}\left(M_{\varphi}\right)\psi\right\rangle=\int_{\varphi^{-1}\left(\Omega\right)}\left|\psi\left(k\right)\right|^{2}\nu\left(\mathrm{d}k\right)\ .$
Hence, if $\Omega\subseteq\mathbb{R}$ has zero Lebesgue measure, then under the conditions of the corollary,
$\left\langle\psi,\chi_{\Omega}\left(M_{\varphi}\right)\psi\right\rangle=0\ .$
In other words, for any $\psi\in L^{2}(\mathbb{T}^{2},\nu)$, $\langle\psi,\chi_{(\cdot)}(M_{\varphi})\psi\rangle$ is absolutely continuous with respect to the Lebesgue measure.
We next provide a result on the strong operator convergence and a version of Fubini’s theorem for (constant fiber) direct integrals, which are also used in our proofs.
Proposition 5.6 (Strong operator convergence)
Let $(A_{n})_{n\in\mathbb{N}}$ be any bounded sequence in $L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$. If
$s-\lim_{n\rightarrow\infty}A_{n}\left(x\right)=A\left(x\right),\qquad x\in\mathcal{X}\text{ },$
then $A\in L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$ and
$s-\lim_{n\rightarrow\infty}\int_{\mathcal{X}}^{\oplus}A_{n}\left(x\right)\,\mu\left(\mathrm{d}x\right)=\int_{\mathcal{X}}^{\oplus}A\left(x\right)\,\mu\left(\mathrm{d}x\right)\ .$
Proof. The assertion is well-known, but for the reader’s convenience we give here its complete proof. For any $\varphi,\psi\in\mathcal{Y}$, it follows from the fact that $A_{n}(x)\psi\rightarrow A(x)\psi$ everywhere and the continuity of $\left\langle\varphi,\cdot\right\rangle\in\mathcal{Y}^{\ast}$ that
${\lim_{n\rightarrow\infty}}\,\left\langle\varphi,A_{n}\left(x\right)\psi\right\rangle_{\mathcal{Y}}=\left\langle\varphi,A\left(x\right)\psi\right\rangle_{\mathcal{Y}}\text{ },\qquad x\in\mathcal{X}\text{ }.$
This shows that $A$ is strongly measurable, because the pointwise limit of a sequence of real-valued measurable functions is measurable as well. Now, let
$M\doteq\sup_{n\in\mathbb{N}}\left\|A_{n}\right\|_{\infty}<\infty\ .$
For $\mu$-a.e. $x\in\mathcal{X}$ and any $\varphi\in\mathcal{Y}$,
$\left\|A_{n}\left(x\right)\varphi\right\|_{\mathcal{Y}}\leq\left\|A_{n}\left(x\right)\right\|_{\mathrm{op}}\left\|\varphi\right\|_{\mathcal{Y}}\leq M\left\|\varphi\right\|_{\mathcal{Y}}\text{ },\qquad n\in\mathbb{N}\text{ }.$ (139)
Taking the limit $n\rightarrow\infty$, one thus gets that, for $\mu$-a.e. $x\in\mathcal{X}$ and any $\varphi\in\mathcal{Y}$,
$\left\|A\left(x\right)\varphi\right\|_{\mathcal{Y}}=\lim_{n\rightarrow\infty}\left\|A_{n}\left(x\right)\varphi\right\|_{\mathcal{Y}}\leq M\left\|\varphi\right\|_{\mathcal{Y}}\text{ }.$ (140)
Hence, $M$ is an essential upper bound for $\{\|A(x)\|_{\mathrm{op}}\}_{x\in\mathcal{X}}$ and, therefore, $A\in L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$. Finally, given any element $\varphi\in L^{2}(\mathcal{X},\mathcal{Y},\mu)$, by (139)–(140) and the triangle inequality, we have the estimate
[…]
Since $A_{n}\left(x\right)\varphi\left(x\right)\rightarrow A\left(x\right)\varphi\left(x\right)$ for all $x\in\mathcal{X}$, we can therefore apply Lebesgue’s dominated convergence theorem to conclude that, for any $\varphi\in L^{2}(\mathcal{X},\mathcal{Y},\mu)$,
$\lim_{n\rightarrow\infty}\left\|\left(\int_{\mathcal{X}}^{\oplus}A_{n}\left(x\right)\,\mu\left(\mathrm{d}x\right)\right)\varphi-\left(\int_{\mathcal{X}}^{\oplus}A\left(x\right)\,\mu\left(\mathrm{d}x\right)\right)\varphi\right\|_{L^{2}\left(\mathcal{X},\mathcal{Y},\mu\right)}=0\ .$
Before proving a version of Fubini’s theorem for constant fiber direct integrals, we fix some terminology concerning the Riemann integral: A partition of the interval $[a,b]$ is a finite set $P=\{t_{0}<t_{1}<\cdots<t_{k}\}$ where $t_{0}=a$ and $t_{k}=b$. The norm of the partition $P$ is the number $|P|=\max_{1\leq i\leq k}(t_{i}-t_{i-1})$. A tagged partition is a pair $P^{\ast}=(P,\xi)$ where $P$ is a partition and $\xi=(\xi_{1},\dotsc,\xi_{k})$ is such that $t_{i-1}\leq\xi_{i}<t_{i}$ for every $i=1,\dotsc,k$. If $P^{\ast}$ is a tagged partition of $[a,b]$, the corresponding Riemann sum for $f:[a,b]\rightarrow\mathcal{Z}$, with $\mathcal{Z}$ being a vector space, is
$\Sigma(f;P^{\ast})=\displaystyle{\sum_{i=1}^{k}}\,(t_{i}-t_{i-1})f(\xi_{i})\in\mathcal{Z}\ .$
Proposition 5.7 (Fubini’s Theorem for direct integrals)
Let $A_{(\cdot)}:[a,b]\rightarrow L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$ be a continuous function. Then:
  1. i.)
    The mapping
    $\mathcal{X}\ni x\mapsto{\int_{a}^{b}}A_{t}\left(x\right)\,\mathrm{d}t\in\mathcal{B}(\mathcal{Y})$
    is an element of $L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$;
  2. ii.)
    Then the mapping
    $[a,b]\ni t\mapsto{\int_{\mathcal{X}}^{\oplus}}A_{t}\left(x\right)\,\mu\left(\mathrm{d}x\right)\in\mathcal{B}\left(\int_{\mathcal{X}}^{\oplus}\mathcal{Y}\,\mu\left(\mathrm{d}x\right)\right)$
    is continuous and
    ${\int_{a}^{b}\int_{\mathcal{X}}^{\oplus}}A_{t}\left(x\right)\,\mu\left(\mathrm{d}x\right)\mathrm{d}t={\int_{\mathcal{X}}^{\oplus}\int_{a}^{b}}A_{t}(x)\,\mathrm{d}t\,\mu\left(\mathrm{d}x\right)\text{ }.$
Proof. If $A_{(\cdot)}:[a,b]\rightarrow L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$ is continuous, then so is $A_{(\cdot)}(x):[a,b]\rightarrow\mathcal{B}(\mathcal{Y})$ for $x\in\mathcal{X}$ $\mu$-a.e. For simplicity we may assume that $A_{(\cdot)}(x)$ is even continuous for all $x\in\mathcal{X}$. If fact, as this is true for $x\in\mathcal{X}$ $\mu$-a.e., for some Borel set $\mathcal{X}_{0}\subseteq\mathcal{X}$ with $\mu(\mathcal{X}_{0})=0$, $\mathbf{1}\left[x\notin\mathcal{X}_{0}\right]A_{(\cdot)}(x)$ is continuous for all $x\in\mathcal{X}$. Note that, for all $t\in[a,b]$ the functions $A_{t}$ and $\mathbf{1}\left[(\cdot)\notin\mathcal{X}_{0}\right]A_{t}$ are strongly mensurable and represent the same element (i.e., equivalence class of strongly mensurable functions $\mathcal{X}\rightarrow\mathcal{B}(\mathcal{Y})$) of $L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$. Moreover, as $[a,b]$ is compact, $\{A_{t}\}_{t\in[a,b]}$ is bounded as a subset of the metric space $\left(L^{\infty}(\mathcal{X},\mathcal{Y},\mu),\|\cdot\|_{\infty}\right)$. Thus,
$\left\|\int_{a}^{b}A_{t}(x)\,\mathrm{d}t\right\|_{\mathrm{op}}$$\leq$ $\int_{a}^{b}\|A_{t}(x)\|_{\mathrm{op}}\,\mathrm{d}t\leq\int_{a}^{b}\|A_{t}\|_{\infty}\,\mathrm{d}t$
[…]
Let $P_{n}^{\ast}$ be a tagged partition whose norm of the corresponding partition $P_{n}$ goes to zero as $n\rightarrow\infty$. Then for every $\varphi,\psi\in\mathcal{Y}$ and $x\in\mathcal{X}$ $\mu$-a.e,
$\left\langle\varphi,\left({\int_{a}^{b}}A_{t}(x)\,\mathrm{d}t\right)\psi\right\rangle_{\mathcal{Y}}={\int_{a}^{b}}\langle\varphi,A_{t}(x)\psi\rangle_{\mathcal{Y}}\,\mathrm{d}t={\lim_{n\rightarrow\infty}}\,\Sigma\left(\langle\psi,A_{(\cdot)}(x)\varphi\rangle_{\mathcal{Y}};P_{n}^{\ast}\right)\text{ }.$
For the first equality we used the fact that the Riemann integral commutes with bounded linear transformations. Observing that for $x\in\mathcal{X}$ $\mu$-a.e. the right-hand side is a pointwise limit of a linear combination of continuous (hence Riemann integrable) functions, this proves assertion (i).
Note that the mapping defined in Assertion (ii) is a composition of two continuous functions, namely: $A_{(\cdot)}:[a,b]\rightarrow L^{\infty}(\mathcal{X},\mathcal{Y},\mu)$ and
$L^{\infty}(\mathcal{X},\mathcal{Y},\mu)\ni B\mapsto\int_{\mathcal{X}}^{\oplus}B(x)\,\mu\left(\mathrm{d}x\right)\in\mathcal{B}\left(\int_{\mathcal{X}}^{\oplus}\mathcal{Y}\,\mu\left(\mathrm{d}x\right)\right).$
Given any $\varphi,\psi\in L^{2}(\mathcal{X},\mathcal{Y},\mu)$, observe that the function
$\mathcal{X}\times[a,b]\ni(x,t)\mapsto f\left(x,t\right)=\langle\varphi(x),A_{t}(x)\psi(x)\rangle_{\mathcal{Y}}\in\mathbb{C}$
is measurable when $A_{(\cdot)}(x)\in C([a,b],\mathcal{B}(\mathcal{Y}))$ for any $x\in\mathcal{X}$.
To prove this, for each $n\in\mathbb{N}$, define the function $f_{n}:\mathcal{X}\times[a,b]\rightarrow\mathbb{C}$ by $f_{n}(x,t)=f(x,s_{t})$, with $s_{t}=\min\{m_{t,n}/n,b\}$ and $m_{t,n}\in\mathbb{Z}$ being such that $(m_{t,n}-1)/n\leq t<m_{t,n}/n$. In particular,
$f_{n}(x,t)={\sum_{m\in\mathbb{Z}:m\geq na}}\mathbf{1}\left[t\in n^{-1}[m-1,m)\right]f(x,\min\{m/n,b\})\ .$
Characteristic functions are measurable on $[a,b]$ and $f(\cdot,t)$ is also measurable on $\mathcal{X}$ for every $t\in[a,b]$. So, $f_{n}$ is measurable for all $n\in\mathbb{N}$. It is easy to check that $m_{t,n}/n\rightarrow t$ for any $t\in[a,b]$, which in turn implies that $f_{n}$ pointwise converges to $f$, as $n\rightarrow\infty$, because of the continuity of $f(x,\cdot)$ for any fixed $x\in\mathcal{X}$. The last continuity property is a direct consequence of the assumption $A_{(\cdot)}(x)\in C([a,b],\mathcal{B}(\mathcal{Y}))$ together with elementary estimates using the Cauchy-Schwarz inequality.
As a consequence, $f$ is measurable on $\mathcal{X}\times[a,b]$. Note also that
${\int_{\mathcal{X}}\int_{a}^{b}}|f(x,t)|\,\mathrm{d}t\,\mu(\mathrm{d}x)\leq(b-a)||\varphi||_{L^{2}(\mathcal{X},\mathcal{Y},\mu)}^{2}||\psi||_{L^{2}(\mathcal{X},\mathcal{Y},\mu)}^{2}\sup_{t\in[a,b]}\|A_{t}\|_{\infty}<\infty\text{ },$
thanks to the Cauchy-Schwarz inequality for both spaces $\mathcal{Y}$ and $L^{2}(\mathcal{X},\mathcal{Y},\mu)$. We can then apply (usual) Fubini’s theorem to obtain
$\left\langle\varphi,\left({\int_{\mathcal{X}}^{\oplus}\int_{a}^{b}}A_{t}(x)\,\mathrm{d}t\,\mu(\mathrm{d}x)\right)\psi\right\rangle_{L^{2}(\mathcal{X},\mathcal{Y},\mu)}={\int_{\mathcal{X}}}\left\langle\varphi(x),\left({\int_{a}^{b}}A_{t}(y)\,\mathrm{d}t\right)\psi(x)\right\rangle_{\mathcal{Y}}\,\mu(\mathrm{d}x)=\\[5.0pt] ={\int_{\mathcal{X}}\int_{a}^{b}}\langle\varphi(x),A_{t}(x)\psi(x)\rangle_{\mathcal{Y}}\,\mathrm{d}t\,\mu(\mathrm{d}x)={\int_{a}^{b}\int_{\mathcal{X}}}\langle\varphi(x),A_{t}(x)\psi(x)\rangle_{\mathcal{Y}}\,\mu(\mathrm{d}x)\,\mathrm{d}t=\\[5.0pt] ={\int_{a}^{b}}\left\langle\varphi,\left({\int_{\mathcal{X}}^{\oplus}}A_{t}(x)\,\mu(\mathrm{d}x)\right)\psi\right\rangle_{\mathcal{Y}}\,\mathrm{d}t=\left\langle\varphi,\left({\int_{a}^{b}\int_{\mathcal{X}}^{\oplus}}A_{t}(x)\,\mu(\mathrm{d}x)\,\mathrm{d}t\right)\psi\right\rangle_{L^{2}(\mathcal{X},\mathcal{Y},\mu)}\ .$
As $\varphi,\psi$ are arbitrary, we arrive at Assertion (ii).
We conclude this short account on constant fiber direct integrals by providing a representation of them as tensor products:
Proposition 5.8 (Direct integrals and tensor products)
There is a unique unitary transformation $\mathrm{U}:L^{2}(\mathcal{X},\mu)\otimes\mathcal{Y}\rightarrow L^{2}(\mathcal{X},\mathcal{Y},\mu)$ such that
[…]
Proof. See [39, Proposition 5.2].
5.4 The Birman-Schwinger Principle
There are various versions of the Birman-Schwinger principle in the literature and we give below the precise version that is used in our proofs. To this end, we first define Birman-Schwinger operators: For any operator $T$ acting on some complex vector space, recall that $\rho(T)\subseteq\mathbb{C}$ denotes its resolvent set. Given two (bounded) operators $T,V$ acting on some complex vector space and any $\lambda\in\rho(T)$, we define the associated Birman-Schwinger operator to be
$\mathrm{B}\left(\lambda\right)\equiv\mathrm{B}\left(\lambda,T,V\right)\doteq V\left(T-\lambda\mathfrak{1}\right)^{-1}V\text{ }.$ (141)
It turns out that, for all $\lambda\in\rho(T)$, $1$ is an eigenvalue of $\mathrm{B}\left(\lambda\right)$ iff $\lambda$ is an eigenvalue of $T-V^{2}$:
Lemma 5.9 (The eigenvalues of Birman-Schwinger operators)
Let $T,V$ be two bounded operators acting on a vector space $\mathcal{X}$ over $\mathbb{C}$. Assume that $\lambda\in\rho(T)$ is an eigenvalue of $T-V^{2}$ and let $\{\varphi_{i}\}_{i\in I}$ denote any basis of the corresponding eigenspace. Define $\gamma_{i}\doteq V\varphi_{i}$, $i\in I$. Then,
$\varphi_{i}=\left(T-\lambda\mathfrak{1}\right)^{-1}V^{2}\varphi_{i}=\left(T-\lambda\mathfrak{1}\right)^{-1}V\gamma_{i}\text{ },\qquad i\in I\text{ },$ (142)
and $\{\gamma_{i}\}_{i\in I}$ is a linearly independent set satisfying
$\mathrm{B}\left(\lambda\right)\gamma_{i}=V\varphi_{i}=\gamma_{i}\text{ },\qquad i\in I\text{ }.$ (143)
Proof. Suppose that $\lambda$ is an eigenvalue of $T-V^{2}$ and let $\{\varphi_{i}\}_{i\in I}$ be a basis of the corresponding eigenspace. Set $\gamma_{i}=V\varphi_{i}$, $i\in I$. Then
$\left(T-\lambda\mathfrak{1}\right)\varphi_{i}=\left(T-V^{2}\right)\varphi_{i}+\left(V^{2}-\lambda\mathfrak{1}\right)\varphi_{i}=\lambda\varphi_{i}+\left(V^{2}-\lambda\mathfrak{1}\right)\varphi_{i}=V^{2}\varphi_{i}$
so that (142) holds true. By (142), $\gamma_{i}$ is a non-zero vector for any $i\in I$ since $\varphi_{i}\neq 0$ for all $i\in I$. As linear transformations map a linearly dependent set onto a linearly dependent set, we conclude that $\{\gamma_{i}\}_{i\in I}$ is a linearly independent set. Equation (143) is a direct consequence of (141) and (142).
This last lemma is explicitly used in the proof of Corollary 4.5 and allows meanwhile to prove the Birman-Schwinger principle for eigenvalues. Below, for any operator $T$, we use the notation $\mathcal{E}_{T}(\lambda)$ for the eigenspace associated with the eigenvalue $\lambda$ of $T$.
Theorem 5.10 (Birman-Schwinger)
Let $T,V$ be two linear operators acting on a vector space $\mathcal{X}$ over $\mathbb{C}$ and $\lambda\in\rho(T)$. Then $\lambda$ is an eigenvalue of $T-V^{2}$ iff $1$ is an eigenvalue of $\mathrm{B}(\lambda)\equiv\mathrm{B}(\lambda,T,V)$. In this case,
$\dim\mathcal{E}_{T-V^{2}}\left(\lambda\right)=\dim\mathcal{E}_{\mathrm{B}\left(\lambda\right)}\left(1\right)\text{ },$
i.e., the corresponding (geometric) multiplicities of eigenvalues are equal to each other.
Proof. If $\lambda$ is an eigenvalue of $T-V^{2}$ then Lemma 5.9 implies that $1$ is an eigenvalue of $\mathrm{B}(\lambda)$ and the eigenspace of $\mathrm{B}(\lambda)$ corresponding to the eigenvalue $1$ has at least dimension $|I|=\dim\mathcal{E}_{T-V^{2}}(\lambda)$. Conversely, if $\{\phi_{j}\}_{j\in J}$ is a basis of the eigenspace of $\mathrm{B}(\lambda)$ corresponding to the eigenvalue $1$ then we set
$\psi_{j}\doteq\left(T-\lambda\mathfrak{1}\right)^{-1}V\phi_{j}\text{ },\qquad j\in J\text{ }.$
Then, by (141),
$\phi_{j}=\mathrm{B}\left(\lambda\right)\phi_{j}=V\psi_{j}\text{ },\qquad j\in J\text{ },$
which implies that $\{\psi_{j}\}_{j\in J}$ is a linearly independent set. Thus,
$\left(T-V^{2}\right)\psi_{j}=\left(T-\lambda\mathfrak{1}\right)\psi_{j}+\left(\lambda\mathfrak{1}-V^{2}\right)\psi_{j}=V\phi_{j}+\left(\lambda\mathfrak{1}-V^{2}\right)\psi_{j}=\lambda\psi_{j}\text{ },\qquad j\in J\text{ },$
and, hence, the eigenspace of $T-V^{2}$ corresponding to the eigenvalue $\lambda$ has at least dimension $|J|=\dim\mathcal{E}_{\mathrm{B}(\lambda)}(1)$.
5.5 Combes-Thomas Estimates
We give here a version of the celebrated Combes-Thomas estimates, first proven in 1973 [41], which is well-adapted to our framework. For the non-expert reader, we provide also its proof, which is relatively short and easy to understand in the particular situation we are interested in.
Fix a countable set $\Lambda$ and a pseudometric $d:\Lambda\times\Lambda\rightarrow\mathbb{R}_{0}^{+}$ on $\Lambda$. Let $\ell^{2}(\Lambda)$ be the (separable) Hilbert space of square summable functions $\Lambda\rightarrow\mathbb{C}$. Similar to (5), its canonical orthonormal basis is defined by
$\mathfrak{e}_{x}\left(y\right)\doteq\delta_{x,y}\ ,\qquad x,y\in\Lambda\ ,$
where $\delta_{\mathfrak{i},\mathfrak{j}}$ is the Kronecker delta. For simplicity, as before, we use the shorter notation $\langle\cdot,\cdot\rangle\equiv\langle\cdot,\cdot\rangle_{\ell^{2}(\Lambda)}$ for its scalar product.
For each bounded operator $T\in\mathcal{B}(\ell^{2}(\Lambda))$ and positive parameter $\mu\in\mathbb{R}_{0}^{+}$, we define the quantity
$\mathbf{S}\left(T,\mu\right)\doteq\sup_{x\in\Lambda}\sum_{y\in\Lambda}\left(\mathrm{e}^{\mu d\left(x,y\right)}-1\right)\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|\in\left[0,\infty\right]\ .$ (144)
Compare with (124). By definition of a pseudometric, the function $d$ is symmetric with respect to the variables $x$ and $y$. The same occurs with the factor $|\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\rangle|$, provided that $T$ is self-adjoint. Thus, in this particular case, $\mathbf{S}(T,\mu)$ is equal to
$\mathbf{S}\left(T,\mu\right)=\sup_{y\in\Lambda}\sum_{x\in\Lambda}\left(\mathrm{e}^{\mu d\left(x,y\right)}-1\right)\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|\in\left[0,\infty\right]\ .$ (145)
The lemma below provides an estimate of the operator norm of $T$ in terms of quantities that are similar to (144) and (145):
Lemma 5.11
For any bounded operator $T\in\mathcal{B}(\ell^{2}(\Lambda))$,
$\left\|T\right\|_{{\mathrm{o}\mathrm{p}}}^{2}\leq\left(\sup_{y\in\Lambda}\sum_{x\in\Lambda}\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|\right)\left(\sup_{x\in\Lambda}\sum_{y\in\Lambda}\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|\right)\ .$
Proof. Assume without loss of generality that the above bound is finite for $T\in\mathcal{B}(\ell^{2}(\Lambda))$. Otherwise the assertion would be trivial. Let $V:\ell^{1}(\Lambda)+\ell^{\infty}(\Lambda)\rightarrow\ell^{1}(\Lambda)+\ell^{\infty}(\Lambda)$ be the mapping defined by
$Vf\left(x\right)\doteq\sum_{y\in\Lambda}f\left(y\right)\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\ ,\qquad x\in\Lambda,\ f\in\ell^{1}\left(\Lambda\right)+\ell^{\infty}\left(\Lambda\right)\ .$
If $f\in\ell^{\infty}(\Lambda)$ then $Vf\in\ell^{\infty}(\Lambda)$ because
$\sup_{x\in\Lambda}\sum_{y\in\Lambda}\left|f\left(y\right)\right|\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|\leq\left\|f\right\|_{\infty}\sup_{x\in\Lambda}\sum_{y\in\Lambda}\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|<\infty\ ,$
while, for any $f\in\ell^{1}(\Lambda)$, we also have $Vf\in\ell^{1}(\Lambda)$ because
$\left\|Vf\right\|_{\ell^{1}(\Lambda)}\sum_{x\in\Lambda}\left|Vf\left(x\right)\right|\leq\sum_{x,y\in\Lambda}\left|f\left(y\right)\right|\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|\leq\left\|f\right\|_{\ell^{1}(\Lambda)}\sup_{y\in\Lambda}\sum_{x\in\Lambda}\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|<\infty\ ,$
using Tonelli’s theorem. It then follows from the Riesz-Thorin theorem [48, Theorem 6.27] that, for any function $f\in\ell^{2}(\Lambda)\subseteq\ell^{1}(\Lambda)+\ell^{\infty}(\Lambda)$,
$\left\|Vf\right\|_{\ell^{2}(\Lambda)}^{2}\leq\left\|f\right\|_{\ell^{2}(\Lambda)}^{2}\left(\sup_{y\in\Lambda}\sum_{x\in\Lambda}\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|\right)\left(\sup_{x\in\Lambda}\sum_{y\in\Lambda}\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle\right|\right)\ .$
Finally, we observe that
$\left(Tf\right)\left(x\right)=\left\langle\mathfrak{e}_{x},Tf\right\rangle=\sum_{y\in\Lambda}f\left(y\right)\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{y}\right\rangle=\left(Vf\right)\left(x\right)$
whenever $x\in\Lambda$ and $f\in\ell^{2}(\Lambda)$.
We now state another, well-known, technical lemma, which is given here for completeness. Recall that, here, $\rho(T)\subseteq\mathbb{C}$ and $\sigma(T)\doteq\mathbb{C}\backslash\rho(T)$ respectively denote the resolvent set and the spectrum of any element $T$ in some unital $C^{\ast}$-algebra (like the space of bounded operators on some Hilbert space). Similar to (123), we use the notation
$\Delta\left(\lambda;T\right)\doteq\min\left\{\left|\lambda-a\right|:a\in\sigma(T)\right\}$ (146)
for the distance between a complex number $\lambda\in\mathbb{C}$ and the spectrum $\sigma(T)$ of any element $T$ in some unital $C^{\ast}$-algebra.
Lemma 5.12 (Norm estimates of resolvents)
[…]
$\left\|\left(T+B-\lambda\mathfrak{1}\right)^{-1}\right\|\leq{\frac{1}{\Delta\left(\lambda;T\right)-\left\|B\right\|}}\ .$
Proof. Assume all conditions of the lemma, in particular that $\|B\|<\Delta(\lambda;T)$. Then,
$\left\|\left(T-\lambda\mathfrak{1}\right)^{-1}B\right\|\leq\Delta\left(\lambda;T\right)^{-1}\left\|B\right\|<1$
and using the Neumann series [40, Lemma 4.24] for $-(T-\lambda\mathfrak{1})^{-1}B$, the element $\mathfrak{1}+(T-\lambda\mathfrak{1})^{-1}B$ is invertible with norm bounded by
$\left\|\left(\mathfrak{1}+\left(T-\lambda\mathfrak{1}\right)^{-1}B\right)^{-1}\right\|$$\leq$ $\sum_{n=0}^{\infty}\left\|\left(T-\lambda\mathfrak{1}\right)^{-1}B\right\|^{n}=\frac{1}{1-\left\|\left(T-\lambda\mathfrak{1}\right)^{-1}B\right\|}$
$\leq$ $\frac{1}{1-\Delta\left(\lambda;T\right)^{-1}\left\|B\right\|}=\frac{\Delta\left(\lambda;T\right)}{\Delta\left(\lambda;T\right)-\left\|B\right\|}\ .$
Finally, one uses the equality
$T+B-\lambda\mathfrak{1}=\left(T-\lambda\mathfrak{1}\right)\left(\mathfrak{1}+\left(T-\lambda\mathfrak{1}\right)^{-1}B\right)$
for any $\lambda\in\rho(T)$ to deduce that $\lambda\in\rho(T+B)$ and
$\left\|\left(T+B-\lambda\mathfrak{1}\right)^{-1}\right\|=\left\|\left(\mathfrak{1}+\left(T-\lambda\mathfrak{1}\right)^{-1}B\right)^{-1}\left(T-\lambda\mathfrak{1}\right)^{-1}\right\|\leq\frac{1}{\Delta\left(\lambda;T\right)-\left\|B\right\|}\ .$
We can now prove the following version of Combes-Thomas estimates:
Theorem 5.13 (Combes-Thomas estimates)
Let $T\in\mathcal{B}(\ell^{2}(\Lambda))$ be a self-adjoint operator. Given $\mu\in\mathbb{R}_{0}^{+}$ and $\lambda\in\mathbb{C}$ with $\Delta(\lambda;T)>\mathbf{S}(T,\mu)$, the following inequality holds true:
$\left|\left\langle\mathfrak{e}_{x},\left(T-\lambda\mathfrak{1}\right)^{-1}\mathfrak{e}_{y}\right\rangle\right|\leq\frac{\mathrm{e}^{-\mu d\left(x,y\right)}}{\Delta\left(\lambda;T\right)-\mathbf{S}\left(T,\mu\right)}\ ,\qquad x,y\in\Lambda\ .$
Proof. Fix $y\in\Lambda$ and $R\in\mathbb{R}^{+}$. Define the function $\varphi:\Lambda\rightarrow[1,\mathrm{e}^{\mu R}]$ by
$\varphi\left(x\right)\doteq\exp\left(\mu\min\left\{d\left(x,y\right),R\right\}\right)\ ,\qquad x\in\Lambda\ .$
Clearly, $\varphi$ and $1/\varphi$ are bounded and the inverse of the multiplication operator $M_{\varphi}\in\mathcal{B}(\ell^{2}(\Lambda))$ by $\varphi$ is nothing else than $M_{1/\varphi}$. Because $\varphi$ is a real-valued function, $M_{\varphi}^{\ast}=M_{\varphi}$ and, for any $x\in\Lambda$, $\mathfrak{e}_{x}$ is of course an eigenvector of $M_{\varphi}$ with associated eigenvalue $\varphi(x)$. In particular,
$\left\langle\mathfrak{e}_{x},M_{\varphi}TM_{\varphi}^{-1}\mathfrak{e}_{z}\right\rangle=\frac{\varphi\left(x\right)}{\varphi\left(z\right)}\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{z}\right\rangle\ ,\qquad x,z\in\Lambda\ .$
Since $(x,z)\mapsto\min\left\{d\left(x,z\right),R\right\}$ is another pseudometric on $\Lambda$, for all $x,z\in\Lambda$, we have that
$\min\left\{d\left(x,y\right),R\right\}-\min\left\{d\left(z,y\right),R\right\}\leq\min\left\{d\left(x,z\right),R\right\}\leq d\left(x,z\right).$
In particular, the operator $B\doteq M_{\varphi}TM_{\varphi}^{-1}-T$ satisfies the bound
$\left|\left\langle\mathfrak{e}_{x},B\mathfrak{e}_{z}\right\rangle\right|\leq\left(\mathrm{e}^{\mu d\left(x,z\right)}-1\right)\left|\left\langle\mathfrak{e}_{x},T\mathfrak{e}_{z}\right\rangle\right|\ ,\qquad x,z\in\Lambda\ .$
By Lemma 5.11 together with Equations (144) and (145), it follows that
$\left\|B\right\|_{{\mathrm{o}\mathrm{p}}}\leq\mathbf{S}\left(T,\mu\right)<\Delta\left(\lambda;T\right)\ .$
Applying now Lemma 5.12, we then arrive at the bound
$\left\|\left(M_{\varphi}TM_{\varphi}^{-1}-\lambda\mathfrak{1}\right)^{-1}\right\|_{{\mathrm{o}\mathrm{p}}}\leq\frac{1}{\Delta\left(\lambda;T\right)-\left\|B\right\|_{{\mathrm{o}\mathrm{p}}}}\leq\frac{1}{\Delta(\lambda;T)-\mathbf{S}(T,\mu)}\ .$
Finally, for any $x\in\Lambda$ such that $d(x,y)\leq R$ we observe from the last upper bound that
$\mathrm{e}^{\mu d\left(x,y\right)}\left|\left\langle\mathfrak{e}_{x},\left(T-\lambda\mathfrak{1}\right)^{-1}\mathfrak{e}_{z}\right\rangle\right|=\frac{\varphi\left(x\right)}{\varphi\left(1\right)}\left|\left\langle\mathfrak{e}_{x},\left(T-\lambda\mathfrak{1}\right)^{-1}\mathfrak{e}_{z}\right\rangle\right|=\left|\left\langle\mathfrak{e}_{x},\left(M_{\varphi}TM_{\varphi}^{-1}-\lambda\mathfrak{1}\right)^{-1}\mathfrak{e}_{z}\right\rangle\right|\\ \leq\frac{1}{\Delta(\lambda;T)-\mathbf{S}(T,\mu)}\ .$
Since $y\in\Lambda$ and $R\in\mathbb{R}^{+}$ are arbitrary parameters, the above inequality in fact holds true for all $x,y\in\Lambda$.
5.6 Elementary Observations
For completeness and the reader’s convenience, we conclude the appendix by given a few elementary results related to the space of bounded operators on a Hilbert space.
Proposition 5.14 (Monotonicity of the inverse on operators)
Let $B$ and $C$ be two positive bounded operators on a Hilbert space with bounded (positive) inverse. If $B\leq C$ then $C^{-1}\leq B^{-1}$.
Proof. The proof is standard. Since it is very short, we give it for completeness. If $B$ and $C$ commute, then
$B^{-1}-C^{-1}=B^{-1}\left(C-B\right)C^{-1}\geq 0\ ,$
because $B^{-1}$, $C^{-1}$ and $C-B$ are positive commuting operators. If $B$ and $C$ do not commute then we observe that $B\leq C$ yields
$C^{-\frac{1}{2}}BC^{-\frac{1}{2}}\leq\mathfrak{1}\text{ },$
which in turn implies that
$C^{\frac{1}{2}}B^{-1}C^{\frac{1}{2}}\geq\mathfrak{1}\text{ },$
because $\mathfrak{1}$ and $C^{-\frac{1}{2}}BC^{-\frac{1}{2}}$ commute. From the last inequality it follows that $B^{-1}\geq C^{-1}$.
Proposition 5.15 (Monotone convergence theorem for operators)
Let $\mathcal{X}$ be any complex Hilbert space. Any increasing (decreasing) monotone net $(A_{i})_{i\in I}$ of self-adjoint elements in $\mathcal{B}(\mathcal{X})$ that is bounded from above (below) has a supremum (infimum) in $\mathcal{B}(\mathcal{X})$. The supremum (infimum) is itself also self-adjoint and is the strong operator limit of the net.
Proof. [40, Proposition 2.17] already tells us that any increasing (decreasing) monotone net $(A_{i})_{i\in I}$ of self-adjoint elements in $\mathcal{B}(\mathcal{X})$ that is bounded from above (below) has a supremum (infimum) $A_{\infty}$ in $\mathcal{B}(\mathcal{X})$. The supremum (infimum) $A_{\infty}$ is itself also self-adjoint and is the weak operator limit of the net. This is proven by using the polarization identity and the Riesz representation theorem together with elementary estimates. To conclude the proof, it remains to show that $A_{\infty}\in\mathcal{B}(\mathcal{X})$ is the limit of the increasing net $(A_{i})_{i\in I}$ also in the strong operator topology. To this end, assume that $(A_{i})_{i\in I}$ is an increasing net and define the (decreasing) net $(B_{i})_{i\in I}$ of positive operators by $B_{i}\doteq A_{\infty}-A_{i}\geq 0$. By construction, this net converges in the weak operator topology to $0\in$ $\mathcal{B}(\mathcal{X})$, which in turn implies that the net $(B_{i}^{1/2})_{i\in I}$ converges in the strong operator topology to $0\in$ $\mathcal{B}(\mathcal{X})$. As the net $(B_{i}^{1/2})_{i\in I}$ is norm-bounded and $B_{i}=B_{i}^{1/2}B_{i}^{1/2}$, we then conclude that also $(B_{i})_{i\in I}$ converges in the strong operator topology to $0\in$ $\mathcal{B}(\mathcal{X})$, that is, $A_{\infty}$ is the strong operator limit of the net $(A_{i})_{i\in I}$. If $(A_{i})_{i\in I}$ is a decreasing net then we consider the increasing net $(-A_{i})_{i\in I}$ to conclude that $(A_{i})_{i\in I}$ has a infimum $A_{\infty}=A_{\infty}^{\ast}\in\mathcal{B}(\mathcal{X})$, which is again the strong operator limit of the net $(A_{i})_{i\in I}$.
Acknowledgments: This work is supported by the COST Action CA18232 financed by the European Cooperation in Science and Technology (COST) and also by the Basque Government through the grant IT1615-22 and BERC 2022-2025 program and by the Ministry of Science and Innovation: PID2020-112948GB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”.